GNU Radius Reference Manual

GNU Radius Reference Manual

version 1.3, 21 July 2004

Sergey Poznyakoff

This manual documents GNU Radius (version 1.3, 21 July 2004). Published by:

GNU Press  Website: www.gnupress.org
a division of the  General: press@gnu.org
Free Software Foundation  Orders: sales@gnu.org
59 Temple Place, Suite 330  Tel: 617-542-5942
Boston, MA 02111-1307 USA  Fax: 617-542-2652

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999, 2000, 2001, 2002, 2003 Free Software Foundation

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.2 or any later version published by the Free Software Foundation; with no Invariant Sections, with the Front-Cover texts being “A GNU Manual”, and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled “GNU Free Documentation License”.

(a) The FSF’s Back-Cover Text is: “You have freedom to copy and modify this GNU Manual, like GNU software. Copies published by the Free Software Foundation raise funds for GNU development.”

Cover art by Etienne Suvasa. Cover design by Jonathan Richard.

Introduction to Radius

GNU Radius is a software package that provides authentication and accounting services. The acronym radius stands for Remote Authentication Dial In User Service and (in that form) usually denotes the underlying protocol name.

Historically, radius servers were used as a means to authenticate the user coming from a dial-in connection, but GNU Radius is much more than an authentication system: it is an advanced, customizable, and extensible system for controlling access to the network.

GNU Radius has several built-in authentication and accounting methods. When these methods are not enough, it allows the administrator to implement any new method she deems convenient.

The GNU Radius package includes the server program, radiusd, which responds to authentication and accounting requests, and a set of accompanying programs designed to monitor the activity of the server and analyze the information it provides.

Overview

To illustrate what GNU Radius does, let’s consider an imaginary internet service provider. Our provider has two network access servers (nases for short)—i.e., two pieces of equipment which directly accept users’ connections—and a core router that connects the ISP’s internal network with the Internet backbone.

When a user connects to a nas, the server must verify that the user is actually registered and that the credentials she has supplied are correct. This first step is called authentication.

Upon authenticating the user, the nas must determine which services the user is permitted to use and to what extent the user may use them. This second step is called authorization.

When the first two stages have been successfully completed, the nas takes the third step and establishes the connection between the user and the main server. This connection is called a user session. For the purposes of accounting, the nas remembers the exact time of the start of the session. When the session is terminated, the duration of the session and the number of bytes transferred are recorded as well.

All three tasks can be accomplished by the use of user and accounting databases on each terminal server. However, this is not convenient, and it is error-prone in that the maintenance of separate databases for the same users is not a trivial task. What is worse, as the number of terminal servers grows, this maintenance problem becomes more difficult.

How Does radius Perform These Tasks?

radius allows an administrator to keep authentication and accounting data in a single place, no matter how many network access servers are actually present. Using radius, nases instead communicate with this central server to perform authentication and accounting, thus easing the burden on the system administrator.

Let’s return to our imaginary ISP. Suppose it runs a radius daemon on its central server. Each nas runs client software to communicate with the radius server by sending radius packets.

An average user session life cycle looks as follows.

A user connects to the nearest nas and supplies his login and password. The nas forms an authentication request and sends it to the radius server.

The radius server verifies the user’s credentials and finds them sufficient. It then retrieves the user’s authorization information from its database, packages it into an acknowledgement packet, and then sends it back to the nas

The nas receives the acknowledgement packet and starts the user session. The information brought with the packet tells the nas to establish a connection between the core router and the user, and to assign the user a certain IP address. Having established the session, the nas informs the radius server by sending it an accounting start packet. The server acknowledges the receipt of the accounting packet.

Now suppose that after some time the user decides to break the connection. The nas notices this and terminates the user’s session. The nas then sends an accounting stop packet to the radius server to mark this event. Again, the server acknowledges the receipt of the packet.

radius Attributes

Attributes are means of passing the information between the nas and the server. Basically, an attribute is an integer number that identifies some piece of information. A set of properties are associated with each attribute, specifying the way to interpret the attribute. The most important property is the data type, which declares the type of data that the attribute identifies (character string, integer number, IP address, or raw binary data).

The information to be transmitted with the request is packaged in a set of attribute-value pairs (or a/v pairs for short). Such pairs consist of attribute numbers and the associated data.

radius Packets

There exist two basic kinds of radius packets: authentication and accounting packets. Each of them is subdivided into requests and replies.

Authentication requests are sent from the nas to the radius server and contain the information necessary to check the identity of the user. The minimum set of data in such packets consists of the user login name, user password, and nas IP or identifier.

Authentication replies are sent by the radius server and contain the reply code and a set of additional attributes. According to their reply code the authentication replies are subdivided into authentication acknowledgements, authentication rejections, and authentication challenges.

An authentication acknowledgement packet is sent to the nas if the credentials supplied with the authentication request were correct. This kind of packet tells the nas to establish a normal user session. The additional attributes in such packets carry the authorization data, i.e., they determine which kind of service the user is to be provided.

An authentication rejection is sent to the nas if the authentication has failed. This packet forbids the nas to provide any service to the user. The additional attributes may carry descriptive text to be displayed as an explanation to the user for the failure of his request.

Finally, an authentication challenge packet is sent to the nas if the supplied credentials did not suffice to establish the authenticity of the user. This means that the dialog between the nas and the radius server continues. As the radius server asks for additional authentication credentials, the nas acts as a liaison, passing server requests to the user and sending user replies back to the server. Such a dialog ends when the radius server sends either an acknowledgement packet or a rejection packet.

An accounting request is sent to the server when the nas wishes to report some event in the user session: the start of the session, session termination, etc. The attributes carry the actual information about the event.

For each accounting request that has been received and successfully processed, the radius server sends back an accounting acknowledgement. This packet carries no attributes, but simply informs the nas that the information it had sent was received.

Occasionally, a radius server may fail to receive incoming requests or may fail to process them due to high server load. In order to prevent such requests from being lost, the nas retransmits the request if no response from the server is received within a predefined interval of time (a timeout interval). Usually the nas is configured in such a way that it continues retransmitting failed requests until either it receives a reply from the server or a predefined number of retries are exhausted, whichever occurs first. Furthermore, a nas may be configured to communicate with a set of backup radius servers. In this case it applies the described process to each server from the set, until one of them responds or the set is exhausted.

4 GNU Radius Reference Manual

1 Naming Conventions

This chapter describes file naming conventions used throughout this document.

Programs from the GNU Radius package use the following directories to store various configuration and log files:

Configuration or database directory A directory where all configuration files are stored.

Log directory A directory where radiusd stores its log files.

Accounting directory A directory where radiusd stores accounting detail files (see Section 7.2 [Detailed Request Accounting], page 81).

Data directory A directory where shared data files are stored, such as Rewrite or Scheme source files.

The default locations of these directories are determined at compile time. Usually these are:

Directory  Short name Default location
Configuration directory  ‘raddb’  /usr/local/etc/raddb
Log directory  ‘radlog’  /var/log
Accounting directory  ‘radacct’  /var/log/radacct
Data directory  ‘datadir’  /usr/local/share/radius/1.3

These locations may differ depending on your local site configuration. Throughout this document we will refer to these directories by their short

names. For example, when we say:

     ... this information is contained in file `raddb/sqlserver'
we actually mean ‘/usr/local/etc/raddb/sqlserver’.

To get the default directory names that your version of Radius was com-

piled with, run radiusd --version.

Locations of these directories may be overridden by specifying the appropriate command line options. For example, any program from the GNU Radius package accepts the command line option ‘-d’ or ‘--directory’, which introduces the configuration directory path.

6 GNU Radius Reference Manual

2 How Radius Operates

The main purpose of GNU Radius is to centralize authentication of users coming from various network stations, pursuant to the radius specification. Its primary usage is for dial-in users, though it can be used for any kind of network connection.

2.1 Attributes

Information carried by radius requests is stored as a list of attribute-value pairs. Each pair consists of an attribute number and an attribute value. The attribute number identifies the type of information the pair carries, and the attribute value keeps the actual data.

The value part of an attribute can contain data of one of the following types: Integer A 32-bit unsigned integer value. IP-number

An IPv4 IP-number.

String A character string up to 253 characters long.

For convenience, the attributes and the values of some frequently used integer attributes are given symbolic names. These names are assigned to attributes and values in the dictionary file (see Section 4.2 [dictionary file], page 39).

Attribute numbers range from 1 to 255. Attributes with numbers greater than 255 are used internally by the server and cannot be sent to the nas.

The vendor-specific attribute number 26 is special, allowing vendors of the nas hardware or software to support their own extended attributes. Section 13.1.26 [Vendor-Specific], page 175.

Each attribute has a set of properties associated with it. The properties are: Usage flags

These flags determine the usage of the attribute in the configuration files ‘huntgroups’, ‘hints’, and ‘users’.

Propagation When a radius server functions in proxy mode, it uses the propagation flag to determine which attributes from the reply packet should be passed back to the requesting nas (see Section 2.4.2.1 [Proxy Service], page 13).

additivity Some configuration rules may cause the addition of new a/v pairs to the incoming request. Before the addition of a new pair, radiusd scans the request to see if it already contains a pair with the same attribute. If it does, the value of the additivity determines the following additional actions:

None The old pair is retained in the request; the new pair is not added to it.

Replace The old pair is retained in the request, but its value is replaced with that of the new pair.

Append The new pair is appended to the end of the pair list.

Attributes are declared in the ‘raddb/dictionary’ file. For a detailed description, see Section 4.2.4 [ATTRIBUTE], page 41. For information about particular attributes, see Chapter 13 [Attribute List], page 165.

2.2 radius Requests

The term request refers to both the authentication/accounting request packet from a nas to a radius server and the response packet that the server sends back to the nas.

Each request contains the following fields:

‘Code’ The code field identifies the type of the request.

‘Identifier’

The number in the range 0–255 used to match the request with the reply.

‘Length’ The length of the request packet.

‘Authenticator’

The 16-byte hash value used to authenticate the packet.

‘Attributes’

The list of attribute-value pairs carrying actual information about the request.

2.2.1 Authentication Requests

A nas sends authentication requests (packets with code field set to Access-Request) to a radius server when a user is trying to connect to that nas. Such requests convey information used to determine whether a user is allowed access to the nas, and whether any special services are requested for that user.

An Access-Request must contain a User-Name attribute Section 13.1.24 [User-Name], page 174. This packet should contain a NAS-IP-Address attribute, a NAS-Identifier attribute, or both. It also must contain either a User-Password attribute or a CHAP-Password attribute. These attributes are passed after being encoded using a method based on the RSA Message Digest Algorithm MD5.

The Access-Request should contain a NAS-Port or NAS-Port-Type attribute or both, unless the type of access being requested does not involve a port or the nas does not distinguish among its ports.

Upon receiving an Access-Request packet for a particular user and authenticating that user, the radius server replies to the nas that has sent the packet with any one of the following packets:

• Access-Accept

• Access-Reject

• Access-Challenge

GNU Radius replies with an Access-Accept packet when it has successfully authenticated the user. Such a reply packet provides the configuration information necessary to begin delivery of service to the user.

GNU Radius replies with an Access-Reject packet when it is unable to authenticate the user. Such a packet may contain a descriptive text encapsulated in one or more Reply-Message attributes. The nas may display this text along with its response to the user.

GNU Radius replies with an Access-Challenge packet when it needs to obtain more information from the user in order to determine the user’s authenticity or to determine the kind of service to be provided to the user.

An Access-Challenge packet may include one or more Reply-Message attributes, and it may or may not include a single State attribute. No other attributes are permitted in an Access-Challenge packet.

Upon receipt of an Access-Challenge, the Identifier field is matched with a pending Access-Request. Additionally, the Response Authenticator field must contain the correct response for the pending Access-Request. In the event of an invalid packet, GNU Radius discards the offending packet and issues the appropriate log message.

If the nas does not support challenge/response, it treats an Access-Challenge as though it had received an Access-Reject instead. Otherwise, upon receipt of a valid Access-Challenge the nas prompts the user for a response, possibly displaying the text message provided in the Reply-Message attributes of the request. It then sends its original Access-Request with a new request id and request authenticator, along with the User-Password attribute replaced by the encrypted user’s response, and including the State attribute from the Access-Challenge, if any.

2.2.2 Accounting Requests

Accounting-Request packets are sent from a nas to a radius server to allow for accounting of a service provided to a user.

Upon receipt of an Accounting-Request packet, the server attempts to record it (see Chapter 7 [Accounting], page 81), and if it succeeds in doing so, it replies with an Accounting-Response packet. Otherwise, it sends no reply, which then causes the nas to retransmit its request within a preconfigured interval of time. Such retransmits will continue until either the server responds with an Accounting-Response packet or a preconfigured number of retransmits is reached, whichever occurs first.

Any attribute valid in an Access-Request or Access-Accept packet is also valid in an Accounting-Request packet, except the following attributes, which are never present in any Accounting-Request packet:

• User-Password
• CHAP-Password
• Reply-Message
• State
Either a NAS-IP-Address or a NAS-Identifier must be present in an

Accounting-Request packet. It should contain either a NAS-Port or a NAS-Port-Type attribute (or both), unless the service does not involve a port or the nas does not distinguish among its ports.

If the Accounting-Request packet includes a Framed-IP-Address, that attribute must contain the actual IP of the user.

There are five types of accounting packets, differentiated by the value of the Acct-Status-Type attribute. These are:

Session Start Packet The session start packet is sent after the user has successfully passed the authentication and has started to receive the requested service. It must contain at least following attributes:

  • Acct-Status-Type = Start
  • User-Name
  • Acct-Session-Id
  • NAS-IP-Address
  • NAS-Port-Id

Session Stop Packet The session stop packet is sent after the user has disconnected. It conveys the information about the duration of the session, number of octets transferred, etc. It must contain at least the following attributes:

  • Acct-Status-Type = Stop
  • User-Name
  • NAS-IP-Address
  • Acct-Session-Id

The last three of them are used to find the corresponding session start packet.

Keepalive Packet The keepalive packet is sent by the nas when it obtains some new information about the user’s session, e.g. it has determined its IP or has changed the connection speed. The packet must contain at least the following attributes:

  • Acct-Status-Type = Alive
  • User-Name
  • NAS-IP-Address
  • Acct-Session-Id

Accounting-Off Packet By sending this packet, the nas requests that radiusd mark all sessions registered from this particular nas as finished. Receiving this packet usually means that the nas is to be shut down, or is about to change its configuration in a way that requires all currently opened sessions to be closed. The packet must contain at least the following attributes:

  • Acct-Status-Type = Accounting-Off
  • NAS-IP-Address

Accounting-On Packet By sending this packet, the nas informs radiusd that it is ready to accept the incoming connections. Usually this packet is sent after startup, or after a major reconfiguration of the nas. It must contain at least the following attributes:

  • Acct-Status-Type = Accounting-On
  • NAS-IP-Address

2.3 Matching Rule

A record in the GNU Radius database describing a particular rule for matching an incoming request is called a matching rule. Each such rule defines an action to be taken when the match occurs.

The matching rule consists of three distinct parts:

Label This is used to identify the rule. The special usernames DEFAULT and BEGIN are reserved. These will be described in detail below.

Left-Hand Side (lhs)

The list of attribute-value pairs used for matching the profile against an incoming request.

Right-Hand Side (rhs)

The list of attribute-value pairs that define the action to be taken if the request matches lhs.

The following GNU Radius configuration files keep data in a matching rule format: ‘hints’, ‘huntgroups’, and ‘users’. Although they keep data in a similar format, the rules that are used to match incoming requests against the contents of these files differ from file to file. The following section describes these rules in detail.

2.4 Processing Requests

Upon receiving a request, radiusd applies to it a number of checks to determine whether the request comes from an authorized source. If these checks succeed, the request is processed and answered. Otherwise, the request is dropped and corresponding error message is issued (see Chapter 8 [Logging], page 85).

The following checks are performed:

Check if the username is supplied.

If the packet lacks the User-Name attribute, it is not processed.

Check if the nas is allowed to speak.

The source IP of the machine that sent the packet is looked up in the ‘clients’ file (see Section 4.3 [clients file], page 44). If no match is found, the request is rejected.

Compute the encryption key.

Using the data from the packet and the shared key value from the ‘clients’ file, Radius computes the MD5 encryption key that will be used to decrypt the value of the User-Password attribute.

Process user-name hints.

User-name hints are special rules that modify the request depending on the user’s name and her credentials. These rules allow an administrator to divide users into distinct groups, each group having its own authentication and/or accounting methods. The user-name hints are stored in ‘raddb/hints’ (see Section 4.6 [hints file], page 49).

Process huntgroup rules.

Huntgroup rules allow an administrator to segregate incoming requests depending on the nas and/or port number they came from. These rules are stored in ‘raddb/huntgroups’ (see Section 4.7 [huntgroups file], page 50).

Determine whether the request must be proxied to another radius server. The requests pertaining to another realm are immediately forwarded to the remote radius server for further processing. See Section 2.4.2 [Proxying], page 13, for the description of this process.

Process individual user profiles This step applies only to authentication requests.

2.4.1 Checking for Duplicate Requests

As described above (see Chapter 2 [Operation], page 7), a nas may decide to retransmit the request under certain circumstances. This behavior ensures that no requests are lost. For example, consider the following scenario:

1. The nas sends a request to the server.

2. The server processes it and sends back the reply.

3. The reply is lost due to a network outage, or the load average of the nas is too high and it drops the response.

4. The nas retransmits the request.

Thus the radius server will receive and process the same request twice. This probably won’t do any harm if the request in question is an authentication one, but for accounting requests it will lead to duplicate accounting. To avoid such an undesirable effect, radiusd keeps a queue of received requests. When an incoming request arrives, radiusd first scans the request queue to see if the request is a duplicate. If so, it drops the request; otherwise, it inserts the request into the queue for processing. After the request is completed, it will still reside in the queue for a preconfigured interval of time (see Section 4.1.3 [auth], page 28, parameter request-cleanup-delay).

By default, radiusd considers two requests to be equal if the following conditions are met:

1. Both requests come from the same nas.

2. They are of the same type.

3. The request identifier is the same for both requests.

4. The request authenticator is the same for both requests.

Additionally, radiusd may be configured to take into account the contents of both requests. This may be necessary, since some nases modify the request authenticator or request identifier before retransmitting the request, so the method described above fails to recognize the request as a duplicate. This extended comparison is described in detail in Section 5.1 [Extended Comparison], page 67.

2.4.2 Proxying

Proxying is a mode of operation where a radius server forwards incoming requests from a nas to another radius server, waits for the latter to reply, and then forwards the reply back to the requesting nas. A common use for such operation mode is to provide roaming between several internet service providers (ISPs). Roaming permits ISPs to share their resources, allowing each party’s users to connect to other party’s equipment. Thus, users traveling outside the area of one ISP’s coverage are still able to access their services through another ISP.

2.4.2.1 Proxy Service

Suppose the ISP ‘Local’ has a roaming arrangement with the ISP ‘Remote’. When the user of ‘Remote’ dials in to the nas of ‘Local’, the nas sends the authentication request to the ‘Local’ radius server. The server then determines that this is a roaming user, stores a copy of the request in its internal queue, and forwards the request to the ‘Remote’ radius server for processing. Thus, the ‘Local’ radius server acts as a client for the ‘Remote’ radius server.

When the ‘Remote’ radius server responds, the ‘Local’ radius server receives the response, and passes it back to the nas. The copy of the request from the server’s queue determines which nas originated the request. Before passing the request back to the nas, the server removes information

specific to the ‘Remote’ site, such as Framed-IP-Address, Framed-Netmask,

etc. Only the attributes marked with a ‘propagation’ flag (see Section 2.1 [Attributes], page 7) are passed back to the nas. After removing site-specific attributes, the ‘Local’ radius server passes the request through its user profiles (see Section 2.4.5 [User Profiles], page 16) to insert any local, site-specific information that might be needed. Finally, it passes the reply back to the nas.

Proxied accounting requests are processed in a similar manner, except that no attribute filtering takes place, as accounting responses do not carry any a/v pairs.

This example illustrates only the simplest proxy chain, consisting of two servers; real-life proxy chains may consist of several servers. For example, our ‘Remote’ radius server might also act as a proxy, forwarding the request to yet another radius server, and so on.

Note that when the accounting request passes through a chain of forwarding servers, the accounting records are stored on all servers in the chain.

2.4.2.2 Realms

GNU Radius determines which server a request must be forwarded to by the request’s authentication realm. There are three kinds of realms:

1. A named realm is the part of a user name following the at sign (‘@’). For example, if the user name is ‘jsmith@this.net’, then ‘this.net’ is the realm. The named realms can be cascaded; e.g., a request with user name ‘jsmith@this.net@remote.net’ will first be forwarded to the radius server of the realm ‘remote.net’, which in turn will forward

    it to ‘this.net’.

2. A default realm defines the server to which the requests for realms not mentioned explicitly in the configuration are forwarded.

3. An empty realm defines the server to which the requests without explicitly named realms are forwarded. If the configuration does not define an empty realm, such requests are processed by the server itself.

GNU Radius keeps the information about the realms it serves in the ‘raddb/realms’ configuration file (see Section 4.8 [realms file], page 50).

2.4.3 Hints

User-name hints are special rules that modify the incoming request depending on the user name and its credentials. Hints are stored as a list of matching rules (see Section 2.3 [Matching Rule], page 11). Upon receiving a request, radiusd scans the hint entries sequentially, comparing each rule’s label with the value of the User-Name attribute from the request. If they coincide, then radiusd appends the contents of the rule’s rhs to the request’s pair list.

The two user names must match exactly in order for a hint to take effect, unless the hint’s checklist contains either the Prefix or the Suffix attribute. The special name ‘DEFAULT’ or ‘DEFAULT%d’ (where %d denotes any decimal number), used as a hint’s label, matches any user name.

Two special attributes, Prefix and Suffix, may be used in lhs to restrict the match to a specified part of a user name. Both are string attributes. The Prefix instructs radiusd to accept the hint only if the user name begins with the given prefix. Similarly, Suffix instructs radiusd to accept the hint only if the user name ends with the given suffix. A hint may contain both Prefix and Suffix attributes.

In addition to these two attributes, a hint’s lhs may contain User-ID and Group attributes.

The following attributes, when used in a hint’s rhs have special meaning. They are not appended to the request pair list. Instead, they are removed after completing their function:

Fall-Through

If this attribute is present and is set to Yes, radiusd continues scanning the hints after processing the current entry. This allows radiusd to apply several hints to a single packet.

Rewrite-Function

If this attribute is present, the specified rewrite function is invoked.

Replace-User-Name

The value of this attribute is expanded (see Section 4.14 [Macro Substitution], page 64) and replaces the value of the User-Name attribute from the request.

Hint rules are defined in the ‘raddb/hints’ file (see Section 4.6 [hints file], page 49).

2.4.4 Huntgroups

Huntgroups are special rules that allow an administrator to provide alternate processing of certain incoming requests depending on the nas IP and port number they come from. These rules are stored as a list of matching rules (see Section 2.3 [Matching Rule], page 11).

Upon receiving a request, radiusd scans this list sequentially until it finds an entry such that the conditions set forth in its lhs are matched by the request. If such an entry is found, radiusd verifies that the request meets the conditions described by rhs. If it does not, the request is rejected. In short, a huntgroup requires that any request matching its lhs must match also its rhs.

The label part of the rule is not used in comparisons; instead it is used to label huntgroups. All entries with the same label form a single huntgroup. The special attribute Huntgroup-Name can be used to request a match against a particular huntgroup (see Section 13.3.12 [Huntgroup-Name], page 186).

Huntgroup rules are defined in the ‘raddb/huntgroups’ file (see Section 4.7 [huntgroups file], page 50).

2.4.5 User Profiles

User profiles are per-user matching rules (see Section 2.3 [Matching Rule], page 11). All incoming authentication requests are compared with the user profiles after they have passed both hints and huntgroups. radiusd selects the user profiles whose label matches the value of the User-Name attribute from the incoming request.

The selected profiles form the list of authentication rules for the request. In order for a profile to be selected, its label must either coincide literally with the User-Name value, or be one of the special labels, DEFAULT or BEGIN.

Rules in an authentication list are ordered as follows: first go all the profiles with the BEGIN label, followed by the profiles whose labels match the User-Name literally, followed finally by the rules labeled with the DEFAULT.^{1}

Within each of the three sublists, the rules preserve the order in which they appear in the ‘raddb/users’ file. Once the list is constructed, it is scanned sequentially until the rule is found whose lhs matches the incoming request. If no such rule is found, the authentication fails. Otherwise, the contents of its rhs are appended to the reply list being constructed. If the rhs of the matched rule contains the attribute Fall-Through with the value Yes, the matching continues. When the list is exhausted, the authentication result is sent back to the nas along with the a/v pairs collected in the reply list.

User profiles are defined in the ‘raddb/users’ file (see Section 4.9 [users file], page 52).

1 For compatibility with other radius implementations, GNU Radius treats profile labels

in the form DEFAULT%d, where %d represents a decimal number, in the same way it

treats DEFAULT labels. The same applies to BEGIN labels.

3 How to Start the Daemon.

When started radiusd uses the configuration values from the following sources (in order of increasing precedence):

• Compiled-in defaults

  • ‘raddb/config’ file.

• Command line arguments

Whenever a command line options has its equivalent in config file the use of this equivalent should be preferred (see Section 4.1 [config file], page 22).

The following command line options are accepted:

‘-A’
‘--log-auth-detail’

Enable detailed authentication logging. When this option is specified each authentication request is logged to the

          file ‘radacct/NASNAME/detail.auth’, where NASNAME is

replaced by the short name of the nas from ‘raddb/naslist’ Chapter 1 [Naming Conventions], page 5. Config file equivalent: auth { detail yes; };.

‘-a DIR’
‘--acct-directory DIR’

Specify accounting directory.

          Config file equivalent: option { acct-dir DIR; };.
‘-b’

‘--dbm’ Enable DBM support.

Config file equivalent: usedbm yes;.

‘-d DIR’
‘--config-directory DIR’
‘--directory D’

Specify alternate configuration directory. Default is

          ‘/usr/local/etc/raddb’.
‘-f’
‘--foreground’

Stay in foreground. We recommend to use it for debugging purposes only.

‘-i IP’
‘--ip-address’

Specifies the ip address radiusd will listen on. If this option is not specified, the program will listen on all IP addresses, assigned to the machine it runs on.

          Config file equivalent: option { source-ip IP; };.

Note that listen statement in ‘raddb/config’ provides a better control over ip addresses to listen on (see Section 4.1.3 [auth], page 28, and see Section 4.1.4 [acct], page 30).

‘-L’
‘--license’

Display GNU General Public License and exit.

‘-l DIR’
‘--logging-directory DIR’

Specify alternate logging directory.

          Config file equivalent: option { log-dir DIR; };.
‘-mb’

‘--mode b’ “Builddbm” mode. Builds a DBM version of a plaintext users database. Section 11.8 [Builddbm], page 130.

‘-mc’

‘--mode c’ Check configuration files and exit. All errors are reported via usual log channels.

‘-mt’

‘--mode t’ Test mode. In this mode radiusd starts an interactive interpreter which allows to test various aspects of its configuration.

‘-N’
‘--auth-only’

Process only authentication requests.

‘-n’
‘--do-not-resolve’

Do not resolve IP addresses for diagnostic output. This can reduce the amount of network traffic and speed up the server. Config file equivalent: option { resolve no };.

‘-p PORTNO’
‘--port PORTNO’

Listen the udp port PORTNO. The accounting port is computed as PORTNO + 1.

‘-P DIR’
‘--pid-file-dir DIR’

Specifies the alternate path for the pidfile.

‘-S’
‘--log-stripped-names’

Log usernames stripped off any prefixes/suffixes.

          Config file equivalent: auth { strip-names yes };.

‘-s’
‘--single-process’

Run in single process mode. This is for debugging purposes only. We strongly recommend against using this option. Use it only when absolutely necessary.

‘-v’
‘--version’

Display program version and compilation options.

‘-x DEBUG_LEVEL’
‘--debug DEBUG_LEVEL’

Set debugging level. DEBUG LEVEL is a comma-separated list of assignments in the forms

     MODULE
     MODULE = LEVEL

where MODULE is the module name or any non-ambiguous assignment thereof, LEVEL is the debugging level in the range 0-100. Section 9.2 [Debugging], page 89 Config file equivalent:

     logging {
             category debug {
                     level DEBUG_LEVEL;
             };
     };
‘-y’
‘--log-auth’

Log authentications. With this option enabled, Radius will log any authentication attempt into its log file Chapter 8 [Logging], page 85.

Config file equivalent: logging { category auth { detail
yes; }; }; .
‘-z’
‘--log-auth-pass’

Log passwords along with authentication information. Do not use this option. It is very insecure, since all users’ passwords will be echoed in the logfile. This option is provided only for debugging purposes. Config file equivalent:

     logging {
             category auth {
                     print-pass yes;
             };
     };

See Section 4.1 [config file], page 22.

20 GNU Radius Reference Manual

4 Radius Configuration Files

At startup, GNU Radius obtains the information vital for its functioning from a number of configuration files. These are normally found in /usr/local/etc/raddb directory, which is defined at configuration time, although their location can be specified at runtime. In the discussion below we will refer to this directory by ‘raddb’. See Chapter 1 [Naming Conventions], page 5.

Each configuration file is responsible for a certain part of the GNU Radius functionality. The following table lists all configuration files along with a brief description of their purposes.

‘config’ Determines the runtime defaults for radiusd, such as the IP address and ports to listen on, the sizes of the request queues, configuration of the SNMP subsystem, fine-tuning of the extension languages, etc.

‘clients’ Lists the shared secret belonging to each nas. It is crucial for the normal request processing that each nas have an entry in this file. The requests from nases that are not listed in ‘clients’ will be ignored, as well as those from the nases that have a wrong value for the shared secret configured in this file.

‘naslist’ Defines the types for the known nases. Its information is used mainly when performing multiple login checking (see Section 6.9 [Multiple Login Checking], page 74).

‘nastypes’

Declares the known nas types. The symbolic type names, declared in this file can be used in ‘naslist’.

‘dictionary’

Defines the symbolic names for radius attributes and attribute values. Only the names declared in this file may be used in the

files ‘users’, ‘hints’ and ‘huntgroups’.
‘huntgroups’

Contains special rules that process the incoming requests basing on the nas IP and port number they come from. These can also be used as a kind of access control list.

‘hints’ Defines the matching rules that modify the incoming request depending on the user name and its credentials.

‘users’ Contains the individual users’ profiles.

‘realms’ Defines the Radius realms and the servers that are responsible for them.

‘access.deny’

A list of usernames that should not be allowed access via Radius.

‘sqlserver’

Contains the configuration for the sql system. This includes the type of sql interface used, the IP and port number of the server and the definition of the sql requests used by radiusd.

‘rewrite’ Contains the source code of functions in Rewrite extension language.

‘menus’ A subdirectory containing the authentication menus.

The rest of this chapter describes each of these files in detail.

4.1 Run-Time Configuration Options —

    ‘raddb/config’

At startup radiusd obtains its configuration values from three places. The basic configuration is kept in the executable module itself. These values are overridden by those obtained from ‘raddb/config’ file. Finally, the options obtained from the command line override the first two sets of options.

When re-reading of the configuration is initiated either by SIGHUP signal or by SNMP channel any changes in the config file take precedence over command line arguments, since ‘raddb/config’ is the only way to change configuration of the running program.

This chapter discusses the ‘raddb/config’ file in detail.

The ‘raddb/config’ consists of statements and comments. Statements end with a semicolon. Many statements contain a block of sub-statements which also terminate with a semicolon.

Comments can be written in shell, C, or C++ constructs, i.e. any of the following represent a valid comment:

     # A shell comment
     /* A C-style
      * multi-line comment
      */
     // A C++-style comment

These are the basic statements:

4.1.1 option block

Syntax:

     option {
             source-ip number ;
             max-requests number ;
             radiusd-user string ;
             exec-program-user string ;
             username-chars string ;
             log-dir string ;
             acct-dir string ;
             resolve bool ;

max-processes number ;
process-idle-timeout number ;
master-read-timeout number ;
master-write-timeout number ;
} ;

Usage

The option block defines the global options to be used by radiusd.

Boolean statements

resolve Determines whether radius should resolve the IP addresses for diagnostic output. Specifying resolve no speeds up the server and reduces the network traffic.

Numeric statements

source-ip

Sets the source ip address. When this statement is not present, the ip address of the first available network interface on the machine will be used as source.

max-requests

Sets the maximum number of the requests in queue.

max-processes

Sets the maximum number of child processes. The default value is 16. If you plan to raise this value, make sure you have enough file descriptors available, as each child occupies four descriptors for its input/output channels.

process-idle-timeout

Sets the maximum idle time for child processes. A child terminates if it does not receive any requests from the main process within this number of seconds. By default, this parameter is 3600 seconds (one hour).

master-read-timeout
master-write-timeout

These two values set the timeout values for the interprocess input/output operations in the main server process. More specifically, master-read-timeout sets the maximum number of seconds the main process will wait for the answer from the subprocess, and master-write-timeout sets the maximum number of seconds the main process will wait for the subprocess’s comunication channel to become ready for input. By default, no timeouts are imposed.

String statements

radiusd-user

Instructs radiusd to drop root privileges and to switch to the real user and group IDs of the given user after becoming daemon. Notice the following implications of this statement:

1. All configuration files must be readable for this user.

2. Authentication type System (see Section 6.5 [System Auth], page 72) requires root privileges, so it cannot be used with radiusd-user. Any ‘raddb/users’ profiles using this authentication type will be discarded.

3. Authentication type PAM (see Section 6.7 [PAM Auth], page 73) may require root provileges. It is reported to always require root privileges on some systems (notably on Solaris).

4. exec-program-user statement (see below) is ignored when

    used with radiusd-user.
exec-program-user

Sets the privileges for the programs executed as a result of Exec-Program and Exec-Program-Wait. The real user and group ids will be retrieved from the ‘/etc/passwd’ entry for the given user.

username-chars

Determines characters that are valid within a username. The alphanumeric characters are always allowed in a username, it is not necessary to specify them in this statement. By default the following characters are allowed in a username: ‘.-_!@#$%^&\/"’. The username-chars statement overrides this default, thus setting:

     username-chars ":"

will restrict the set of allowed characters to the alphanumeric characters and colon. If you wish to expand the default character set, you will have to explicitly specify it in the username-chars argument, as shown in the example below:

     username-chars ".-_!@#$%^&\\/\":"

(Notice the use of escape character ‘\’).

log-dir Specifies the logging directory.

acct-dir Specifies the accounting directory.

4.1.2 logging block

Syntax:

logging {
prefix-hook string ;

             suffix-hook string ;
             category category_spec {
                     channel channel_name ;
                     print-auth bool ;
                     print-pass bool ;
                     print-failed-pass bool ;
                     level debug_level ;
             } ;
             channel channel_name {
                     file string ;
                     syslog facility . priority ;
                     print-pid bool ;
                     print-category bool ;
                     print-cons bool ;
                     print-level bool ;
                     print-priority bool ;
                     print-tid bool;
                     print-milliseconds bool;
                     prefix-hook string ;
                     suffix-hook string ;
             };
     } ;

Usage

The logging statement describes the course followed by radiusd’s logging information.

The parts of this statement are discussed below.

4.1.2.1 Logging hooks

Most diagnostic messages displayed by radiusd describe some events that occured while processig a certain incoming request. By default they contain only a short summary of the event. Logging hooks are means of controlling actual amount of information displayed in such messages. They allow you to add to the message being displayed any relevant information from the incoming request that caused the message to appear.

A hook is a special Rewrite function that takes three arguments and returns a string. There are two kinds of logging hooks: prefix and suffix. Return value from the prefix hook function will be displayed before the actual log message, that of the suffix hook function will be displayed after the message.

Furthermore, there may be global and channel-specific hooks. Global hooks apply to all categories, unless overridden by category-specific hooks. Global prefix hook is enabled by prefix-hook statement appearing in the logging block. Global suffix hook is enabled by suffix-hook statement. Both statements take as their argument the name of corresponding Rewrite function.

For detailed information about writing logging hooks, See Section 10.2.7 [Logging Hook Functions], page 105.

4.1.2.2 category statement

Each line of logging information generated by radiusd has an associated category. The logging statement allows each category of output to be controlled independently of the others. The logging category is defined by category name and a severity. category name determines what part of radiusd daemon is allowed to send its logging information to this channel. It can be any of main, auth, acct, proxy, snmp. priority determines the minimum priority of the messages displayed by this channel. The priorities in

ascending order are: debug, info, notice, warn, err, crit, alert, emerg.

The full category specification, denoted by the category_spec in the above section, can take any of the following three forms:

category name Print the messages of given category.

priority Print messages of all categories, abridged by given priority. If the priority is prefixed with ‘=’, only messages with given priority will be displayed. If it is prefixed with ‘!’, the messages with priority other than the specified will be displayed. Otherwise, the messages with priorities equal to or greater than the specified will be displayed.

category name . priority Print the messages of given category, abridged by given priority. The priority may be prefixed with either ‘=’ or ‘!’ as described above. The dot (‘.’) separates the priority from the category name, it may be surrounded by any amount of whitespace.

Additional category options valid for auth category are:

print-auth

Log individual authentications.

print-pass

Include passwords for successful authentications. It is very insecure, since all users’ passwords will be echoed in the logfile. This option is provided only for debugging purposes.

print-failed-pass

Include passwords for failed authentications.

4.1.2.3 channel statement

Channels represent methods for recording logging information. Each channel has a unique name, and any categories which specify that name in a channel statement will use that channel.

radiusd can write logging information to files or send it to syslog. The file statement sends the channel’s output to the named file (see Chapter 1 [Naming Conventions], page 5). The syslog statement sends the channel’s output to syslog with the specified facility and severity.

Channel options modify the data flowing through the channel:

print-pid

Add the process id of the process generating the logging information.

print-cons

Also send the logging information to the system console.

print-category

Add the category name to the logging information.

print-priority
print-level

Add the priority name to the logging information.

print-milliseconds

Print timestamp with milliseconds.

prefix-hook

Declares the name of Rewrite function used as logging prefix hook for that channel (see Section 4.1.2.1 [hooks], page 25). This overrides any global prefix hook.

suffix-hook

Declares the name of Rewrite function used as logging suffix hook for that channel (see Section 4.1.2.1 [hooks], page 25). This overrides any global suffix hook.

4.1.2.4 Example of the logging statement

     logging {
             channel default {
                     file "radius.log";
                     print-category yes;
                     print-priority yes;
             };
             channel info {
                     file "radius.info";
                     print-pid yes;
                     print-cons yes;
                     print-priority yes;
             };
             channel notice {
                     syslog auth.notice;
             };
             category auth {

                     print-auth yes;
                     print-failed-pass yes;
             };
             category notice {
                     channel notice;
             };
             category info {
                     channel info;
             };
             category debug {
                     channel info;
                     level radiusd=1,files;
             };
             category *.!debug {
                     channel default;
             };
     };

4.1.3 auth statement

Syntax:

     auth {
             listen ( addr-list | no );
             forward addr-list;
             port number ;
             max-requests number ;
             time-to-live number ;
             request-cleanup-delay number ;
             detail bool ;
             strip-names bool ;
             checkrad-assume-logged bool ;
             password-expire-warning number ;
             compare-atribute-flag character ;
             trace-rules bool ;
             reject-malformed-names bool ;
     } ;

Usage:

The auth statement configures the parameters of the authentication service.

listen statement

This statement determines on which addresses radiusd will listen for incoming authentication requests. Its argument is a comma-separated list of items in the form ip:port-number. ip can be either an IP address in familiar “dotted-quad” notation or a hostname. :port-number part may be omitted, in which case the default authentication port is assumed.

If the listen statement is omitted, radiusd will accept incoming requests from any interface on the machine.

The special value no disables listening for authentication requests. The following example configures radius to listen for the incoming re-

quests on the default authentication port on the address 10.10.10.1 and on port 1645 on address 10.10.11.2.

listen 10.10.10.1, 10.10.11.2:1645;

forward statement

This statement enables forwarding of the requests to the given set of servers. Forwarding is an experimental feature of GNU Radius, it differs from proxying in that the requests are sent to the remote server (or servers) and processed locally. The remote server is not expected to reply.

This mode is intended primarily for debugging purposes. It could also be useful in some very complex and unusual configurations.

Numeric statements

port Sets the number of which udp port to listen on for the authentication requests.

max-requests

Sets the maximum number of authentication requests in the queue. Any surplus requests will be discarded.

time-to-live

Sets the request time-to-live in seconds. The time-to-live is the time to wait for the completion of the request. If the request job isn’t completed within this interval of time it is cleared, the corresponding child process killed and the request removed from the queue.

request-cleanup-delay

Sets the request cleanup delay in seconds, i.e. determines how long will the completed authentication request reside in the queue.

password-expire-warning

Sets the time interval for password expiration warning. If user’s password expires within given number of seconds, radiusd will send a warning along with authentication-acknowledge response. Default is 0.

Boolean statements

detail When set to true, radiusd will produce the detailed log of each

received packet in the file ‘radacct/nasname/detail.auth’.

The format of such log files is identical to the format of detailed accounting files (see Section 7.2 [Detailed Request Accounting], page 81).

strip-names

Determines whether radiusd should strip any prefixes/suffixes off the username before logging.

checkrad-assume-logged

See Section 4.1.11 [mlc], page 39, for the description of this setting. It is accepted in auth for compatibility with previous versions of GNU Radius.

trace-rules

Enables tracing of the configuration rules that were matched during processing of each received authentication request. See Section 9.1 [Rule Tracing], page 87, for detailed information about this mode.

reject-malformed-names

Enables sending access-reject replies for the access-accept requests that contain an invalid value in User-Name attribute. By default such requests are discarded without answering. See the description of username-chars (see Section 4.1.1 [Option statement], page 22).

Character statement

compare-attribute-flag

The argument to this statement is a character from ‘1’ through ‘9’. This statement modifies the request comparison method for authentication requests. See Section 5.1 [Extended Comparison], page 67, for a detailed description of its usage.

4.1.4 acct statement

Syntax:

acct {
listen ( addr-list | no );
forward addr-list ;
port number ;
detail bool;
system bool;
max-requests number ;
time-to-live number ;
request-cleanup-delay number ;
compare-atribute-flag character ;
trace-rules bool ;
} ;

Usage:

The acct statement configures the parameters of the accounting service.

listen statement

This statement determines on which addresses radiusd will listen for incoming accounting requests. Its argument is a comma-separated list of items in the form ip:port-number. ip can be either an IP address in familiar “dottedquad” notation or a hostname. :port-number part may be omitted, in which case the default accounting port is assumed.

If the listen statement is omitted, radiusd will accept incoming requests from any interface on the machine.

The special value no disables listening for accounting requests.

The following example configures radius to listen for the incoming requests on the default accounting port on the address 10.10.10.1 and on port 1646 on address 10.10.11.2.

     listen 10.10.10.1, 10.10.11.2:1646;

forward statement

This statement enables forwarding of the requests to the given set of servers. Forwarding is an experimental feature of GNU Radius, it differs from proxying in that the requests are sent to the remote server (or servers) and processed locally. The remote server is not expected to reply.

This mode is intended primarily for debugging purposes. It could also be useful in some very complex and unusual configurations.

Numeric statements

port Sets the number of which port to listen for the authentication requests.

max-requests

Sets the maximum number of accounting requests in the queue. Any surplus requests will be discarded.

time-to-live

Sets the request time-to-live in seconds. The time-to-live is the time to wait for the completion of the request. If the request job isn’t completed within this interval of time it is cleared, the corresponding child process killed and the request removed from the queue.

request-cleanup-delay

Sets the request cleanup delay in seconds, i.e. determines how long will the completed account request reside in the queue.

Boolean statements

detail When set to no, disables detailed accounting (see Section 7.2 [Detailed Request Accounting], page 81). system When set to no, disables system accounting (see Section 7.1

[System Accounting], page 81). Notice, that this will disable simultaneous use checking as well, unless you supply an alternative mlc method (currently sql, See Section 6.9 [Multiple Login Checking], page 74, for the detailed discussion of this).

trace-rules

Enables tracing of the configuration rules that were matched during processing of each received accounting request. See Section 9.1 [Rule Tracing], page 87, for detailed information about this mode.

Character statement

compare-attribute-flag

The argument to this statement is a character from ‘1’ through ‘9’. This statement modifies the request comparison method for authentication requests. See Section 5.1 [Extended Comparison], page 67, for a detailed description of its usage.

4.1.5 usedbm statement

Syntax:

usedbm ( yes | no ) ;

Usage

The usedbm statement determines whether the DBM support should be enabled.

no Do not use DBM support at all.

yes Use only the DBM database and ignore ‘raddb/users’.

4.1.6 snmp statement

Syntax:

snmp {
port portno ;
listen ( addr-list | no );
max-requests number ;
time-to-live number ;
request-cleanup-delay number ;
ident string ;
community name ( rw | ro ) ;
network name network [ network ... ] ;
acl {
        allow network_name community_name ;
        deny network_name ;
} ;

             storage {
                     file filename ;
                     perms number ;
                     max-nas-count number ;
                     max-port-count number ;
             } ;
     };

Usage

The snmp statement configures the SNMP service.

listen statement

The listen statement determines on which addresses radiusd will listen for incoming SNMP requests. The argument is a comma-separated list of items in the form ip:port-number. The ip can be either an IP address in familiar “dotted-quad” notation or a hostname. The :port-number part may be omitted, in which case the default SNMP port (161) is used.

If the listen statement is omitted, radiusd will accept incoming requests from any interface on the machine.

The special value no disables listening for SNMP requests.

The following example configures radius to listen for the incoming SNMP requests on the default SNMP port on the address 10.10.10.1 and on port 4500 on address 10.10.11.2.

     listen 10.10.10.1, 10.10.11.2:4500;

Numeric statements

port Sets the number of which port to listen for the SNMP requests.

max-requests

Sets the maximum number of SNMP requests in the queue. Any surplus requests will be discarded.

time-to-live

Sets the request time-to-live in seconds. The time-to-live is the time to wait for the completion of the request. If the request job isn’t completed within this interval of time it is cleared, the corresponding child process killed and the request removed from the queue.

request-cleanup-delay

Sets the request cleanup delay in seconds, i.e. determines how long will the completed SNMP request reside in the queue.

String statements

ident Sets the SNMP server identification string.

Community and network definitions

community name ( rw | ro )

Defines the community name as read-write (rw) or read-only (ro).

network name network [ network ... ]

Groups several networks or hosts under one logical network name.

Access-Control List definitions

allow network_name community_name

allow hosts from the group network name access to community community name.

deny NETWORK_NAME

Deny access to SNMP service from any host in the group network name.

Storage control

GNU Radius stores the SNMP monitoring data in an area of shared memory mapped to an external file. This allows all subprocesses to share this information and to accumulate the statistics across invocations of the daemon.

The storage statement controls the usage of the storage for the SNMP data.

file Sets the file name for the SNMP storage file. Unless the filename begins with a ‘/’ it is taken as relative to the current logging directory.

perms Sets the access permissions for the storage file. Notice, that this statement does not interpret its argument as octal by default, so be sure to prefix it with ‘0’ to use an octal value.

max-nas-count

Sets maximum number of NASes the storage file is able to handle. Default is 512. Raise this number if you see the following message in your log file:

 reached SNMP storage limit for the number of
 monitored NASes: increase max-nas-count
max-port-count

Sets maximum number of ports the storage file is able to handle. Default is 1024. Raise this number if you see the following message in your log file:

 reached SNMP storage limit for the number of
 monitored ports: increase max-port-count

4.1.7 rewrite statement.

(This message will disappear, once this node revised.)

Syntax:

rewrite {
stack-size number ;
load-path string ;
load string ;
};

Numeric statements

stack-size

Configures runtime stack size for Rewrite. The number is the size of stack in words. The default value is 4096.

String statements

load-path

Add specified pathname to the list of directories searched for rewrite files.

load Loads the specified source file on startup. Unless string is an absolute pathname, it will be searched in directories set up by load-path statement.

Loading

The default load path is `RADDB':`DATADIR'/rewrite.

4.1.8 guile statement

(This message will disappear, once this node revised.) The guile statement allows to configure server interface with Guile.

Syntax

guile {
debug bool ;
load-path string ;
load string ;
load-module string [ string ... ] ;
eval expression [ expression ... ] ;
gc-interval number ;
outfile string ;
};

Usage

Boolean statements

debug When set to yes, enables debugging evaluator and backtraces on Guile scripts.

Numeric statements

gc-interval

Configures the forced garbage collections. By default the invocation of the garbage collector is run by the internal Guile mechanism. However, you may force Radius to trigger the garbage collection at fixed time intervals. The gc-interval statement sets such interval in seconds. For more information about Guile memory management system in general and garbage collections in particular, see section “Memory Management and Garbage Collection” in The Guile Reference Manual.

String statements

eval Evaluates its argument as Scheme expression.

load-path

Adds specified pathname to %load-path variable.

load Loads the specified source file on startup.

load-module

Loads the specified Scheme module on startup. This statement takes an arbitrary number of arguments. The first argument specifies the name of the module to load, the rest of arguments is passed to the module initialization funtion. Module initialization function is a function named ‘module-init’, where module is the module name. Arguments are converted using usual Guile rules, except that the ones starting with a dash (‘-’) are converted to keyword arguments.

outfile Redirects the standard output and standard error streams of the Guile functions to the given file. Unless the filename starts with ‘/’, it is taken relative to the current logging directory.

See Section 10.3 [Guile], page 115, for a detailed description of Guile extensions interface.

4.1.9 message statement

The message statement allows to set up the messages that are returned to the user with authentication-response packets.

Syntax

message {
account-closed string ;
password-expired string ;
password-expire-warning string ;
access-denied string ;
realm-quota string ;
multiple-login string ;
second-login string ;
timespan-violation string ;
};

All variables in message block take a string argument. In string you can use the usual C backslash notation to represent non-printable characters. The use of %C{} and %R{} sequences is also allowed (see Section 4.14 [Macro Substitution], page 64).

String statements

account-closed

This message will be returned to the user whose account is administratively closed.

password-expired

This message will be returned to the user whose password has expired.

password-expire-warning

This is a warning message that will be returned along with an authentication-acknowledge packet for the user whose password will expire in less than n seconds. The value of n is set by

password-expire-warning variable in auth block. See Sec-

tion 4.1.3 [auth], page 28. In this string, you can use the %R{Password-Expire-Days} substitution, to represent the actual number of days left to the expiration date. The default is

     Password Will Expire in %R{Password-Expire-Days} Days\r\n
access-denied

This message is returned to the user who supplies an incorrect password or a not-existent user-name as his authentication credentials.

realm-quota

This message is returned when the user is trying to log in using a realm, and number of users that are currently logged in from this realm reaches maximum value. For a description of realms, see Section 2.4.2.2 [Realms], page 14.

multiple-login

This message is returned to the user, who has logged in more than allowed number of times. For description of how to set the maximum number of concurrent logins, see Section 13.3.25 [Simultaneous-Use], page 192.

second-login

This is a special case of multiple-login, which is used when the user’s login limit is 1.

timespan-violation

This message is returned to the user who is trying to login outside of allowed time interval. For description of how to limit user’s login time, see Section 13.3.14 [Login-Time], page 187.

4.1.10 filters statement

The filters statement configures user-defined external filters. See Section 10.1 [Filters], page 95, for the detailed discussion of external filters.

Syntax

     filters {
             filter ident {
                     exec-path path ;
                     error-log filename ;
                     common bool [max-wait];
                     auth {
                             input-format fmt ;
                             wait-reply bool ;
                     };
                     acct {
                             input-format fmt ;
                             wait-reply bool ;
                     };
             };
             ...
     };

Each filter directive defines a new filter. The ident argument declares the name of the filter. This string must be used in Exec-Program-Wait or Acct-Ext-Program attributes to trigger invocation of this filter (see Section 13.3.7 [Exec-Program-Wait], page 181).

Usage

exec-path path Absolute path to the filter program.

error-log filename Redirect error output from the filter program to filename. If the filename does not start with a slash, it is taken relative to the current logging directory (see Section 4.1.1 [option], page 22).

auth acct These compound statements define authentication and accounting parts of this filter. Any one of them may be missing. The two statements allowed within auth and acct blocks are:

input-format fmt Format of the input line for this filter. Usually this string uses %C{} notations (see Section 4.14 [Macro Substitution], page 64). You can also use the return value from a rewrite function as input line to the filter. To do so, declare:

                         input-format "=my_func()";

where my func is the name of the rewrite function to invoke. The function must return string value.

wait-reply bool If the filter prints a single line of output for each input line, set this to yes. Otherwise, if the filter produces no output, use wait-reply no.

4.1.11 mlc statement

Syntax

mlc {
method (system|sql);
checkrad-assume-logged bool;
};

Usage

Mlc statement configures multiple login checking subsystem (see Section 6.9 [Multiple Login Checking], page 74).

method Sets the method of retrieving information about the currently open sessions. Currently two methods are implemented. Setting method to system will use system accounting database (see Section 7.1 [System Accounting], page 81). This is the default method. Setting it to sql will use sql database.

checkrad-assume-logged radiusd consults the value of this variable when the nas does not responds to checkrad queries (see Section 6.9 [Multiple Login Checking], page 74). If this variable is set to yes, the daemon will proceed as if the nas returned “yes”, i.e. it will assume the user is logged in. Otherwise radiusd assumes the user is not logged in.

4.2 Dictionary of Attributes — ‘raddb/dictionary’

The dictionary file ‘raddb/dictionary’ defines the symbolic names for radius attributes and their values (see Section 2.1 [Attributes], page 7). The file consists of a series of statements, each statement occupies one line.

In the detailed discussion below we use the following meta-syntactic characters:

number Denotes a decimal, octal or hexagesimal number. Usual C conventions are honored, i.e. if number starts with ‘0x’ or ‘0X’ it is read as a hex number, if it starts with ‘0’ it is read as an octal number, otherwise it is read as a decimal one.

type Denotes an attribute type. These are valid attribute types:

string A string type.

integer An integer type.

ipaddr ip address in a dotted-quad form.

date A date in the format: "MON DD CCYY", where MON is the usual three-character abbreviation, DD is day of month (1-31), CCYY is the year, including the century.

4.2.1 Comments

Comments are introduced by a pound sign (‘#’). Everything starting from the first occurrence of ‘#’ up to the end of line is ignored.

4.2.2 $INCLUDE Statement

Syntax

     $INCLUDE `filename'

Usage

The $INCLUDE statement causes the contents of the file ‘filename’ to be read in and processed. The file is looked up in the Radius database directory, unless its name starts with a slash.

4.2.3 VENDOR Statement

Syntax

     VENDOR  vendor-name vendor-id

Usage

A VENDOR statement defines the symbolic name vendor-name for vendor identifier vendor-id. This name can subsequently be used in ATTRIBUTE statements to define Vendor-Specific attribute translations. See Section 13.1.26 [Vendor-Specific], page 175.

Example

     VENDOR  Livingston  307

4.2.4 ATTRIBUTE statement

Syntax

     ATTRIBUTE  name  number  type [vendor] [flags]

Usage

The ATTRIBUTE statement defines the internal representation of an attribute: its symbolic name, data type and syntactical usage. Its parts have the following meaning:

name The attribute name.

number The attribute ID (number).

type The attribute type.

vendor Vendor name for vendor-specific attributes. For usual attributes this field is empty or contains a dash (‘-’). The latter usage is for compatibility with previos version of GNU Radius

flags Flags, defining attribute properties (see Section 2.1 [Attributes], page 7).

The attribute property flags consist of a sequence of letters, whose meaning is determined by the following rules:^{1}

1. The attribute usage is described by three pairs of symbols, enclosed in square brackets. Each pair describes how the attribute can be used in each of three configuration files. The first pair corresponds to ‘raddb/users’, the second one corresponds to ‘raddb/hints’, and the third one corresponds to ‘raddb/huntgroups’. Within each pair, the letter ‘L’ in first position means that the attribute is allowed in lhs of a rule. The letter ‘R’ in second position means that the attribute is allowed in rhs of a rule. The absence of any of these letters is indicated by dash (‘-’). Thus, the following usage specification:

                 [L--RLR]

means that the attribute may be used in lhs of a rule in ‘raddb/users’, in rhs of a rule in ‘raddb/hints’, and in both sides of a rule in

    ‘raddb/huntgroups’.

2. The attribute additivity is described by one of the following letters:

1 The flags are optional for compatibility with previous versions of GNU Radius. If they

are omitted, the default is ‘[LRLRLR]+’

= Additivity = Replace

+ Additivity = Append

N Additivity = None

3. The presence of letter ‘P’ in property flags raises the propagation bit.

4. Letter ‘l’ (lower-case ell) enables logging the given attribute in detail file (see Section 7.2 [Detailed Request Accounting], page 81). This is meaningful only for internal attributes, i.e. the ones whose decimal value is greater than 255 (see Section 13.3 [Radius Internal Attributes], page 178). By default such attributes do not appear in detailed logs. The flag ‘l’ reverts this behavior.

5. Letter ‘E’ marks attributes encrypted as described in RFC 2138. Currently these are User-Password and CHAP-Password.

6. Letter ‘T’ marks attribute encrypted according to RFC 2868.

7. The characters from ‘1’ to ‘9’ denote nine user-defined flags (see Section 5.1 [Extended Comparison], page 67).

Example

     ATTRIBUTE  Service-Type  6 integer - [LR-RLR]=P

This statement declares that the attribute number 6 will be referred to by the symbolic name ‘Service-Type’. The attribute is of integer data type and it may be used in any part of matching rules, except in lhs of a ‘raddb/hints’ rule. The additivity of Service-Type is set to ‘Replace’. The attribute will be propagated through the proxy chain.

4.2.5 Blocks of Vendor-Specific Attributes

Syntax

     BEGIN VENDOR vendor-name [vendor-id]
     ...
     END

Usage

The BEGIN keyword marks start of the block of definitions of vendor-specific attributes. The block is terminated by END keyword, optionally followed by an arbitrary number of words, which are regarded as a comment. The block may contain any valid dictionary declarations, except other blocks: nesting of declaration blocks is not allowed.

If vendor-id is absent, the value of vendor ID is looked up in the internal table of vendors; therefore, it must be defined before BEGIN statement (see Section 4.2.3 [VENDOR], page 40).

BEGIN---END block alters the handling of ATTRIBUTE statements within it. If ATTRIBUTE statement does not contain an explicit vendor-id specification, the value of vendor-id is used instead.

For compatibility with FreeRadius an alternative syntax is also supported:

     BEGIN-VENDOR vendor-name
     ...
     END-VENDOR vendor-name

Such compatibility blocks must appear only ater the declaration of vendorname (see Section 4.2.3 [VENDOR], page 40).

Example

The following is the usual way of definig vendor-specific attributes:

     VENDOR          Livingston      307
     ATTRIBUTE       LE-Terminate-Detail     2       string  Livingston
     ATTRIBUTE       LE-Advice-of-Charge     3       string  Livingston

The following two examples show the alternative ways:

     VENDOR Livingston 307
     BEGIN VENDOR Livingston
     ATTRIBUTE       LE-Terminate-Detail     2       string
     ATTRIBUTE       LE-Advice-of-Charge     3       string
     END
     BEGIN VENDOR Livingston 307
     ATTRIBUTE       LE-Terminate-Detail     2       string
     ATTRIBUTE       LE-Advice-of-Charge     3       string
     END

These three examples are completely equivalent to each other.

4.2.6 ALIAS statement

Syntax

     ALIAS name  alt-name

Usage

The ALIAS statement defines an altenative name alt-name for attribute name. The latter should already be defined, otherwise an error occurs.

Example

     ALIAS User-Password Password

4.2.7 PROPERTY statement

Syntax

     PROPERTY  name  flags
     PROPERTY  name  +flags [-flags ...]

Usage

The PROPERTY statement redefines property flags for attribute name. The attribute must be defined, otherwise an error occurs. The PROPERTY statement has two forms. In first form, it takes a single argument, representing new property flags for the attribute. In its second form it takes any number of arguments, each of them preceeded by ‘+’ sign, inidicating addition of properties, or by ‘-’ sign, indicating removal of these.

See Section 4.2.4 [ATTRIBUTE], page 41, for the discussion of attribute property flags.

Example

The following example defines that the attribute User-Password may be used only on left-hand side of a ‘raddb/users’ entry, and that it is transmitted in encrypted form.

     PROPERTY  User-Password [L-----]E

Next example illustrates adding and removing attribute properties:

     PROPERTY  My-Attrib     +P -=

it adds propagation bit (‘P’) and removes ‘replace’ additivity from My-Attrib attribute.

4.2.8 VALUE Statement

Syntax

     VALUE   Attribute-Translation       Value-Translation       number

Usage

The VALUE statement assigns a translation string to a given value of an integer attribute. Attribute-Translation specifies the attribute and the Value-Translation specifies the name assigned to the value number of this attribute.

Example

The following assigns the translation string ‘Login-User’ to the value 1 of the attribute ‘Service-Type’.

     VALUE  Service-Type  Login-User  1

4.3 Clients List — ‘raddb/clients’

The ‘raddb/clients’ lists nases which are allowed to make authentication requests. As usual, the ‘#’ character introduces a comment. Each record in the file consists of two fields, separated by whitespace. The fields are:

NAS name Specifies a hostname or ip address of the nas.

Key Lists the encryption key shared between the server and this nas.

If the set of nases share the same encryption key, there are two ways to list it in ‘raddb/clients’. First, if these nases lie in a single network, you can specify this network address in NAS name field, e.g.:

     10.10.10.0/27   seCRet

Notice also that specifying full netmask after the ‘/’ character is also allowed, so that the above example could also be written as follows:

     10.10.10.0/255.255.255.224   seCRet

Otherwise, the keyword DEFAULT may be used as NAS name. This notation will match any ip address, so it should be used with caution.

4.3.1 Example of ‘clients’ file

     # This is a list of clients which are allowed to make authentication
     # requests.
     # Each record consists of two fields:
     #       i.  Valid hostname.
     #       ii. The shared encryption key for this hostname.
     #
     #Client Name            Key
     #----------------       -------------------
     myhost.dom.ain          guessme
     merlin                  emrys
     11.10.10.10             secRet
4.4 NAS List — ‘raddb/naslist’

The ‘raddb/naslist’ file contains a list of nases known to the Radius server. Each record in the file consist of the following four fields, the first two being mandatory, the last two being optional:

NAS name Specifies either a hostname or ip address for a single nas or a CIDR net block address for a set of nases. The word ‘DEFAULT’ may be used in this field to match any nas.^{2}

Short Name This field defines a short name under which this nas will be listed in logfiles. The short name is also used as a name of the subdirectory where the detailed logs are stored.

Type Specifies the type of this nas. Using this value radiusd determines the way to query nas about the presence of a given user on it (see Section 6.9 [Multiple Login Checking], page 74). The two special types: ‘true’ and ‘false’, can be used to disable nas querying. When the type field contains ‘true’, radiusd assumes the user is logged in to the nas, when it contains ‘false’, radiusd assumes the user is not logged in. Otherwise, the type is used as a link to ‘nastypes’ entry (see Section 4.5 [nastypes file], page 47). If this field is not present ‘true’ is assumed.

2 Logins from DEFAULT nases are not reflected in SNMP variables.

Arguments Additional arguments describing the nas. Multiple arguments must be separated by commas. No intervening whitespace is allowed in this field.

There are two groups of nas arguments: nas-specific arguments and nasquerying arguments. Nas-specific arguments are used to modify a behavior of radiusd when sending or receiving the information to or from a particular nas.

Nas-querying arguments control the way radiusd queries a nas for confirmation of a user’s session (see Section 6.9 [Multiple Login Checking], page 74). These arguments override the ones specified in ‘nastypes’ and can thus be used to override the default values.

The nas-specific arguments currently implemented are: broken pass

This is a boolean argument that controls the encryption of user passwords, longer than 16 octets. By default, radiusd uses method specified by rfc 2865. However some nases, most notably max ascend series, implement a broken method of encoding long passwords. This flag instructs radiusd to use broken method of password encryption for the given nas.

compare-auth-flag=flag Instructs radius to use attributes marked with a given user-defined flag when comparing authentication requests. It overrides compare-attribute-flag (see Section 4.1.3 [auth], page 28) for this particular nas. See Section 5.1 [Extended Comparison], page 67, for a detailed description of its usage.

compare-acct-flag=flag Instructs radius to use attributes marked with a given userdefined flag when comparing accounting requests. It overrides compare-attribute-flag (see Section 4.1.4 [acct], page 30) for this particular nas. See Section 5.1 [Extended Comparison], page 67, for a detailed description of its usage.

See Section 2.4.1 [Checking Duplicates], page 12, for general description of request comparison methods.

For the list of nas-querying arguments, See Section 4.5 [Full list of allowed arguments], page 47.

4.4.1 Example of ‘naslist’ file

# raddb/naslist: contains a list of Network Access Servers

     #
     # Each record consists of following fields:
     #
     #       i.      A valid hostname or IP address for the client.
     #       ii.     The short name to use in the logfiles for this NAS.
     #       iii.    Type of device. Valid values are `true', `false' and
     #               those defined in raddb/nastypes file.
     # NAS Name              Short Name      Type
     #----------------       ----------      ----
     myhost.dom.ain          myhost          unix
     merlin                  merlin          max
     11.10.10.10             arthur          livingston

4.5 NAS Types — ‘raddb/nastypes’

The ‘raddb/nastypes’ file describes the ways to query nases about active user sessions.

4.5.1 Syntax of ‘raddb/nastypes’

(This message will disappear, once this node revised.)

Syntax

Each record consists of three fields separated by any amount of whitespace. The fields are:

Type Type of the nas which is described in this record.

Method Method to use to query a nas of given type.

Arguments Arguments to pass to this method. Each argument is a pair arg=value, where arg is its name and value is a value assigned to it. The list of predefined argument names follows. Note, that no intervening whitespace is allowed in this field.

Methods

Version 1.3 of GNU Radius supports following querying methods: finger, snmp, external and guile. .

Arguments

In the discussion below n means numeric and s string value.

The following arguments are predefined:

Common for all methods

function=s Specifies the check function to use with this method (see Section 10.2.5 [Login Verification Functions], page 103). This argument must be present. For description of how this function is applied, see Section 6.9 [Multiple Login Checking], page 74.

port=n Use port number n instead of the default for the given method.

Method snmp

password=s Use community s instead of the default. This argument must be present.

retries=n Retry n times before giving up.

timeout=n Timeout n seconds on each retry.

Method finger

timeout=n Give up if the nas does not respond within n seconds.

notcp tcp=0 Disable the use of T/TCP for hosts with a broken TCP imple-

mentation.

arg=subst Send subst to finger, instead of username. subst must be one of macro variables, described below.

Macro variables

The following macro-variables are recognized and substituted when encountered in the value pair of an argument:

‘%u’ Expands to username.

‘%s’ Expands to session id.

‘%d’ Expands to session id converted to decimal representation.

‘%p’ Expands to port number.

‘%P’ Expands to port number + 1.

4.5.2 Example of nastypes file.

Note, that in the following example the long lines are broken into several lines for readability.

     # Type     Method          Args
     # ----     ------          ----

unix       finger       function=check_unix
max-f      finger       function=check_max_finger
max        snmp         oid=.1.3.6.1.4.1.529.12.3.1.4.%d,
                function=check_snmp_u
as5300-f   finger       function=check_as5300_finger
as5300     snmp         oid=.1.3.6.1.4.1.9.9.150.1.1.3.1.2.%d,
                function=check_snmp_u
livingston snmp         oid=.1.3.6.1.4.1.307.3.2.1.1.1.5.%P,
                function=check_snmp_s

4.5.3 Standard nas types

The ‘nastypes’ shipped with version 1.3 of GNU Radius defines following nas types:

unix — UNIX boxes running Finger This type suits for unix boxes running finger service able to return information about dial-up users active on them. To enable finger checking of a unix host add following to your ‘naslist’ file:

     #Hostname       Shortname   Type
     #--------       ---------   ----
     nas.name        T           unix

max-f — MAX Ascend with Finger Use this type if you have MAX Ascend terminal server that answers finger queries. The ‘naslist’ entry for such nas will look like:

     #Hostname       Shortname   Type  Flags
     #--------       ---------   ----  -----
     nas.name        T           max-f broken_pass

Note the use of broken_pass flag. It is needed for most MAX Ascend servers (see Section 4.4 [naslist file], page 45).

max — MAX Ascend, answering SNMP Use this type if you have MAX Ascend terminal server that answers SNMP queries. The ‘naslist’ entry for such nas will look like:

     #Hostname       Shortname   Type  Flags
     #--------       ---------   ----  -----
     nas.name        T           max-f broken_pass,community=comm

Replace comm with your actual SNMP community name.

as5300-f — Cisco AS5300 running finger as5300 — Cisco AS5300 answering SNMP livingston — Livingston Portmaster

Type livingston queries portmaster using SNMP.

4.6 Request Processing Hints — ‘raddb/hints’

The ‘raddb/hints’ file is used to modify the contents of the incoming request depending on the username. For a detailed description of this, See Section 2.4.3 [Hints], page 14.

The file contains data in Matching Rule format (see Section 2.3 [Matching Rule], page 11).

Notice, that versions of GNU Radius up to 1.0 allowed to use only a subset of attributes in the check list of a ‘hints’ entry, namely:

  • Suffix
  • Prefix
  • Group
  • User-ID

This requirement has been removed in version 1.0.

4.6.1 Example of ‘hints’ file

     ## If the username starts with `U', append the UUCP hint
     DEFAULT         Prefix = "U", Strip-User-Name = No
                     Hint = "UUCP"
     ## If the username ends with `.slip', append the SLIP service data
     ## and remove the suffix from the user name.
     DEFAULT         Suffix = ".slip",
                        Strip-User-Name = Yes
                     Hint = "SLIP",
                        Service-Type = Framed-User,
                        Framed-Protocol = SLIP

4.7 Huntgroups — ‘raddb/huntgroups’

The ‘raddb/huntgroups’ contains the definitions of the huntgroups. For a detailed description of huntgroup concept, See Section 2.4.4 [Huntgroups], page 15.

The file contains data in Matching Rule format (see Section 2.3 [Matching Rule], page 11).

4.7.1 Example of ‘huntgroups’ file.

     ## This defines the packet rewriting function for the server 11.10.10.11
     DEFAULT NAS-IP-Address = 11.10.10.11, Rewrite-Function = "max_fixup"
             NULL

4.8 List of Proxy Realms — ‘raddb/realms’

The ‘raddb/realms’ file lists remote Radius servers that are allowed to communicate with the local Radius server (see Section 2.4.2 [Proxying], page 13).

Each record consists of up to three fields, separated by whitespace. Two of them are mandatory. The fields are:

Realm name Specifies the name of the realm being defined, i.e. part of the login name after the ‘@’ symbol. There are three special forms of this field. The name ‘NOREALM’ defines the empty realm, i.e. lines marked with this name will match user names without any realm suffix. The name ‘DEFAULT’ defines the default realm (see Section 2.4.2.2 [Realms], page 14). The lines with this realm name will match any user name, not matched by any other line in

‘raddb/realms’.

Remote server list A comma-separated list of remote servers to which the requests for this realm should be forwarded. Each item in the list is:

     servername[:auth-port[:acct-port]]

Optional auth-port and acct-port are the authentication and accounting port numbers. If acct-port is omitted, it is computed as auth-port + 1. If auth-port is omitted, the default authentication port number is used. The servers from this list are tried in turn until any of them replies or the list is exhausted, whichever occurs first. The timeout value and number of retries for each server are set via timeout and retry flags (see below). There may be cases where you would wish a particular realm to be served by the server itself. It is tempting to write

     # Wrong!
     realm.name      localhost

however, this will not work. The special form of the server list is provided for this case. It is the word ‘LOCAL’. The correct configuration line for the above case will thus be:

     # Use this to declare a locally handled realm
     realm.nam       LOCAL

Flags (optional)

The flags meaningful in ‘raddb/realms’ are

ignorecase Boolean value. When set, enables case-insensitive comparison

of realm names. For example, if a realm were defined as

     myrealm.net     remote.server.net:1812  ignorecase

then user name ‘user@MyREAlm.NeT’ will match this definition.

52 GNU Radius Reference Manual

strip Boolean value. Controls whether the realm name should be stripped off the username before forwarding the request to the remote server. Setting strip enables stripping, setting nostrip disables it. Default is to always strip user names. quota=num

Set maximum number of concurrent logins allowed from this realm to the given value (num).

timeout Number of seconds to wait for reply from the remote server before retransmitting the request.

retries Number of attempts to connect a server. If the server does not respond after the last attempt, the next server from the list is tried.

auth Proxy only authentication requests.

acct Proxy only accounting requests.

4.8.1 Example of ‘realms’ file

Example 1.

# Realm                 Remote server[:port]            flags
#----------------       ---------------------           --------
that.net                radius.that.net                 nostrip
dom.ain                 server.dom.ain:3000             strip,quota=20
remote.net              srv1.remote.net,srv2.remote.net

Example 2.

# Realm                 Remote server[:port]            flags
#----------------       ---------------------           --------
NOREALM                 radius.server.net
that.net                radius.that.net                 nostrip
dom.ain                 server.dom.ain:3000             strip,quota=20

4.9 User Profiles — ‘raddb/users’

File ‘raddb/users’ contains the list of User Profiles. See Section 2.4.5 [User Profiles], page 16, for a description of its purpose.

4.9.1 Example of ‘users’ file

## The following entry is matched when the user appends ``.ppp'' to his
## username when logging in.
## The suffix is removed from the user name, then the password is
## looked up in the SQL database.
## Users may log in at any time. They get PPP service.
DEFAULT Suffix = ".ppp",
                Auth-Type = SQL,
                Login-Time = "Al",
                Simultaneous-Use = 1,
                Strip-User-Name = Yes
        Service-Type = Framed-User,
                Framed-Protocol = PPP

Chapter 4: Radius Configuration Files 53

## This is for SLIP users.
## This entry is matched when the auth request matches ``SLIP'' hint
DEFAULT Hint = "SLIP",
                Auth-Type = Mysql
        Service-Type = Framed-User
                Framed-Protocol = SLIP
## The following authenticates users using system passwd files.
## The users are allowed to log in from 7:55 to 23:05 on any weekday,
## except the weekend, and from 07:55 to 12:00 on Sunday.
## Only one login is allowed per user.
## The program telauth is used to further check the authentication
## information and provide the reply pairs
## Note the use of backslashes to split a long line.
DEFAULT Auth-Type = System,
                Login-Time = "Wk0755-2305,Su0755-1200",
                Simultaneous-Use = 1
        Exec-Program-Wait = "/usr/local/sbin/telauth \
                             %C{User-Name} \
                             %C{Calling-Station-Id} \
                             %C{NAS-IP-Address} \
                             %C{NAS-Port-Id}"
## This particular user is authenticated via PAM. He is presented a
## choice from `raddb/menus/menu1' file.
gray    Auth-Type = Pam
        Menu = menu1

4.10 List of Blocked Users — ‘raddb/access.deny’

The ‘raddb/access.deny’ file contains a list of user names which are not allowed to log in via Radius. Each user name is listed on a separate line. As usual, the ‘#’ character introduces an end-of-line comment.

4.11 SQL Configuration — ‘raddb/sqlserver’

The ‘raddb/sqlserver’ file configures the connection to sql server.

The file uses simple line-oriented ‘keyword --- value’ format. Comments are introduced by ‘#’ character.

The ‘sqlserver’ statements can logically be subdivided into following groups: SQL Client Parameters, configuring the connection between sql client and the server, Authentication Server Parameters, Authorization Parameters, and Accounting server parameters.

4.11.1 SQL Client Parameters

These parameters configure various aspects of connection between sql client and the server.

interface iface-type

Specifies the sql interface to use. Currently supported values for iface-type are mysql and postgres. Depending on this, the default communication port number is set: it is 3306 for interface mysql and 5432 for interface postgres. Use of this statement is only meaningful when the package was configured with both

‘--with-mysql’ and ‘--with-postgres’ option.
server string

Specifies the hostname or ip address of the sql server.

port number

Sets the sql communication port number. It can be omitted if your server uses the default port.

login string

Sets the sql user login name.

password password

Sets the sql user password.

keepopen bool

Specify whether radiusd should try to keep the connection open. When set to no (the default), radiusd will open new connection before the transaction and close it right after finishing it. We recommend setting keepopen to yes for heavily loaded servers, since opening the new connection can take a substantial amount of time and slow down the operation considerably.

idle_timeout number

Set idle timeout in seconds for an open sql connection. The connection is closed if it remains inactive longer that this amount of time.

4.11.2 Authentication Server Parameters

(This message will disappear, once this node revised.) These parameters configure the sql authentication. The general syntax

is:

doauth bool

When set to yes, enables authentication via sql. All auth_ keywords are ignored if doauth is set to no.

auth_db string

Specifies the name of the database containing authentication information.

auth_query string

Specifies the sql query to be used to obtain user’s password from the database. The query should return exactly one string value — the password.

group_query string

Specifies the query that retrieves the list of user groups the user belongs to. This query is used when Group or Group-Name attribute appears in the lhs of a user’s or hint’s profile.

auth_success_query string

This query is executed when an authentication succeeds. See Section 6.10 [Auth Probing], page 76, for the detailed discussion of its purpose.

auth_failure_query string

This query is executed upon an authentication failure. See Section 6.10 [Auth Probing], page 76, for the detailed discussion of its purpose.

Example of Authentication Server Parameters

Let’s suppose the authentication information is kept in the tables passwd

and groups.

The passwd table contains user passwords. A user is allowed to have different passwords for different services. The table structure is:

     CREATE TABLE passwd (
       user_name           varchar(32) binary default '' not null,
       service             char(16) default 'Framed-PPP' not null,
       password            char(64)
     );

Additionally, the table groups contains information about user groups a particular user belongs to. Its structure is:

     CREATE TABLE groups (
       user_name           char(32) binary default '' not null,
       user_group          char(32)
     );

The queries used to retrieve the information from these tables will then look like:

     auth_query  SELECT password
                 FROM passwd
                 WHERE user_name = '%C{User-Name}'
                 AND service = '%C{Auth-Data}'
     group_query SELECT user_group
                 FROM groups
                 WHERE user_name = '%C{User-Name}'

It is supposed, that the information about the particular service a user is wishing to obtain, will be kept in Auth-Data attribute in lhs of a user’s profile.

4.11.3 Authorization Parameters

These parameters define queries used to retrieve the authorization information from the sql database. All the queries refer to the authentication database.

check_attr_query string

This query must return a list of triplets:

                 attr-name, attr-value, opcode

The query is executed before comparing the request with the profile entry. The values returned by the query are added to lhs of the entry. opcode here means one of valid operation codes: ‘=’, ‘!=’, ‘<’, ‘>’, ‘<=’, ‘>=’.

reply_attr_query string

This query must return pairs:

                 attr-name, attr-value

The query is executed after a successful match, the values it returns are added to the rhs list of the matched entry, and are therefore returned to the nas in the reply packet.

Example of Authorization Parameters

Suppose your attribute information is stored in a sql table of the following structure:

     CREATE TABLE attrib (
       user_name varchar(32) default '' not null,
       attr      char(32) default '' not null,
       value     char(128),
       op enum("=", "!=", "<", ">", "<=", ">=") default null
     );

Each row of the table contains the attribute-value pair for a given user. If op field is NULL, the row describes rhs (reply) pair. Otherwise, it describes a lhs (check) pair. The authorization queries for this table will look as follows:

     check_attr_query  SELECT attr,value,op \
                       FROM attrib \
                       WHERE user_name='%u' \
                       AND op IS NOT NULL
     reply_attr_query  SELECT attr,value \
                       FROM attrib \
                       WHERE user_name='%u' \
                       AND op IS NULL

Now, let’s suppose the ‘raddb/users’ contains only one entry:

     DEFAULT Auth-Type = SQL
             Service-Type = Framed-User

And the attrib table contains following rows:

user name  attr  value  op

jsmith         NAS-IP-Address     10.10.10.1      =

jsmith         NAS-Port-Id        20              <=
jsmith         Framed-Protocol    PPP             NULL
jsmith         Framed-IP- Address  10.10.10.11     NULL

Then, when the user jsmith is trying to authenticate, the following happens:

1. Radius finds the matching entry (DEFAULT) in the ‘raddb/users’.

2. It queries the database using the check_attr_query. The triplets it returns are then added to the lhs of the profile entry. Thus, the lhs will contain:

         Auth-Type = SQL,
         NAS-IP-Address = 10.10.10.1,
         NAS-Port-Id <= 20

3. Radius compares the incoming request with the lhs pairs thus obtained. If the comparison fails, it rejects the authentication. Note that the Auth-Type attributes itself triggers execution of auth_query, described in the previous section.

4. After a successful authentication, Radius queries the database, using reply_attr_query, and adds its return to the list of rhs pairs. The rhs pairs will then be:

         Service-Type = Framed-User,
         Framed-Protocol = PPP,
         Framed-IP-Address = 10.10.10.11

This list is returned to the nas along with the authentication accept packet.

Thus, this configuration allows the user jsmith to use only nas 10.10.10.1, ports from 1 to 20 inclusive. If the user meets these conditions, he is allowed to use ppp service, and is assigned ip address 10.10.10.11.

4.11.4 Accounting Parameters

To perform the sql accounting radiusd needs to know the database where it is to store the accounting information. This information is supplied by the following statements:

doacct bool

When set to yes enables sql accounting. All acct_keywords are ignored if doacct is set to no.

acct_db string

Specifies the name of the database where the accounting information is to be stored.

Further, radiusd needs to know which information it is to store into the database and when. Each of five accounting request types (see Section 2.2.2 [Accounting Requests], page 9) has a sql query associated with it. Thus, when radius receives an accounting request, it determines the query to use by the value of Acct-Status-Type attribute.

Following statements define the accounting queries:

acct_start_query string

Specifies the sql query to be used when Session Start Packet is received. Typically, this would be some INSERT statement (see Section 4.11.4.1 [Queries], page 58).

acct_stop_query string

Specifies the sql query to be used when Session Stop Packet is received. Typically, this would be some UPDATE statement.

acct_stop_query string

Specifies the sql query to be executed upon arrival of a Keepalive Packet. Typically, this would be some UPDATE statement.

acct_nasup_query string

Specifies the sql query to be used upon arrival of an Accounting Off Packet.

acct_nasdown_query string

Specifies the sql query to be used when a nas sends Accounting On Packet.

None of these queries should return any values.

Three queries are designed for use by multiple login checking mechanism (see Section 6.9 [Multiple Login Checking], page 74):

mlc_user_query string

A query retrieving a list of sessions currently opened by the given user.

mlc_realm_query string

A query to retrieve a list of sessions currently open for the given realm.

mlc_stop_query string

A query to mark given record as hung.

4.11.4.1 Writing SQL Accounting Query Templates

Let’s suppose you have an accounting table of the following structure:

Chapter 4: Radius Configuration Files 59

CREATE TABLE calls (
status              int(3),
user_name           char(32),
event_date_time     datetime DEFAULT '0000-00-00 00:00:00' NOT NULL,
nas_ip_address      char(17),
nas_port_id         int(6),
acct_session_id     char(16) DEFAULT '' NOT NULL,
acct_session_time   int(11),
acct_input_octets   int(11),
acct_output_octets  int(11),
connect_term_reason int(4),
framed_ip_address   char(17),
called_station_id   char(32),
calling_station_id  char(32)
);

On receiving the Session Start Packet we would insert a record into this table with status set to 1. At this point the columns acct_session_

time, acct_input_octets, acct_output_octets as well as connect_term_

reason are unknown, so we will set them to 0:

# Query to be used on session start
acct_start_query     INSERT INTO calls \
                   VALUES(%C{Acct-Status-Type},\
                          '%u',\
                          '%G',\
                          '%C{NAS-IP-Address}',\
                          %C{NAS-Port-Id},\
                          '%C{Acct-Session-Id}',\
                          0,\
                          0,\
                          0,\
                          0,\
                          '%C{Framed-IP-Address}',\
                          '%C{Called-Station-Id}',\
                          '%C{Calling-Station-Id}')

Then, when the Session Stop Packet request arrives we will look up the record having status = 1, user_name matching the value of User-Name

attribute, and acct_session_id matching that of Acct-Session-Id at-

tribute. Once the record is found, we will update it, setting

status = 2
acct_session_time = value of Acct-Session-Time attribute
acct_input_octets = value of Acct-Input-Octets attribute
acct_output_octets = value of Acct-Output-Octets attribute
connect_term_reason = value of Acct-Terminate-Cause attribute

Thus, every record with status = 1 will represent the active session and every record with status = 2 will represent the finished and correctly closed record. The constructed acct_stop_query is then:

     # Query to be used on session end
     acct_stop_query      UPDATE calls \
                          SET status=%C{Acct-Status-Type},\
                              acct_session_time=%C{Acct-Session-Time},\
                              acct_input_octets=%C{Acct-Input-Octets},\
                              acct_output_octets=%C{Acct-Output-Octets},\
                              connect_term_reason=%C{Acct-Terminate-Cause} \
                          WHERE user_name='%C{User-Name}' \
                          AND status = 1 \
                          AND acct_session_id='%C{Acct-Session-Id}'

Upon receiving a Keepalive Packet we will update the information stored

with acct_start_query:
     acct_alive_query  UPDATE calls \
                       SET acct_session_time=%C{Acct-Session-Time},\
                           acct_input_octets=%C{Acct-Input-Octets},\
                           acct_output_octets=%C{Acct-Output-Octets},\
                           framed_ip_address=%C{Framed-IP-Address} \
                       WHERE user_name='%C{User-Name}' \
                       AND status = 1 \
                       AND acct_session_id='%C{Acct-Session-Id}'

Further, there may be times when it is necessary to bring some nas down. To correctly close the currently active sessions on this nas we will define a acct_nasdown_query so that it would set status column to 2 and update

acct_session_time in all records having status = 1 and nas_ip_address

equal to ip address of the nas. Thus, all sessions on a given nas will be closed correctly when it brought down. The acct_session_time can be computed as difference between the current time and the time stored in

event_date_time column:
     # Query to be used when a NAS goes down, i.e. when it sends
     # Accounting-Off packet
     acct_nasdown_query UPDATE calls \
                        SET status=2,\
                            acct_session_time=unix_timestamp(now())-\
                                    unix_timestamp(event_date_time) \
                        WHERE status=1 \
                        AND nas_ip_address='%C{NAS-IP-Address}'

We have not covered only one case: when a nas crashes, e.g. due to a power failure. In this case it does not have a time to send Accounting-Off request and all its records remain open. But when the power supply is restored, the nas will send an Accounting On packet, so we define a acct_

nasup_query to set status column to 3 and update acct_session_time in

all open records belonging to this nas. Thus we will know that each record having status = 3 represents a crashed session. The query constructed will be:

     # Query to be used when a NAS goes up, i.e. when it sends
     # Accounting-On packet
     acct_nasup_query   UPDATE calls \
                        SET status=3,\
                            acct_session_time=unix_timestamp(now())-\
                                    unix_timestamp(event_date_time) \
                        WHERE status=1 \
                        AND nas_ip_address='%C{NAS-IP-Address}'

If you plan to use SQL database for multiple login checking (see Section 6.9 [Multiple Login Checking], page 74), you will have to supply at least two additional queries for retrieving the information about currently active sessions for a given user and realm (see Section 6.9.1 [Retrieving Session Data], page 74). Each of these queries must return a list consisting of 5-element tuples:

     user-name, nas-ip-address, nas-port-id, acct-session-id

For example, in our setup these queries will be:

     mlc_user_query SELECT user_name,nas_ip_address,\
                           nas_port_id,acct_session_id \
                    FROM calls \
                    WHERE user_name='%C{User-Name}' \
                    AND status = 1
     mlc_realm_query SELECT user_name,nas_ip_address,\
                            nas_port_id,acct_session_id \
                     FROM calls \
                     WHERE realm_name='%C{Realm-Name}'

While performing multiple login checking radiusd will eventually need to close hung records, i.e. such records that are marked as open in the database (status=1, in our setup), but are actually not active (See Section 6.9.2 [Verifying Active Sessions], page 75, for the description of why it may be necessary). It will by default use acct_stop_query for that, but it has a drawback that hung records will be marked as if they were closed correctly. This may not be suitable for accounting purposes. The special query mlc_ stop_query is provided to override acct_stop_query. If we mark hung records with status=4, then the mlc_stop_query will look as follows:

     mlc_stop_query UPDATE calls \
                    SET status=4,\
                     acct_session_time=unix_timestamp(now())-\
                                       unix_timestamp(event_date_time) \
                    WHERE user_name='%C{User-Name}' \
                      AND status = 1 \
                      AND acct_session_id='%C{Acct-Session-Id}'

4.12 Rewrite functions — ‘raddb/rewrite’

The file ‘raddb/rewrite’ contains definitions of Rewrite extension functions. For information regarding Rewrite extension language See Section 10.2 [Rewrite], page 98.

4.13 Login Menus — ‘raddb/menus’

The menus is a way to allow user the choice between various services he could be provided. The menu functionality is enabled when Radius is compiled

with ‘--enable-livingston-menus’ option.

A user is presented a menu after it is authenticated if the rhs of his profile record consists of a single a/v pair in the form:

     Menu = <menu-name>

The menu files are stored in directory ‘raddb/menus’.

4.13.1 A menu file syntax.

A menu file is a text file containing a menu declaration and any number of choice descriptions. The menus can be nested to an arbitrary depth.

A comment is introduced by a ‘#’ character. All characters from this one up to the end of line are discarded.

The menu declaration is contained between the words ‘menu’ and ‘end’. Each of these must be the only word on a line and must start in column 1.

Choice descriptions follow the menu declaration. Each description starts with a line containing choice identifier. A choice identifier is an arbitrary word identifying this choice, or a word ‘DEFAULT’. It is followed by commaseparated list of a/v pairs which will be returned to the server when a user selects this choice.

4.13.2 An example of menu files

Single-Level Menu

Suppose the following file is stored under ‘raddb/menus/menu1’:

     menu
             *** Welcome EEE user! ***
     Please select an option:
             1. Start CSLIP session
             2. Start PPP session
             3. Quit
             Option:
     end
     # CSLIP choice
     # Framed-IP-Address of 255.255.255.254 indicates that the NAS should
     # select an address for the user from its own IP pool.
     1
             Service-Type = Framed-User,
             Framed-Protocol = SLIP,
             Framed-IP-Address = 255.255.255.254,
             Termination-Menu = "menu1"
     # PPP choice

     2
             Service-Type = Framed-User,
             Framed-Protocol = PPP,
             Framed-IP-Address = 255.255.255.254,
             Termination-Menu = "menu1"
     # A special menu EXIT means abort the session
     3
             Menu = "EXIT"
     # Return to this menu if no valid choice have been entered
     DEFAULT
             Menu = "menu1"

Now, suppose the ‘raddb/users’ contains the following profile entry:

     DEFAULT Auth-Type = System
             Menu = "menu1"

and user ‘jsmith’ has a valid system account and tries to log in from some nas. Upon authenticating the user, the Radius server sees that his reply pairs contain the Menu attribute. Radius then sends Access-Challenge packet to the nas with the text of the menu in it. The ‘jsmith’ then sees on his terminal:

             *** Welcome EEE user! ***
     Please select an option:
             1. Start CSLIP session
             2. Start PPP session
             3. Quit
             Option:

He then enters ‘2’. The nas sends the Access-Request packet to the server, which sees that user wishes to use option 2 and replies to the nas with an Access-Accept packet containing the following attributes:

             Service-Type = Framed-User,
             Framed-Protocol = PPP,
             Framed-IP-Address = 255.255.255.254,
             Termination-Menu = "menu1"

The Termination-Menu in this list makes sure the same process will continue when ‘jsmith’ logs out, i.e. he will be presented the same menu again until he enters choice ‘3’ which will disconnect him.

Nested menus

In this example, the ‘other’ choice refers to the menu above.

     menu
             *** Welcome here! ***
     Please enter an option:
             ppp     ---     Start PPP session
             telnet  ---     Begin guest login session
             other   ---     Select other option
             Enter your choice:

     end
     ppp
             Service-Type = Framed-User,
             Framed-Protocol = PPP
     telnet
             Service-Type = Login-User,
             Login-IP-Host = 10.11.11.7,
             Login-Service = Telnet,
             Login-TCP-Port = 23
     other
             Menu = "menu1"
     DEFAULT
             menu = "menu2"

4.14 Macro Substitution

Some statements in the configuration files need to use the actual values of the attributes supplied with the request. These are:

  • Exec-Program and Exec-Program-Wait assignments in ‘users’ data-

base

• sql query templates in ‘sqlserver’

In these statements the following macros are replaced by the value of corresponding attributes:

%Cnum (num is a decimal number). This variable is replaced by the value of attribute number ‘num’. The attribute is looked up in the incoming request pairlist.

%C{attr-name}

This is replaced by the value of attribute named ‘attr-name’. The attribute is looked up in the incoming request pairlist.

%Rnum (num is a decimal number). This variable is replaced by the value of attribute number ‘num’. The attribute is looked up in the reply pairlist.

%R{attr-name}

This is replaced by the value of attribute named ‘attr-name’. The attribute is looked up in the reply pairlist.

%D This is replaced by current date/time (localtime).

%G This is replaced by current date/time in GMT.

The exact substitution procedure varies depending on the type of the attribute referenced by macro. If the attribute is of string or date type, radiusd first checks if the resulting substitution should be quoted. It does so by looking at the character immediately preceeding ‘%’. If it is a single or double quote, then radiusd assumes the macro must be quoted and replaces it by an appropriately modified attribute value. The purpose of the modification is to ensure that no characters within the expanded string would conflict with the quoting characters. In particular, radiusd searches the attribute value for any of the characters ‘\’, ‘'’, ‘"’ and prepends a ‘\’ to any occurrence of these. For example, suppose that attribute NAS-Identifier has the value ‘A "new" host’. Then:

     nasid=%C{NAS-Identifier} ↦ nasid=A "new" host
     nasid="%C{NAS-Identifier}" ↦ nasid="A \"new\" host"
     nasid=%\C{NAS-Identifier} ↦ nasid=A \"new\" host

The last example illustrates the use of backslash character to force string quoting in the absense of explicit quotation marks.

If an integer attribute reference is quoted, radiusd looks up the string translation of its value in the dictionary (see Section 4.2.8 [VALUE], page 44) and uses this string as a replacement. If no translation is found, the numeric value is used. The following example assumes that the value of Acct-

Terminate-Cause attribute is 10:
     reason=%C{Acct-Terminate-Cause} ↦ reason=10
     reason='%C{Acct-Terminate-Cause}' ↦ reason='NAS-Request'
     reason=%\C{Acct-Terminate-Cause} ↦ reason=NAS-Request

Again, a backslash after percent sign can be used to force dictionary lookup.

The “‘{}’ form” allows to specify default value for the substitution. The default value will be used when no such attribute is encountered in the pairlist. The syntax for specifying the default value resembles that of shell environment variables.

The substitution %C{attr-name:-defval} is expanded to the value of attr-name attribute, if it is present in the request pairlist, and to defval otherwise. For example:

             %C{Acct-Session-Time:-0}

will return the value of Acct-Session-Time attribute or 0 if it doesn’t exist in the request pairlist.

The substitution %C{attr-name:=defval} is expanded to the value of attr-name attribute. If this attribute is not present in the request pairlist, it will be created and assigned the value defval. E.g.:

             %C{Acct-Session-Time:=0}

The substitution %C{attr-name:?message} is expanded to the value of attr-name attribute, if it is present. Otherwise the diagnostic message “attrname: message” is issued to the log error channel, and string “message” is returned.

The substitution %C{attr-name:+retval} is expanded to empty string if the attribute attr-name is present in the referenced pairlist. Otherwise it is expanded to retval.

You can also use the following shortcuts:

%p Port number

%n nas ip address %f Framed ip address %u User name %c Callback-Number %i Calling-Station-Id %t MTU %a Protocol (SLIP/PPP) %s Speed (Connect-Info attribute)

5 Request Comparison Methods

The basic notions about comparison of the incoming requests and why it is necessary were given in Section 2.4.1 [Checking Duplicates], page 12. This chapter concentrates on extended methods of request comparison and on the configuration issues.

5.1 Extended Comparison

The default comparison method may fail to recognize duplicate requests. if the originating nas has modified the request authenticator or request identifier before retransmitting the request. If you happen to use such nases, you will have to enable extended request comparison to compensate for their deficiencies.

The extended request comparison consists in comparing the contents of both requests. However, blindly comparing each a/v pair from both requests won’t work, since many attributes do change their values between successive retransmits. Therefore, radiusd uses only comparable attribute, i.e. a userdefined subset of such attributes that can safely be used in comparison. Thus, extended request comparison works as follows:

1. The comparable attributes are extracted from each request. They form two sorted attribute lists.

2. If lengths of both lists differ, the requests are considered different.

3. Otherwise, the value of each a/v pair from the first list is compared against that of the corresponding a/v pair from the second list. If at least one a/v pair differs, then the requests are considered different. Notice, that values of Password and CHAP-Password are decoded prior to comparison.

To use the extended comparison, follow the procedure below:

1. Select user-defined attribute properties.

The syntax of dictionary file allows for nine user-defined properties, denoted by characters ‘1’ through ‘9’. You should select one of them to mark comparable attributes for authentication and another one to mark those for accounting. It is strongly suggested that you use PROPERTY statement in your main dictionary file (see Section 4.2.7 [PROPERTY], page 43), instead of modifying ATTRIBUTE statements in the underlying dictionary files. See Section 4.2.4 [ATTRIBUTE], page 41, for detailed description of attribute property flags.

2. To enable the extended comparison for requests coming from any nas, declare extended comparison flags in ‘raddb/config’. To enable the extended comparison for authentication requests, add to your auth block the statement

             compare-attribute-flag flag;

The flag is the same symbol you used in the dictionary to mark comparable attributes for authentication. To enable the extended comparison for accounting requests, insert

compare-attribute-flag statement into the acct block.

3. To enable the extended comparison for requests coming from selected nases, declare extended comparison flags in ‘raddb/naslist’. Add the following statement to the declaration of those nases, that require using the extended comparison (in flags column):

             compare-auth-flag=flag,compare-acct-flag=flag

See Section 4.4 [naslist file], page 45, for a description of naslist file syntax.

5.1.1 An example of extended comparison configuration

In this example configuration, the user-defined flag ‘1’ marks authentication comparable attributes, and the flag ‘2’ marks the accounting comparable attributes.

‘raddb/dictionary’
PROPERTY       User-Name               +12
PROPERTY       Password                +1
PROPERTY       NAS-Port-Id             +12
PROPERTY       State                   +1
PROPERTY       Called-Station-Id       +12
PROPERTY       Calling-Station-Id      +12
PROPERTY       Acct-Status-Type        +2
PROPERTY       Acct-Session-Id         +2
PROPERTY       Acct-Session-Time       +2
‘raddb/config’
        auth {
                max-requests 127;
                request-cleanup-delay 2;
                compare-attribute-flag 1;
        };
        acct {
                max-requests 127;
                request-cleanup-delay 2;
                compare-attribute-flag 2;
        };

5.1.2 List of attributes that can be declared comparable.

The following attributes can be declared as comparable:

• User-Name

  • Password
  • CHAP-Password
  • NAS-Port-Id
  • State
  • Called-Station-Id
  • Calling-Station-Id
  • NAS-Identifier
  • Acct-Status-Type
  • Acct-Session-Id
  • Acct-Session-Time
  • User-UID
  • User-GID

Notice that this list is by no means an exhaustive one. Depending on a particular nas other attributes may be safe to be used in comparisons, or, vice-versa, some attributes from this list may not be used. You should carefully analyze packets coming from your nas before deciding which attributes to mark as comparable.

5.2 Fine-Tuning the Request Queue

As described in Section 2.4.1 [Checking Duplicates], page 12, each request is added to the request queue when radiusd starts processing it and is removed from there a certain amount of time after its processing was finished. The configuration parameter request-cleanup-delay defines how long each already processed request is kept in the queue. Its value must be synchronized with the nas settings.

Each nas allows to configure two parameters: Ntimeout The amount of time in seconds during which the nas is waiting

for a response from radius server.

Nretries The number of times the nas tries to re-send the request if it received no response from the radius server.

Of course, these parameters are named differently for different makes of nases. Refer to your nas documentation to find out where these values are configured.

In general, these parameters must satisfy the following relation:

          request-cleanup-delay = Nretries * Ntimeout + const

where const is an empirical constant that depends on the average time of processing a single request. Usually its value lies between 0 and 10 seconds.

For example, if the configuration of your nas sets

        Nretries = 3
        Ntimeout = 10

then your raddb/config should contain:

     auth {
             request-cleanup-delay 40;
     };
     acct {
             request-cleanup-delay 40;
     };

Notice the duplication of request-cleanup-delay: radiusd uses distinct values for authentication and accounting requests, however most existing nases do not make such distinction.

6 Authentication

An Authentication Type specifies which credentials the user is required to supply in order to be authenticated and where the user’s authentication data are stored. It is defined by the value of Auth-Type attribute in lhs of a ‘users’ entry.

6.1 Accept Authentication Type

Accept is the simplest authentication type. Users with this authentication type will be authenticated successfully without checking any credentials. Actually this means that only username is required for authentication.

This authentication type is used for each ‘users’ entry, whose lhs contains

     Auth-Type = Accept

This authentication type can be used for guest accounts, e.g. the following profile in ‘users’:

     guest   Auth-Type = Accept,
                     Simultaneous-Use = 10
             Service-Type = Framed-User,
                     Framed-Protocol = PPP

allows up to 10 simultaneous guest PPP accounts. To log in using such guest account it is sufficient to use username ‘guest’ and any password.

6.2 Reject Authentication Type

The Reject authentication type causes the request to be rejected unconditionally. It can be used to disable a user account (For another method of disabling user accounts, see Section 4.10 [access.deny file], page 53).

This authentication type is used for each ‘users’ entry, whose lhs contains

     Auth-Type = Reject

6.3 Local Password Authentication Type

The Local Password authentication type allows to keep plaintext user passwords. Although the use of this authentication type is strongly discouraged for security reasons, this is the only authentication type that can be used with CHAP authentication.

There are two ways of using this authentication type

Specifying Passwords in users File.

To keep the plaintext passwords in ‘users’ file, the profile entry must follow this pattern:

     user-name  Auth-Type = Local,
                          User-Password = plaintext

The plaintext is the user’s plaintext password. Obviously, user-name may

not be DEFAULT nor BEGIN.

Specifying Passwords in SQL Database.

     user-name   Auth-Type = Local,
                           Password-Location = SQL

When the user is authenticated using such profile, its password is retrieved from the authentication database using auth_query. The configuration of SQL authentication is described in detail in Section 4.11.2 [Authentication Server Parameters], page 54.

6.4 Encrypted Password Authentication Type

The Encrypted Password type allows to keep user’s passwords encrypted via DES or MD5 algorithm. There are two ways of using this authentication type.

Specifying Passwords in users File.

     user-name  Auth-Type = Crypt-Local,
                          User-Password = crypt-pass

The Crypt-Password is a shortcut for the above notation:

     user-name  Crypt-Password = crypt-pass

Specifying Passwords in SQL Database.

     user-name   Auth-Type = Crypt-Local,
                           Password-Location = SQL

Using this profile, the user’s password is retrieved from the authentication database using auth_query. The configuration of SQL authentication is described in detail on Section 4.11.2 [Authentication Server Parameters], page 54.

The shortcut for this notation is Auth-Type = SQL.

In any case, the passwords used with this authentication type must be either DES or MD5 hashed.

6.5 System Authentication Type

The System authentication type requires that the user have a valid system account on the machine where the radius server is running. The use of this type is triggered by setting

     Auth-Type = System

in the lhs of a ‘users’ entry.

6.6 SQL Authentication Type

Setting Auth-Type = SQL or Auth-Type = Mysql in the lhs of a ‘users’ entry

is a synonym for

     Auth-Type = Crypt-Local, Password-Location = SQL

and is provided as a shortcut and for backward compatibility with previous versions of GNU Radius.

For description of SQL authentication, see Section 6.4 [Encrypted Password Auth], page 72. The configuration of SQL subsystem is described in Section 4.11 [sqlserver file], page 53.

6.7 PAM Authentication Type

PAM authentication type indicates that a user should be authenticated using PAM (Pluggable Authentication Module) framework. The simplest way of usage is:

     Auth-Type = PAM

Any user whose ‘users’ profile contains the above, will be authenticated via PAM, using service name ‘radius’. If you wish to use another service name, set it using Auth-Data attribute, e.g.:

     Auth-Type = PAM,
         Auth-Data = pam-service

6.8 Defining Custom Authentication Types

The are three ways to define custom authentication types:

1. Write a PAM module.

2. Use a Guile procedure.

3. Use an external program

You can write a PAM module implementing the new authentication type. Then, specifying Auth-Type = PAM allows to apply it (see Section 6.7 [PAM Auth], page 73).

Alternatively, you may write a Scheme procedure implementing the new authentication type. To apply it, use Scheme-Procedure attribute in rhs. The Auth-Type = Accept can be used in lhs if the whole authentication burden is to be passed to the Scheme procedure. For example, if one wrote a procedure my-auth, to apply it to all users, one will place the following profile in his ‘users’ file:

     DEFAULT  Auth-Type = Accept
              Scheme-Procedure = "my-auth"

For a discussion of how to write Scheme authentication procedures, See Section 10.3.2 [Authentication with Scheme], page 116.

The third way to implement your own authentication method is using an external program. This is less effective than the methods described above, but may be necessary sometimes. To invoke the program, use the following statement in the rhs of ‘users’ entry:

     Exec-Program-Wait = "progname args"

The progname must be the full path to the program, args — any arguments it needs. The usual substitutions may be used in args to pass any request attributes to the program (see Section 4.14 [Macro Substitution], page 64).

For a detailed description of Exec-Program-Wait attribute and an example of its use, see Section 13.3.7 [Exec-Program-Wait], page 181.

6.9 Multiple Login Checking

The number of sessions a user can have open simultaneously can be restricted by setting Simultaneous-Use attribute in the user’s profile lhs (see Section 13.3.25 [Simultaneous-Use], page 192). By default the number of simultaneous sessions is unlimited.

When a user with limited number of simultaneous logins authenticates himself, Radius counts the number of the sessions that are already opened by this user. If this number is equal to the value of Simultaneous-Use attribute the authentication request is rejected.

This process is run in several stages. First, Radius retrieves the information about currently opened sessions from one of its accounting databases. Then, it verifies whether all these sessions are still active. This pass is necessary since an open entry might be a result of missing Stop request. Finally, the server counts the sessions and compares their count with the value of

Simultaneous-Use attribute.

The following subsections address each stage in detail.

6.9.1 Retrieving Session Data

Radius retrieves the list of sessions currently opened by the user either from the system database (see Section 7.1 [System Accounting], page 81), or from the SQL database (see Section 7.3 [SQL Accounting], page 83). The system administrator determines which method to use.

By default, system accounting database is used. Its advantages are simplicity and ease of handling. It has, however, a serious deficiency: the information is kept in the local files. If you run several radius servers, each of them has no easy way of knowing about the sessions initiated by other servers.

This problem is easy to solve if you run SQL accounting (see Section 7.3 [SQL Accounting], page 83). In this case, each radius server stores the data in your SQL database and can easily retrieve them from there.

To enable use of SQL database for multiple login checking, do the following:

In your ‘raddb/config’ file set:

     mlc {
         method sql;
     };

In your ‘raddb/sqlserver’ file, specify the queries for retrieving the information about open sessions and, optionally, a query to close an existing open record.

There are two queries for retrieving the information: mlc_user_query returns the list of sessions opened by the user, mlc_realm_query returns the list of sessions opened for the given realm. Each of them should return a list of 5-element tuples^{1}:

     user-name, nas-ip-address, nas-port-id, acct-session-id
   Here is an example of mlc_user_query and mlc_realm_query:
     mlc_user_query SELECT user_name,nas_ip_address,\
                           nas_port_id,acct_session_id \
                    FROM calls \
                    WHERE user_name='%C{User-Name}' \
                    AND status = 1
     mlc_realm_query SELECT user_name,nas_ip_address,\
                            nas_port_id,acct_session_id \
                     FROM calls \
                     WHERE realm_name='%C{Realm-Name}'

Apart from these two queries you may also wish to provide a query for closing a hung record. By default, radiusd will use acct_stop_query. If you wish to override it, supply a query named mlc_stop_query, for example:

     mlc_stop_query UPDATE calls \
                    SET status=4,\
                     acct_session_time=unix_timestamp(now())-\
                                       unix_timestamp(event_date_time) \
                    WHERE user_name='%C{User-Name}' \
                      AND status = 1 \
                      AND acct_session_id='%C{Acct-Session-Id}'

See Section 4.11.4.1 [Queries], page 58, for detailed information on how to write these queries.

6.9.2 Verifying Active Sessions

Whatever database radiusd uses, an open entry in it does not necessary mean that the corresponding session is still being active. So, after retrieving the information about user sessions, Radius verifies on corresponding nases whether these are actually active.

For each entry in the session list, if its nas acknowledges the session, the session count is incremented. Otherwise, such entry is marked as closed in the database and is not counted.

There may also be cases when the nas is unreachable due to some reasons. In such cases the Radius behavior is determined by the value of checkrad-assume-logged in ‘config’ file Section 4.1.3 [auth statement (raddb/config)], page 28. If the value is yes, Radius assumes the session is still active and increases the session count, otherwise it proceeds as if the nas returned negative reply.

1 This interface is likely to change in future versions

To query a nas, Radius first looks up its type and additional parameters in ‘naslist’ file (see Section 4.4 [naslist file], page 45). There are two predefined nas types that cause Radius to act immediately without querying tne nas: the special type ‘true’ forces Radius to act as if the nas returned 1, the type ‘false’ forces it to act as if the nas returned 0. If the type is neither of this predefined types, Radius uses it as a look up key into the ‘nastypes’ file (see Section 4.5 [nastypes file], page 47) and tries to retrieve an entry which has matching type. If such entry does not exist, Radius issues the error message and acts accordingly to the value of configuration variable checkrad-assume-logged. Otherwise, Radius determines the query method to use from the second field of this entry, and constructs method arguments by appending arguments from the ‘naslist’ entry to those of nastypes entry. Note, that the former take precedence over the latter, and can thus be used to override default values specified in ‘nastypes’.

Having determined the query method and its argument, Radius queries nas and analyzes its output by invoking a user-supplied Rewrite function. The function to use is specified by the function= argument to the method. It is called each time a line of output is received from the nas (for finger queries) or a variable is received (for SNMP queries). The process continues until the function returns 1 or the last line of output is read or a timeout occurs whichever comes first.

If the user-function returns 1 it is taken to mean the user’s session is now active at the nas, otherwise, if it replies 0 or if the end of output is reached, it is taken to mean the user’s session is not active.

The syntax conventions for user-supplied functions are described in detail in Section 10.2.5 [Login Verification Functions], page 103.

6.10 Controlling Authentication Probes

Authentication probe is an attempt of a user to use other user’s account, by guessing his password. The obvious indication of an authentication probe is appearence of several consecutive authentication failures for the same user. Of course, if the intruder is given sufficient number of such probes he will sooner or later succeed in finding the actual password. The conventional method to prevent this from occurring is to keep failure counters for each user and to lock the account when its failure countrer reaches a predefined limit. Notice that a legitimate user may fail (sometimes even several times in sequence) in entering his password so, two important points should always be observed. First, failure counters record the number of consecutive authentication failures and they are reset after each successive authentication.

Secondly, the maximum number of allowed consecutive failures should be set sufficiantly high.

The version 1.3 offers two ways for controlling authentication probes: using external programs and using special SQL queries.

To control authentication probes using external programs, use the com-

bination of Exec-Program-Wait and Auth-Failure-Trigger. The program

specified by Auth-Failure-Trigger is executed each time an authentication attempt failed. When both attributes are used together, the program invoked by Auth-Failure-Trigger can update the failure counter, and the one invoked by Exec-Program-Wait can compare the counter value with the predefined limit and reject authentication when both values become equal. Such approach is most useful in conjunction with BEGIN profile.

Let’s suppose the program ‘/sbin/check_failure’ accepts a user name and returns 1 if the failure counter for this user has reached maximum allowed value. Otherwise it returns 0 and clears the counter. Another program, ‘/sbin/count_failure’ increases failure counter value for the given user name. Assuming our basic authentication type is ‘PAM’, the ‘raddb/users’ file will look as follows:

     BEGIN   NULL
             Exec-Program-Wait = "/sbin/check_failure  %C{User-Name}",
             Auth-Failure-Trigger = "/sbin/count_failure %C{User-Name}",
                     Fall-Through = Yes
     DEFAULT Auth-Type = PAM
             Service-Type = Framed-User,
                     Framed-Protocol = PPP
     [... Other profiles ...]

The BEGIN profile will be executed before any other profile. It will add

to the rhs Exec-Program-Wait and Auth-Failure-Trigger attributes and

then radiusd will proceed to finding a matching profile (due to Fall-Through attribute). When such profile is found, the user will be authenticated according to the method set up by the profile’s Auth-Type attribute. If authentication fails, ‘/sbin/count_failure’ will be called and the user name passed to it as the argument. Otherwise, ‘/sbin/check_failure’ will be invoked.

To complete the example, here are working versions of both programs. Failure counters for each user name are kept in separate file in ‘/var/log/radius/fails’ directory. Both programs are written in bash.

The /sbin/count failure program

     #! /bin/bash
     test $# -eq 1 || exit 1
     MAXFAIL=8
     REGDIR=/var/log/radius/fails
     if [ -r "$REGDIR/$1" ]; then
       read COUNT < "$REGDIR/$1"
       COUNT=$((COUNT+1))
     else
       COUNT=1
     fi
     echo $COUNT > "$REGDIR/$1"
     # End of /sbin/count_failure

The /sbin/check failure program

     #! /bin/bash
     test $# -eq 1 || exit 1
     MAXFAIL=8
     REGDIR=/var/log/radius/fails
     if [ -r "$REGDIR/$1" ]; then
       read COUNT < "$REGDIR/$1"
       if [ $COUNT -ge $MAXFAIL ]; then
         echo "Reply-Message=\"Too many login failures. Your account is locked\""
         exit 1
       else
         rm "$REGDIR/$1"
       fi
     fi
     exit 0
     # End of check_failure

Another way of controlling authentication probes is by using SQL database to store failure counters. Two queries are provided for this purpose in

‘raddb/sqlserver’ file: auth_success_query is executed upon each suc-

cessful authentication, and auth_failure_query is executed upon each authentication failure. Both queries are not expected to return any values. One obvious purpose of auth_failure_query would be to update failure counters and that of auth_success_query would be to clear them. The auth_query or group_query should then be modified to take into account the number of authentication failures.

The default SQL configuration GNU Radius is shipped with provides a working example of using these queries. Let’s consider this example.

First, we create a special table for keeping authentication failure counters for each user:

     CREATE TABLE authfail (
       # User name this entry refers to
       user_name           varchar(32) binary default '' not null,
       # Number of successive authentication failures for this user
       count               int,
       # Timestamp when this entry was last updated
       time                datetime DEFAULT '1970-01-01 00:00:00' NOT NULL,
       # Create a unique index on user_name
       UNIQUE uname (user_name)
     );

The query auth_fail_query will increment the value of count column for the user in question:

     auth_failure_query UPDATE authfail \
                        SET count=count+1,time=now() \
                        WHERE user_name='%C{User-Name}'
   The query auth_success_query will clear count:
     auth_success_query UPDATE authfail \
                        SET count=0,time=now() \
                        WHERE user_name='%C{User-Name}'

Now, the question is: how to use this counter in authentication? The answer is quite simple. First, let’s create a special group for all the users whose authentication failure counter has reached its maximum value. Let this group be called ‘*LOCKED_ACCOUNT*’. We’ll add the following entry to

‘raddb/users’:
     DEFAULT Group = "*LOCKED_ACCOUNT*",
                     Auth-Type = Reject
             Reply-Message = "Your account is currently locked.\n\
     Please, contact your system administrator\n"

which will reject all such users with an appropriate reply message.

The only thing left now is to rewrite group_query so that it returns

‘*LOCKED_ACCOUNT*’ when authfail.count reaches its maximum value.

Let’s say this maximum value is 8. Then the following query will do the job:

     group_query       SELECT user_group FROM groups \
                       WHERE user_name='%u' \
                       UNION \
                       SELECT CASE WHEN (SELECT count > 8 FROM authfail \
                                                      WHERE user_name='%u')
                              THEN '*LOCKED_ACCOUNT*' END

The default configuration comes with these queries commented out. It is up to you to uncomment them if you wish to use SQL-based control over authentication failures.

Notice the following important points when using this approach:

1. Your SQL server must support UNION. Earlier versions of MySQL lacked this support, so if you run MySQL make sure you run a reasonably new version (at least 4.0.18).

2. Both auth_failure_query and auth_success_query assume the data-

base already contains an entry for each user. So, when adding a new user to the database, make sure to insert an appropriate record into authfails table, e.g.

     INSERT INTO authfail VALUES('new-user',0,now());

7 Accounting

By default GNU Radius supports three types of accounting. Any additional accounting methods can be defined using extension mechanisms.

The accounting methods are applied to a request in a following sequence:

1. System accounting

2. Detailed request accounting

3. sql accounting

4. Custom accounting Any method can be enabled or disabled. Thus, you can even disable them

all, thereby disabling accounting altogether.

Notice, that the multiple login checking scheme relies on accounting being enabled. By default it uses system accounting, but can also be configured to use sql accounting. So, if you disable system accounting and still wish to use reliable multiple login checking, make sure you configure radiusd to use sql for this purpose. See Section 6.9 [Multiple Login Checking], page 74, for the detailed information about the subject.

If any accounting type in this sequence fails, the accounting is deemed to fail and all subsequent methods are not invoked.

7.1 System Accounting

Radius keeps files ‘radutmp’ and ‘radwtmp’ in its logging directory and stores the accounting data there. The utilities radwho and radlast can be used to list information about users’ sessions.

This accounting method is enabled by default. To disable it, use system no statement in ‘raddb/config’. See Section 4.1.4 [acct], page 30, for more information. Please notice that disabling this authentication method will disable multiple login checking as well. Refer to Section 6.9 [Multiple Login Checking], page 74, for the detailed discussion of this.

7.2 Detailed Request Accounting

Radius stores the detailed information about accounting packets it receives in files ‘radacct/nasname/detail’ (see Chapter 1 [Naming Conventions], page 5), where nasname is replaced with the short name of the nas from the ‘raddb/naslist’ file (see Section 4.4 [naslist file], page 45).

By default, this accounting type is always enabled, provided that ‘radacct’ directory exists and is writable (see Chapter 1 [Naming Conventions], page 5). To turn the detailed accounting off, use the detail statement in the ‘config’ file. For more information about it, see Section 4.1.4 [acct], page 30.

The accounting detail files consist of a record for each accounting request. A record includes the timestamp and detailed dump of attributes from the packet, e.g.:

     Fri Dec 15 18:00:24 2000
             Acct-Session-Id = "2193976896017"
             User-Name = "e2"
             Acct-Status-Type = Start
             Acct-Authentic = RADIUS
             Service-Type = Framed-User
             Framed-Protocol = PPP
             Framed-IP-Address = 11.10.10.125
             Calling-Station-Id = "+15678023561"
             NAS-IP-Address = 11.10.10.11
             NAS-Port-Id = 8
             Acct-Delay-Time = 0
             Timestamp = 976896024
             Request-Authenticator = Unverified
     Fri Dec 15 18:32:09 2000
             Acct-Session-Id = "2193976896017"
             User-Name = "e2"
             Acct-Status-Type = Stop
             Acct-Authentic = RADIUS
             Acct-Output-Octets = 5382
             Acct-Input-Octets = 7761
             Service-Type = Framed-User
             Framed-Protocol = PPP
             Framed-IP-Address = 11.10.10.125
             Acct-Session-Time = 1905
             NAS-IP-Address = 11.10.10.11
             NAS-Port-Id = 8
             Acct-Delay-Time = 0
             Timestamp = 976897929
             Request-Authenticator = Unverified

Notice that radiusd always adds two pseudo-attributes to detailed listings. Attribute Timestamp shows the UNIX timestamp when radiusd has received the request. Attribute Request-Authenticator shows the result of checking the request authenticator. Its possible values are:

Verified The authenticator check was successful.

Unverified The authenticator check failed. This could mean that either the request was forged or that the remote nas and radiusd do not agree on the value of the shared secret.

None The authenticator check is not applicable for this request type.

Notice also that the so-called internal attributes by default are not logged in the detail file. Internal attributes are those whose decimal value is greater than 255. Such attributes are used internally by radius and cannot be transferred via radius protocol. Examples of such attributes are Fall-Through, Hint and Huntgroup-Name. See Section 13.3 [Radius Internal Attributes], page 178, for detailed listing of all internal attributes. The special attribute flag l (lower-case ell) may be used to force logging of such attributes (see Section 4.2.4 [ATTRIBUTE], page 41).

7.3 sql Accounting

The sql accounting method is enabled when Radius is configured with ‘--enable-sql’ option and the ‘sqlserver’ file in its configuration directory is properly set up (see Section 4.11 [sqlserver file], page 53).

This version of GNU Radius (1.3) supports MySQL and PostgreSQL servers. It also supports odbc, which can be used to build interfaces to another database management systems.

With this accounting method enabled, radiusd will store the information about accounting requests in the configured sql database. The accounting method is fully configurable: the Radius administrator defines both the types of requests to be accounted and the information to be stored into the database (see Section 4.11 [sqlserver file], page 53).

7.4 Defining Custom Accounting Types

If the built-in accounting methods do not meet your requirements, you can implement your own. There are two ways of doing so:

1. Using a Guile procedure.

2. Using an external program To use a Guile procedure for accounting, the name of the procedure must

be specified as a value to the Scheme-Acct-Procedure attribute in the rhs list of a ‘hints’ entry, e.g.:

     DEFAULT NULL Scheme-Acct-Procedure = "my-acct"

For a detailed description of Scheme accounting procedures, see Section 10.3.3 [Accounting with Scheme], page 117.

Another way of implementing your own accounting method is using an external program. This is less effective than the methods described above, but may be necessary sometimes. To invoke the program, use the following statement in the lhs of the ‘hints’ entry:

     Acct-Ext-Program = "progname args"

The progname must be the full path to the program, and args any arguments it needs. The usual substitutions may be used in args to pass any request attributes to the program (see Section 4.14 [Macro Substitution], page 64).

For a detailed description of Acct-Ext-Program, see Section 13.3.1 [Acct-Ext-Program], page 178.

84 GNU Radius Reference Manual

8 Logging

GNU Radius reports every event worth mentioning. The events are segregated by their severity level. Radius discerns the following levels (in order of increasing severity):

Debug The debug messages (Section 9.2 [Debugging], page 89). Auth Under this level every authentication attempt is logged. This is

enabled by setting

                 level auth;

in the category auth statement of the ‘config’ file.

Proxy Messages regarding proxy requests (see Section 2.4.2 [Proxying], page 13).

Info Informational messages.

Notice Normal, but significant conditions.

Warning Warning conditions. These mean some deviations from normal work.

Error Error conditions. Usually these require special attention.

CRIT Critical conditions due to which Radius is no longer able to continue working. These require urgent actions from the site administrator.

By default, all messages in all levels are output to the file ‘radlog/radius.log’. In addition, messages in level CRIT are also duplicated to the system console. These defaults can be overridden using logging statement in the ‘raddb/config’ file. (See Section 4.1.2 [logging statement], page 24, for the description of logging statement syntax; see Chapter 1 [Naming Conventions], page 5 for information about the locations of different Radius configuration files.)

86 GNU Radius Reference Manual

9 Problem Tracking

9.1 Rule Tracing

If you have more than one entry in your ‘users’ file it is not always obvious which of the entries were used for authentication. The authentication data flow becomes even harder to understand if there are some complex rules in

the ‘hints’ and ‘huntgroups’ files.

The rule tracing mode is intended to help you find out the exact order of the rules that each request matched during processing. The mode is toggled by trace-rules statement in auth or acct block of your ‘config’ file. When rule tracing mode is on for a given type of requests, radiusd will display the data flow diagram for each processed request of this type. The diagram is output on info logging category, it represents the list of rules in reverse chronological order. Each rule is represented by its location in the form filename:line. To make the output more compact, if several rules appear in the same configuration file, their locations are listed as a comma-separated list of numbers after the file name. Furthermore, if the configuration files have the same path prefix, then only the first file name appears with the full prefix.

Here is an example of trace rule diagram:

     Oct 31 11:37:17 [28322]: Auth.info: (Access-Request foo 170 bar):
     rule trace: /etc/raddb/users:157,22,3; huntgroups:72; hints:34

This diagram means, that the authentication request from server ‘foo’ for user ‘bar’ with ID 170 matched the following rules

File name  Line number

‘/etc/raddb/hints’  34
‘/etc/raddb/huntgroups’  72
‘/etc/raddb/users’  3
‘/etc/raddb/users’  22
‘/etc/raddb/users’  157

As a practical example, let’s suppose you have the following setup. There are three classes of users:

1. Users from group “root” are authenticated using system password database and get rlogin access to the server 192.168.10.1

2. Users from group “staff” are also authenticated using system password database, but they are granted only telnet access to the server 192.168.10.2

3. Finally, the rest of users is authenticated against SQL database and get usual PPP access. In addition, users from the first two classes are accounted using custom

Scheme procedure staff-acct.

The configuration files for this setup are showed below: Contents of ‘hints’:

DEFAULT  Group = "root"
         Scheme-Acct-Procedure = "staff-acct",
                   Hint = "admin"
DEFAULT  Group = "staff"
         Scheme-Acct-Procedure = "staff-acct",
                   Hint = "staff"

Contents of file ‘users’:

DEFAULT Auth-Type = SQL,
              Simultaneous-Use = 1
        Service-Type = Framed-User,
              Framed-Protocol = PPP
DEFAULT Hint = "admin",
             Auth-Type = System
        Service-Type = Login-User,
             Login-IP-Host = 192.168.0.1,
             Login-Service = Rlogin
DEFAULT Hint = "staff",
              Auth-Type = System,
              Simultaneous-Use = 1
         Service-Type = Login-User,
              Login-IP-Host = 192.168.0.2,
              Login-Service = Telnet

Now, let’s suppose that user ‘svp’ is in the group ‘staff’ and is trying to log in. However, he fails to do so and in radiusd logs you see:

 Nov 06 21:25:24: Auth.notice: (Access-Request local 61 svp):
   Login incorrect [svp]

Why? To answer this question, you add to auth block of your ‘config’ the statement

trace-rules yes;

and ask user ‘svp’ to retry his attempt. Now you see in your logs:

 Nov 06 21:31:24: Auth.notice: (Access-Request local 13 svp):
   Login incorrect [svp]
 Nov 06 21:31:24: Auth.info: (Access-Request local 13 svp):
   rule trace: /etc/raddb/users:1, hints: 5

This means that the request for ‘svp’ has first matched rule on the line 1 of file ‘hints’, then the rule on line 1 of file ‘users’. Now you see the error: the entries in ‘users’ appear in wrong order! After fixing it your ‘users’ looks like:

     DEFAULT Hint = "admin",
                  Auth-Type = System
             Service-Type = Login-User,
                  Login-IP-Host = 192.168.0.1,
                  Login-Service = Rlogin
     DEFAULT  Hint = "staff",
                   Auth-Type = System,
                   Simultaneous-Use = 1
              Service-Type = Login-User,
                   Login-IP-Host = 192.168.0.2,
                   Login-Service = Telnet
     DEFAULT Auth-Type = SQL,
                   Simultaneous-Use = 1
             Service-Type = Framed-User,
                   Framed-Protocol = PPP

Now, you ask ‘svp’ to log in again, and see:

     Nov 06 21:35:14: Auth.notice: (Access-Request local 42 svp):
       Login OK [svp]
     Nov 06 21:35:14: Auth.info: (Access-Request local 42 svp):
       rule trace: /etc/raddb/users:7, hints: 5

Let’s also suppose that user ‘plog’ is not listed in groups “root” and “staff”, so he is supposed to authenticate using SQL. When he logs in, you see in your logs:

     Nov 06 21:39:05: Auth.notice: (Access-Request local 122 plog):
       Login OK [svp]
     Nov 06 21:39:05: Auth.info: (Access-Request local 122 plog):
       rule trace: /etc/raddb/users:14

9.2 Debugging

GNU Radius provides extensive debugging features. These are enabled either by the ‘--debug’ (‘-x’) command line option to radiusd (see Chapter 3 [Invocation], page 17), or by the level statement in the debug category (see Section 4.1.2 [logging statement], page 24). Both cases require as an argument a valid debug specification.

A debug specification sets the module for which the debugging should be enabled and the debugging level. The higher the level is, the more detailed information is provided. The module name and level are separated by an equal sign. If the level is omitted, the highest possible level (100) is assumed. The module name may be abbreviated to the first N characters, in which case the first matching module is selected. Several such specifications can be specified, in which case they should be separated by commas. For example, the following is a valid debug specification:

             proxy.c=10,files.c,config.y=1

It sets debug level 10 for module proxy.c, 100 for files.c, and 1 for

config.y.

The modules and debugging levels are subject to change from release to release.

9.3 Test Mode

Test mode is used to test various aspects of radius configuration, without starting the daemon. To enter test mode, run

     radiusd -mt

You will see usual radiusd diagnostics and the following two lines:

     ** TEST SHELL **
     (radiusd) _

The string ‘** TEST SHELL **’ indicates that radiusd has entered test mode, the string ‘(radiusd)’ is the shell prompt, indicating that radiusd is waiting for your commands.

The syntax of test shell command resembles that of Bourne shell: each command consists of a list of words separated by any amount of whitespace. Each word is either a sequence of allowed word characters (i.e. alphabetical characters, decimal digits, dashes and underscores), or any sequence of characters enclosed in a pair of double quotes. The very first word is a command verb, the rest of words are arguments to this command verb. A command verb may be used in its full form, in its abbreviated form (i.e. you may type only several first characters of the verb, the only condition being that they do not coincide with another command verb), or in it’s short form.

The first command you should know is help (or, in its short form, h). This command takes no arguments and displays the short summary of all the available commands. Here is an example of its output:

  (radiusd) help
  h       help                           Print this help screen
  q       query-nas NAS LOGIN SID PORT [IP]
                                         Query the given NAS
  g       guile                          Enter Guile
  rs      rewrite-stack [NUMBER]         Print or set the Rewrite
                                         stack size
  r       run-rewrite FUNCTION(args..)   Run given Rewrite function
  s       source FILE                    Source the given Rewrite file
  t       timespan TIMESPAN [DOW [HH [MM]]]
                                         Check the timespan interval
  d       debug LEVEL                    Set debugging level
  rd      request-define [PAIR [,PAIR]]  Define a request
  rp      request-print                  Print the request
  quit    quit                           Quit the shell

Each line of the output consists of three fields. The first field shows the short command form. The second one lists its full form and its arguments, optional arguments being enclosed in square brackets. The third field contains short textual description of the command.

[Test Shell Command] query-nas nas login sid port [ip]

[Test Shell Abbreviation] q

Queries the given nas about the session described by its arguments. This command is useful in testing simultaneous login verification (see Section 6.9 [Multiple Login Checking], page 74. Its arguments are

nas Specifies the nas to query. It cn be its short name as defined in ‘raddb/naslist’, or its fully qualified domain name, or its ip address.

login Name of the user whose session should be verified.

sid Session ID.

port Port number on the nas.

ip Framed ip address, assigned to the user.

The command displays the following result codes:

0 The session is not active.

1 The session is active

-1 Some error occurred.

[Test Shell Command] guile [Test Shell Abbreviation] g

Enter Guile shell. The command is only available if the package has been compiled with Guile support. For more information, See Section 10.3 [Guile], page 115.

[Test Shell Command] rewrite-stack [number] [Test Shell Abbreviation] rs

Prints or sets the Rewrite stack size.

[Test Shell Command] run-rewrite function(args . . .)

[Test Shell Abbreviation] r

Runs given Rewrite function and displays its return value. Function arguments are specified in the usual way, i.e. as a comma-separated list of Rewrite tokens.

If the function being tested operates on request contents (see Section 10.2.4 [Rewriting Incoming Requests], page 99), you may supply the request using request-define command (see below).

[Test Shell Command] source file [Test Shell Abbreviation] s

Reads and compiles (“source”) the given Rewrite file. The command prints ‘0’ if the file was compiled successfully. Otherwise, it prints ‘1’ and any relevant diagnostics.

[Test Shell Command] timespan timespan [day-of-week [hour
      [minutes]]]

[Test Shell Abbreviation] t

Checks whether the given time falls within the timespan interval. Its first argument, timespan, contains the valid radiusd timespan specification (see Section 13.3.14 [Login-Time], page 187). Rest of arguments define the time. If any of these is omitted, the corresponding value from current local time is used.

day-of-week Ordinal day of week number, counted from 0. I.e.: Sunday – 0, Monday – 1, etc.

hour Hours counted from 0 to 24.

minutes Minutes.

The following set of samples illustrates this command:

Chapter 9: Problem Tracking 93

(radiusd) timespan Wk0900-1800
ctime: Tue Dec  2 16:08:47 2003
inside Wk0900-1800: 6720 seconds left
(radiusd) timespan Wk0900-1800 0
ctime: Sun Nov 30 16:09:03 2003
OUTSIDE Wk0900-1800: 60660 seconds to wait
(radiusd) timespan Wk0900-1800 0 12 30
ctime: Sun Nov 30 12:30:13 2003
OUTSIDE Wk0900-1800: 73800 seconds to wait
(radiusd) timespan Wk0900-1800 1 05 00
ctime: Mon Dec  1 05:00:33 2003
OUTSIDE Wk0900-1800: 14400 seconds to wait
(radiusd) timespan Wk0900-1800 1 09 10
ctime: Wed Jan  7 22:09:41 2004
OUTSIDE Wk0900-1800: 39060 seconds to wait
(radiusd) timespan Wk0900-1800 1 09 10
ctime: Mon Dec  1 09:10:44 2003
inside Wk0900-1800: 31800 seconds left
(radiusd)

[Test Shell Command] debug level [Test Shell Abbreviation] d

Set debugging level. Level is any valid debug level specification (see Section 9.2 [Debugging], page 89).

[Test Shell Command] request-define [pair [,pair]] [Test Shell Abbreviation] rd

Define a request for testing Rewrite functions. The optional arguments are a comma-separated list of a/v pairs. If they are omitted, the command enters interactive mode, allowing you to enter the desired a/v pairs, as in the following example:

(radiusd) request-define
Enter the pair list. End with end of file
[radiusd] User-Name = smith, User-Password = guessme
[radiusd] NAS-IP-Address = 10.10.10.1
[radiusd] NAS-Port-Id = 34
[radiusd]
(radiusd)

Notice that any number of a/v pairs may be specified in a line. To finish entering the request, either type an ^{⟨}\underline{\overline{EOF}}^{⟩} character or enter an empty line.

[Test Shell Command] request-print [Test Shell Abbreviation] rp

Prints the request, defined by request-define.

     (radiusd) request-print
         User-Name = (STRING) smith
         User-Password = (STRING) guessme
         NAS-IP-Address = (IPADDR) 10.10.10.1
         NAS-Port-Id = (INTEGER) 34
     (radiusd)

[Test Shell Command] quit

Immediately quits the shell.

10 Extensions

The use of extension language allows extending the functionality of GNU Radius without having to modify its source code. The two extension languages supported are Rewrite and Scheme. Use of Rewrite is always enabled. Use of Scheme requires Guile version 1.4 or higher.

10.1 Filters

The simplest way to extend the functionality of Radius is to use filters. A filter is an external program that communicates with Radius via its standard input and output channels.

10.1.1 Getting Acquainted with Filters

Suppose we wish to implement an authentication method based on the user name and the user’s calling station ID. We have a database of user names with valid IDs, and the new method should authenticate a user only if the combination {user name, id} is found in this database.

We write a filter program that reads its standard input line by line. Each input line must consist of exactly two words: the user name and the calling station ID. For each input line, the program prints 0 if the {user name, id} is found in the database and 1 otherwise. Let’s suppose for the sake of example that the database is a plaintext file and the filter is written in a shell programming language. Then it will look like

     #! /bin/sh
     DB=/var/db/userlist
     while read NAME CLID
     do
         if grep "$1:$2" $DB; then
             echo "0"
         else
             echo "1"
         fi
     done

10.1.2 Declaring the Filter

Here is how this filter is declared in the ‘raddb/config’ file:

     filters {
         filter check_clid {
             exec-path "/usr/libexec/myfilter";
             error-log "myfilter.log";
             auth {
                 input-format "%C{User-Name}
                 %C{Calling-Station-Id}";
                 wait-reply yes;

        };
    };
};

Let’s analyze this declaration line by line:

1. filters {

This keyword opens the filters declaration block. The block may contain several declarations.

2. filter check_clid {

This line starts the declaration of this particular filter and names it

‘check_clid’.
3. exec-path "/usr/libexec/myfilter";

This line tells radiusd where to find the executable image of this filter.

4. error-log "myfilter.log";

The diagnostic output from this filter must be redirected to the file ‘myfilter.log’ in the current logging directory

5. auth {

This filter will process authentication requests.

6. input-format "%C{User-Name} %C{Calling-Station-Id}";

Define the input line format for this filter. The %C{} expressions will be replaced by the values of the corresponding attributes from the incoming request (see Section 4.14 [Macro Substitution], page 64).

7. wait-reply yes;

radiusd will wait for the reply from this filter to decide whether to authenticate the user.

10.1.3 Invoking the Filter from a User Profile

To invoke this filter from the user profile, specify its name prefixed with ‘|’ in the value of Exec-Program-Wait attribute, like this:

DEFAULT Auth-Type = System,
                Simultaneous-Use = 1
        Exec-Program-Wait = "|check_clid"

10.1.4 Adding Reply Attributes

Apart from simply deciding whether to authenticate a user, the filter can also modify the reply pairs.

#! /bin/sh
DB=/var/db/userlist
while read NAME CLID
do
    if grep "$1:$2" $DB; then
        echo "0 Service-Type = Login, Session-Timeout = 1200"

Chapter 10: Extensions 97

else
    echo "1 Reply-Message =
          \"You are not authorized to log in\""
fi
done

10.1.5 Accounting Filters

Let’s suppose we further modify our filter to also handle accounting requests. To discern between the authentication and accounting requests we’ll prefix each authentication request with the word ‘auth’ and each accounting request with the word ‘acct’. Furthermore, the input line for accounting requests will contain a timestamp.

Now, our filter program will look as follows:

#! /bin/sh
AUTH_DB=/var/db/userlist
ACCT_DB=/var/db/acct.db
while read CODE NAME CLID DATE
do
case CODE
auth)
    if grep "$1:$2" $DB; then
        echo "0 Service-Type = Login, \
              Session-Timeout = 1200"
    else
        echo "1 Reply-Message = \
              \"You are not authorized to log in\""
    fi
acct)
    echo "$CODE $NAME $CLID $DATE" >> $ACCT_DB
done

Its declaration in the ‘raddb/config’ will also change:

filter check_clid {
exec-path "/usr/libexec/myfilter";
error-log "myfilter.log";
auth {
    input-format "auth %C{User-Name}
                  %C{Calling-Station-Id}";
    wait-reply yes;
};
acct {
    input-format "acct %C{User-Name}
                  %C{Calling-Station-Id} %D";
    wait-reply no;
};
};

(The input-format lines are split for readability. Each of them is actually one line).

Notice wait-reply no in the acct statement. It tells radiusd that it shouldn’t wait for the response on accounting requests from the filter.

10.1.6 Invoking the Accounting Filter

To invoke the accounting filter, specify its name prefixed with a vertical bar character as a value of Acct-Ext-Program in our ‘raddb/hints’ file. For example:

     DEFAULT NULL
             Acct-Ext-Program = "|check_clid:

10.2 Rewrite

Rewrite is the GNU Radius extension language. Its name reflects the fact that it was originally designed to rewrite the broken request packets so they could be processed as usual (see Section 10.2.4 [Rewriting Incoming Requests], page 99). Beside this basic use, however, Rewrite functions are used to control various aspects of GNU Radius functionality, such as verifying the activity of user sessions, controlling the amount of information displayed in log messages, etc (see Section 10.2.3 [Interaction with Radius], page 99).

10.2.1 Syntax Overview

The syntax of Rewrite resembles that of C. Rewrite has two basic data types: integer and string. It does not have global variables; all variables are automatic. The only exceptions are the a/v pairs from the incoming request, which are accessible to Rewrite functions via the special notation

%[attr].

10.2.2 Quick Start

As an example, let’s consider the following Rewrite function:

     string
     foo(integer i)
     {
         string rc;
         if (i % 2)
             rc = "odd";
         else
             rc = "even";
         return "the number is " + rc;
     }

The function takes an integer argument and returns the string ‘the number is odd’ or ‘the number is even’, depending on the value of i. This illustrates the fact that in Rewrite the addition operator is defined on the string type. The result of such operation is the concatenation of operands.

Another example is a function that adds a prefix to the User-Name attribute:

     integer
     px_add()
     {
             %[User-Name] = "pfx-" + %[User-Name];
             return 0;
     }

This function manipulates the contents of the incoming request; its return value has no special meaning.

10.2.3 Interaction with Radius

A Rewrite function can be invoked in several ways, depending on its purpose. There are three major kinds of Rewrite functions:

• Functions used to rewrite the incoming requests.

• Functions designed for simultaneous login verification.

• Functions used to generate Radius attribute values.

• Logging hooks.

10.2.4 Rewriting Incoming Requests

The need for rewriting the incoming requests arises from the fact that some nases are very particular about the information they send with the requests. There are cases when the information they send is hardly usable or even completely unusable. For example, a Cisco as5300 terminal server used as a voice-over IP router packs a lot of information into its Acct-Session-Id attribute. Though the information stored there is otherwise relevant, it makes proper accounting impossible, since the Acct-Session-Id attributes in the start and stop packets of the same session become different, and thus Radius cannot determine the session start to which the given session stop request corresponds (see Section 13.2.7 [Acct-Session-Id], page 177).

In order to cope with such nases, GNU Radius is able to invoke a Rewrite function upon arrival of the packet and before processing it further. This function can transform the packet so that it obtains the form prescribed by rfcs and its further processing becomes possible.

For example, in the case of the as5300 router, a corresponding Rewrite function parses the Acct-Session-Id attribute; breaks it down into fields; stores them into proper attributes, creating them if necessary; and finally replaces Acct-Session-Id with its real value, which is the same for the start and stop records corresponding to a single session. Thus all the information that came with the packet is preserved, but the packet itself is made usable for proper accounting.

A special attribute, Rewrite-Function, is used to trigger invocation of a Rewrite function. Its value is a name of the function to be invoked.

When used in a ‘naslist’ profile, the attribute causes the function to be invoked when the incoming request matches the huntgroup (see Section 2.4.4 [Huntgroups], page 15). For example, to have a function fixup invoked for each packet from the nas 10.10.10.11, the following huntgroup rule may be used:

     DEFAULT  NAS-IP-Address = 11.10.10.11
              Rewrite-Function = "fixup"

The Rewrite-Function attribute may also be used in a ‘hints’ rule. In this case, it will invoke the function if the request matches the rule (see Section 2.4.3 [Hints], page 14). For example, this ‘hints’ rule will cause the function to be invoked for each request containing the user name starting with ‘P’:

     DEFAULT  Prefix = "P"
              Rewrite-Function = "fixup"

Note that in both cases the attribute can be used either in lhs or in rhs pairs of a rule.

The packet rewrite function must be declared as having no arguments and returning an integer value:

     integer fixup()
     {
     }

The actual return value from such a function is ignored, the integer return type is just a matter of convention.

The following subsection present some examples of packet rewrite functions.

10.2.4.1 Examples of Various Rewrite Functions

The examples found in this chapter are working functions that can be used with various existing nas types. They are taken from the ‘rewrite’ file contained in distribution of GNU Radius.

1. Port rewriting for max ascend terminal servers

Some max ascend terminal servers pack additional information into the NAS-Port-Id attribute. The port number is constructed as XYYZZ, where X = 1 for digital, X = 2 for analog, YY is the line number (1 for first PRI/T1/E1, 2 for second, and so on), and ZZ is the channel number (on the PRI or channelized T1/E1).

The following rewrite functions are intended to compute the integer port number in the range (1 .. portcnt), where portcnt represents the real number of physical ports available on the nas. Such a port number can be used, for example, to create a dynamic pool of IP addresses (see Section 13.1.8 [Framed-IP-Address], page 167).

     /*
      * decode MAX port number
      * input: P        --  The value of NAS-Port-Id attribute
      *        portcnt  --  number of physical ports on the NAS
      */
     integer
     max_decode_port(integer P, integer portcnt)
     {
         if (P > 9999) {
             integer s, l, c;
             s = P / 10000;
             l = (P - (10000 * s))/100;
             c = P - ((10000 * s) + (100 * l));
             return (c-1) + (l-1) * portcnt;
         }
         return P;
     }
     /*
      * Interface function for MAX terminal server with 23 ports.
      * Note that it saves the received NAS-Port-Id attribute in
      * the Orig-NAS-Port-Id attribute. The latter must be
      * defined somewhere in the dictionary
      */
     integer
     max_fixup()
     {
         %[Orig-NAS-Port-Id] = %[NAS-Port-Id];
                                       # Preserve original data
         %[NAS-Port-Id] = max_decode_port(%[NAS-Port-Id], 23);
         return 0;
     }

2. Session id parsing for Cisco AS 5300 series

Cisco voip ios encodes a lot of other information into its Acct-Session-Id. The pieces of information are separated by ‘/’ characters. The part of Acct-Session-Id up to the first ‘/’ character is the actual session id.

On the other hand, its accounting packets lack NAS-Port-Id, though they may contain the vendor-specific pair with code 2 (vendor pec 9), which is a string in the form ‘ISDN 9:D:999’ (‘9’ represents any decimal digit). The number after the last ‘:’ character can be used as a port number.

The following code parses Acct-Session-Id attribute and stores the information it contains in various other attributes, generates a normal Acct-Session-Id, and attempts to generate a NAS-Port-Id attribute.

/*
* The port rewriting function for Cisco AS5300 used for
* VoIP. This function is used to generate NAS-Port-Id pair
* on the basis of vendor-specific pair 2. If the latter is
* in the form "ISDN 9:D:999" (where each 9 represents a
* decimal digit), then the function returns the number
* after the last colon. This is used as a port number.
*/
integer
cisco_pid(string A)
{
   if (A =~
       ".*\([0-9][0-9]*\):
        [A-Z0-9][A-Z0-9]*:\([0-9][0-9]*\)") {
       return (integer)\2;
   }
   return -1;
}
/*
* This function parses the packed session id.
* The actual sid is the number before the first slash
* character.  Other possibly relevant fields are also
* parsed out and saved in the Voip-* A/V pairs. The latter
* should be defined somewhere in the dictionary.
* Note that the regular expression in this example
* spans several lines for readability. It should be on one
* line in real file.
*/
string
cisco_sid(string S)
{
  if (S =~ "\(.[^/]*\)/[^/]*/[^/]*/\([^/]*\)/\([^/]*\)/
            \([^/]*\)/\([^/]*\)/\([^/]*\)/\([^/]*\)
            /\([^/]*\).*") {
       %[Voip-Connection-ID] = \2;
       %[Voip-Call-Leg-Type] = \3;
       %[Voip-Connection-Type] = \4;
       %[Voip-Connect-Time] = \5;
       %[Voip-Disconnect-Time] = \6;
       %[Voip-Disconnect-Cause] = \7;
       %[Voip-Remote-IP] = \8;
       return \1;
  }
  return S;
}

/*
 * Normalize cisco AS5300 packets
 */
integer
cisco_fixup()
{
    integer pid;
    if ((pid = cisco_pid(%[Cisco-PRI-Circuit])) != -1) {
        if (*%[NAS-Port-Id])
            %[Orig-NAS-Port-Id] = %[NAS-Port-Id];
        %[NAS-Port-Id] = pid;
    }
    if (*%[Acct-Session-Id]) {
        %[Orig-Acct-Session-Id] = %[Acct-Session-Id];
        %[Acct-Session-Id] = cisco_sid(%[Acct-Session-Id]);
    }
    return 0;
}

3. User-name rewriting for nt machines

Users coming from Windows nt machines often authenticate themselves as ‘NT_DOMAIN\username’. The following function selects the user-name part and stores it in the User-Name attribute:

integer
login_nt(string uname)
{
    integer i;
    if ((i = index(uname, '\\')) != -1)
        return substr(uname, i+1, -1);
    return uname;
}
integer
nt_rewrite()
{
    %[Orig-User-Name] = %[User-Name];
    %[User-Name] = login_nt(%[User-Name]);
    return 0;
}

10.2.5 Login Verification Functions

A login verification function is invoked to process the output from the nas. This process is described in Section 6.9 [Multiple Login Checking], page 74. The function to be invoked for given nas is defined by a function flag in the ‘raddb/nastypes’ or ‘raddb/naslist’ file (see Section 4.5 [nastypes file], page 47). It must be defined as follows:

[Function Template] integer check (string str, string name, integer

pid, string sid)

Its arguments are:

str Input string. If the query method is finger, this is the string of output received from the nas with trailing newline stripped off. If the query method is snmp, it is the received variable value converted to its string representation.

name User name.

pid Port id of the session.

sid Session id.

The function should return non-0 if its arguments match the user’s session, and 0 otherwise.

10.2.5.1 Examples of Login Verification Functions

As an example, let’s consider the function for analyzing a line of output from a standard UNIX finger service. In each line of finger output the first field contains the user name; the third field, the The function must return 1 if the three fields match the input user name and port and session ids:

     integer
     check_unix(string str, string name, integer pid, string sid)
     {
         return field(str, 1) == name
                && field(str, 3) == pid
                && field(str, 7) == sid;
     }

The next example is a function to analyze a line of output from an SNMP query returning a user name. This function must return 1 if the entire input line matches the user name:

     integer
     check_username(string str, string name, integer pid, string sid)
     {
         return str == name;
     }

10.2.6 Attribute Creation Functions

These are the functions used to create Radius reply attributes. An attribute creation function can take any number of arguments. The type of its return is determined by the type of Radius attribute the value will be assigned to. To invoke the function, write its name in the a/v pair of the rhs in the

‘raddb/users’ file, e.g.:
     DEFAULT Auth-Type = SQL
             Service-Type = Framed-User,
                 Framed-IP-Address = "=get_ip_addr(10.10.10.1)"

The function get_ip_addr will be invoked after successful authentication and it will be passed the IP 10.10.10.1 as its argument. An example of a useful function that can be invoked this way is

     integer
     get_ip_address(integer base)
     {
         return base + %[NAS-Port-Id] - %[NAS-Port-Id]/16;
     }

10.2.7 Logging Hook Functions

A logging hook functions should be declared as follows:

[Function Template] string hook (integer reqtype, string nasid,

integer reqid) reqtype Type of the request. It can be converted to string using

request_code_string function (see Section 10.2.8.7 [Builtin Functions], page 111).

nasid nas identifier from ‘raddb/naslist’, or its host name if not declared there

reqid Request identifier.

Notice that the hook function shall not produce any side effects, in particular it shall not modify the incoming request in any way.

Following is an example prefix hook function that formats the incoming request data:

     string
     compat_log_prefix(integer reqtype, string nas, integer id)
     {
             string result;
             return "(" + request_code_string(reqtype) + " "
                        + nas + " " + (string)id + " " + %[User-Name] + ")";
     }

Here is a sample log produced by radiusd before and after enabling this function:

     Auth.notice: Login OK [jsmith]
     ...
     Auth.notice: (AUTHREQ nas-2 251 jsmith): Login OK [jsmith]

10.2.8 Full Syntax Description

10.2.8.1 Rewrite Data Types

There are only two data types: integer and string, the two being coercible to each other in the sense that a string can be coerced to an integer if it contains a valid ascii representation of a decimal, octal, or hex number, and an integer can always be coerced to a string, the result of such coercion being the ascii string that is the decimal representation of the number.

10.2.8.2 Rewrite Symbols

A symbol is a lexical token. The following symbols are recognized:

Arithmetical operators These are ‘+’, ‘-’, ‘*’, ‘/’ representing the basic arithmetical operations, and ‘%’ meaning remainder.

Comparison operators These are: ‘==’, ‘!=’, ‘<’, ‘<=’, ‘>’, ‘>=’ with the same meaning they have in C. Special operators are provided for regularexpression matching. The binary operator ‘=~’ returns true if its left-hand-side operand matches the regular expression on its right-hand side (see Section 10.2.8.6 [Regular Expressions], page 110). ‘!~’ returns true if its left-hand-side operand does not match the regexp on its right-hand side. The right-hand-side operand of ‘!~’ or ‘=~’ must be a literal string, i.e., the regular expression must be known at compile time.

Unary operators The unary operators are ‘-’ and ‘+’ for unary plus and minus, ‘!’ for boolean negation, and ‘*’ for testing for the existence of an attribute.

Boolean operators These are ‘&&’ and ‘||’.

Parentheses ‘(’ and ‘)’ These are used to change the precedence of operators, to introduce type casts (type coercions), to declare functions, and to pass actual arguments to functions.

Curly braces (‘{’ and ‘}’)

These are used to delimit blocks of code.

Numbers Numbers follow the usual C convention for integers. A number consisting of a sequence of digits is taken to be octal if it begins with ‘0’ (digit zero), and decimal otherwise. If the sequence of digits is preceded by ‘0x’ or ‘0X’, it is taken to be a hexadecimal integer.

IP Numbers IP numbers are represented by a standard numbers-and-dots notation. IP numbers do not constitute a separate data type, rather they are in all respects similar to initeger numbers.

Characters These follow the usual C convention for characters, i.e., they consist either of an ascii character itself or of its value, enclosed in a pair of singlequotes. The character value begins with ‘\’ (backslash) and consists either of three octal or of two hexadecimal digits. A character does not form a special data type; it is represented internally by an integer.

Quoted strings These follow slightly modified C conventions for strings. A string is a sequence of characters surrounded by double quotes, as in ‘"..."’. In a string, the double quote character ‘"’ must be preceeded by a backslash ‘\’. A ‘\’ and an immediately following newline are ignored. Following escape sequences have special meaning:

\a Audible bell character (ASCII 7)

\b Backspace (ASCII 8)

\e Escape character (ASCII 27)

\f Form feed (ASCII 12)

\n Newline (ASCII 10)

\r Carriage return (ASCII 13)

\t Horizontal tab (ASCII 9)

\\ Backslash

\ooo (‘o’ represents an octal digit) A character whose ASCII value is represented by the octal number

          ‘ooo’.

\xHH \XHH (‘H’ represents a hex digit) A character whose ASCII

value is represented by the hex number ‘HH’.

\( Two characters ‘\(’.

\) Two characters ‘\)’.

If the character following the backslash is not one of those specified, the backslash is ignored.

Attribute values The incoming request is passed implicitly to functions invoked via the Rewrite-Function attribute. It is kept as an associative array, whose entries can be accessed using the following syntax:

     `%[' attribute-name `]'
     `%[' attribute-name `]' `(' n `)'

The first form returns the value of the attribute attribute-name. Here attribute-name should be a valid Radius dictionary name (see Section 4.2 [dictionary file], page 39). The second form returns the value of the nth attribute of type attribute-name. The index n is counted from zero, so

             %[attribute-name](0)

is equivalent to

             %[attribute-name]

Identifiers Identifiers represent functions and variables. These are described in the next sub-subsection.

Regexp group references

A sequence of characters in the form

     `\number'

refers to the contents of parenthesized group number number obtained as a result of the last executed ‘=~’ command. The regexp group reference has always string data type. For example:

     string
     basename(string arg)
     {
         if (arg =~ ".*/\(.*\)\..*")
             return \1;
         else
             return arg;
     }

This function strips from arg all leading components up to the last slash character, and all trailing components after the last dot character. It returns arg unaltered if it does not contain slashes and dots. It is roughly analogous to the system basename utility.

10.2.8.3 Rewrite Identifiers

A valid identifier is a string of characters meeting the following requirements:

1. It starts with either a lower- or an uppercase letter of the Latin alphabet or either of the following symbols: ‘_’, ‘$’.

2. It consists of alphanumeric characters, underscores(‘_’), and dollar signs (‘$’).

10.2.8.4 Rewrite Declarations

Function declarations

A Rewrite function is declared as follows:

type function-name (parameter-list)

where type specifies the return type of the function, function-name declares the symbolic name of the function, and parameter-list declares the formal parameters to the function. It is a comma-separated list of declarations in the form

     type parm-name

type being the parameter type, and parm-name being its symbolic name. Both function-name and parm-name should be valid identifiers.

Variable declarations

There are no global variables in Rewrite. All variables are local. The local variables are declared right after the opening curly brace (‘{’) and before any executable statements. The declaration syntax is

     type ident_list ;

Here ident list is either a valid Rewrite identifier or a comma-separated list of such identifiers.

Note that, unlike in C, no assignments are allowed in variable declarations.

10.2.8.5 Rewrite Statements

The Rewrite statements are: expressions, assignments, conditional statements, and return statements. A statement is terminated by a semicolon.

Expressions

An expression is one of the following:

• A variable identifier

• A type coercion expression

• An arithmetic expression

• A boolean expression

• An assignment

• A function call

• A delete statement

Type coercion

The type coercion is like a type cast in C. Its syntax is

     `(' type `)' ident

The result of type coercion is as follows:

type  Variable type  Resulting conversion

integer integer No conversion. This results in the same integer value. integer string If the string value of the variable is a valid ascii representation of the integer number (either decimal, octal, or hex), it is converted to the integer; otherwise the result of the conversion is undefined.

string integer The ascii representation (in decimal) of the integer number.

string string No conversion. This results in the same string value.

Assignment

An assignment is

     ident = expression ;

The variable ident is assigned the value of expression.

Function calls

These take the form

     ident ( arg-list )

where ident is the identifier representing the function, and arg-list is a comma-separated list of expressions supplying actual arguments to the function. The number of the expressions must correspond exactly to the number of formal parameters in the function definition. The function that ident references can be either a compiled function or a built-in function.

‘delete’ statement

The ‘delete’ statement is used to delete an attribute or attributes from the incoming request. Its syntax is:

     delete attribute-name;
     delete attribute-name(n);

The first variant deletes all the attributes of the given type. The second variant deletes only the nth occurrence of the matching attribute.

10.2.8.6 Regular Expressions

Rewrite uses POSIX regular expressions (See section “Regular Expression Library” in Regular Expression Library, for the detailed description of these).

You control the exact type of regular expressions by the use of the pragmatic comment regex. Its syntax is as follows:

     #pragma regex option-list

Option-list is a whitespace-separated list of options. Each option is one of the following words prefixed with ‘+’ or ‘-’:

extended Use POSIX extended regular expression syntax when interpreting regular expressions. icase Do not differentiate case. Subsequent regular expression com-

parisons will be case insensitive.

newline Match-any-character operators don’t match a newline. A non-matching list (‘[^...]’) not containing a newline does not match a newline. Match-beginning-of-line operator (‘^’) matches the empty string immediately after a newline. Match-end-of-line operator (‘$’) matches the empty string immediately before a newline.

Prefixing an option with ‘+’ means to enable the corresponding behavior. Prefixing it with ‘-’ means to disable it. Thus, the following statement:

     #pragma regex +extended +icase

will enable extended POSIX regular expressions using case-insensitive comparison.

Using the following comment:

     #pragma regex -extended

will switch to the basic POSIX regular expressions.

The settings of a regex pragmatic comment remain in force up to the end of current source file, or to the next regex comment, whichever occurs first.

For compatibility with previous versions, GNU Radius uses the following defaults:

     #pragma regex -extended -icase -newline

i.e. all regular expressions are treated as basic POSIX, comparison is casesensitive.

10.2.8.7 Rewrite Built-in Functions

The following built-in functions are provided:

[Function] integer length (string s)

Returns the length of the string s.

        length("string") ⇒ 6

[Function] integer index (string s, integer c)

Returns the index of the first occurrence of the character c in the string s. Returns −1 if no such occurrence is found.

        index("/raddb/users", 47) ⇒ 0
        index("/raddb/users", 45) ⇒ −1

[Function] integer rindex (string s, integer i)

Returns the index of the last occurrence of the character c in the string s. Returns −1 if no such occurrence is found.

        rindex("/raddb/users", 47) ⇒ 6

[Function] string substr (string s, integer start, integer length)

Returns the substring of s of length at most length starting at position start.

     substr("foo-bar-baz", 3, 5) ⇒ "-bar-"

All character positions in strings are counted from 0.

[Function] string field (string buffer, integer n)

This function regards the buffer argument as consisting of fields separated with any amount of whitespace. It extracts and returns the nth field. n is counted from 1.

     field("GNU's not UNIX", 1) ⇒ "GNU's"
     field("GNU's not UNIX", 2) ⇒ "not"
     field("GNU's not UNIX", 3) ⇒ "UNIX"
     field("GNU's not UNIX", 4) ⇒ ""

[Function] integer logit (string msg)

Outputs its argument to the Radius log channel info. Returns 0. For debugging purposes.

[Function] integer inet_aton (string str)

Converts the Internet host address str from the standard numbers-and-dots notation into the equivalent integer in host byte order.

     inet_aton("127.0.0.1") ⇒ 2130706433

[Function] string inet_ntoa (integer ip)

Converts the Internet host address ip given in host byte order to a string in standard numbers-and-dots notation.

     inet_ntoa(2130706433) ⇒ "127.0.0.1"

[Function] integer htonl (integer n)

Converts the integer n, regarded as long, from host to network byte order.

[Function] integer ntohl (integer n)

Converts the integer n, regarded as long, from network to host byte order.

[Function] integer htons (integer n)

Converts the integer n, regarded as short, from host to network byte order.

[Function] integer ntohs (integer n)

Converts the integer n, regarded as short, from network to host byte order.

[Function] string gsub (string regex, string repl, string str)

For each substring matching the regular expression regex in the string str, substitute the string repl, and return the resulting string.

     gsub("s","S","strings")
         ⇒ "StringS"
     gsub("[0-9][0-9]*","N","28 or 29 days")

         ⇒ "N or N days"
     gsub("[()'\"]","/","\"a\" (quoted) 'string'")
         ⇒ "/a/ /quoted/ /string/"

[Function] string qprn (string str)

Replace all non-printable characters in string S by their corresponding hex value preceeded by a percent sign. Return the resulting string. Printable are alphabetical characters, decimal digits and dash (‘-’). Other characters are considered non-printable. For example:

     qprn("a string/value") ⇒ "a%20string%2Fvalue"
[Function] string quote_string (string str)

Replace all non-printable characters in string str by their three-digit octal code prefixed with a backslash, or by their C escape notation, as appropriate. Non-printable characters depend on the locale settings. For example, suppose that the current locale is set to ISO-8859-1 (a so called “Latin-1” character set) and ⋆ represents a tab character. Then:

     quote_string("François contains non⋆printable chars")
       ⇒ "Fran\347ois contains non\tprintable chars"
[Function] string unquote_string (string str)

Replace C escape notations in string str with corresponding characters using current locale. For example, for ISO-8859-1 locale:

     unquote_string("Fran\347ois") ⇒ "François"

[Function] string toupper (string str)

Returns the copy of the string str with all alphabetical characters converted to upper case. For example:

     toupper("a-string") ⇒ "A-STRING"

[Function] string tolower (string str)

Returns the copy of the string str with all alphabetical characters converted to lower case. For example:

     tolower("A-STRING") ⇒ "a-string"
[Function] string request_code_string (integer code)

Converts integer RADIUS request code to its textual representation as per RFC 3575. This function is useful in logging hooks (see Section 4.1.2.1 [hooks], page 25).

     request_code_string(4) ⇒ "Accounting-Request"

Native Language Support

The native language support is provided via the functions described below. These functions are interfaces to GNU gettext library. For the information about general concepts and principles of Native Language Support, please refer to section “gettext” in GNU gettext utilities.

The default current textual domain is ‘radius’.

[Function] string textdomain (string domain)

Sets the new value for the current textual domain. This domain is used by the functions gettext and ngettext. Returns the name of the previously used domain.

[Function] string gettext (string msgid)

[Function] string _(string msgid)

The function returns the translation of the string msgid if it is available in the current domain. If it is not available, the argument itself is returned. The second form of this function provides a traditional shortcut notation. For a detailed description of the GNU gettext interface, refer to section “Interface to gettext” in GNU gettext utilities.

[Function] string dgettext (string domain, string msgid)

Returns the translation of the string msgid if it is available in the domain domain. If it is not available, the argument itself is returned.

[Function] string ngettext (string msgid_singular, string
      msgid_plural, integer number)

The ngettext function is used to translate the messages that have singular and plural forms. The msgid singular parameter must contain the singular form of the string to be converted. It is also used as the key for the search in the catalog. The msgid_plural parameter is the plural form. The parameter number is used to determine the plural form. If no message catalog is found msgid singular is returned if number == 1, otherwise msgid plural. For a detailed description of the GNU gettext interface for the plural translation, refer to section “Additional functions for plural forms” in GNU gettext utilities.

[Function] string dngettext (string domain, string msg_sing,

string msg_plur, integer number) Similar to ngettext, but searches translation in the given domain.

Request Accessors

The following functions are used to read some internal fields of a radius request.

[Function] Integer request_source_ip ()

Returns source ip address of the currently processed request. This function can be used to add NAS-IP-Address attribute to the requests lacking one, e.g.:

     integer
     restore_nas_ip()
     {
             if (!*%[NAS-IP-Address])
                     %[NAS-IP-Address] = request_source_ip();

                return 0;
        }
[Function] Integer request_source_port ()

Returns the source udp port.

[Function] Integer request_id ()

Returns the request identifier.

[Function] Integer request_code ()

Returns the request code.

10.3 Guile

The name Guile stands for GNU’s Ubiquitous Intelligent Language for Extensions. It provides a Scheme interpreter conforming to the R4RS language specification. This section describes use of Guile as an extension language for GNU Radius. It assumes that the reader is sufficiently familiar with the Scheme language. For information about the language, refer to section “Top” in Revised(4) Report on the Algorithmic Language Scheme. For more information about Guile, see section “Overview” in The Guile Reference Manual.

Scheme procedures can be called for processing both authentication and accounting requests. The invocation of a Scheme procedure for an authentication request is triggered by the Scheme-Procedure attribute; the invocation for an accounting request is triggered by the Scheme-Acct-Procedure attribute. The following sections address these issues in more detail.

10.3.1 Data Representation

a/v pair lists are the main object Scheme functions operate upon. Scheme is extremely convenient for representation of such objects. A Radius a/v pair is represented by a Scheme pair; e.g.,

             Session-Timeout = 10

is represented in Guile as

             (cons "Session-Timeout" 10)

The car of the pair can contain either the attribute dictionary name or the attribute number. Thus, the above pair may also be written in Scheme as

             (cons 27 10)

(because Session-Timeout corresponds to attribute number 27).

Lists of a/v pairs are represented by Scheme lists. For example, the Radius pair list

             User-Name = "jsmith",
                     User-Password = "guessme",
                     NAS-IP-Address = 10.10.10.1,
                     NAS-Port-Id = 10

is written in Scheme as

             (list
               (cons "User-Name" "jsmith")
               (cons "User-Password" "guessme")
               (cons "NAS-IP-Address" "10.10.10.1")
               (cons "NAS-Port-Id" 10))

10.3.2 Authentication with Scheme

The Scheme procedure used for authentication must be declared as follows:

[Function Template] auth-function request-list check-list reply-list Its arguments are:

request-list The list of a/v pairs from the incoming request

check-list The list of a/v pairs from the lhs of the profile entry that matched the request

reply-list The list of a/v pairs from the rhs of the profile entry that matched the request

The function return value determines whether the authentication will succeed. The function must return either a boolean value or a pair. The return of #t causes authentication to succeed. The return of #f causes it to fail.

For a function to add something to the reply a/v pairs, it should return a pair in the form

         (cons return-code list)

where return-code is a boolean value of the same meaning as described above. list is a list of a/v pairs to be added to the reply list. For example, the following function will always deny the authentication, returning an appropriate message to the user:

     (define (decline-auth request-list check-list reply-list)
       (cons #f
             (list
              (cons "Reply-Message"
                    "\r\nSorry, you are not
                     allowed to log in\r\n"))))

As a more constructive example, let’s consider a function that allows the authentication only if a user name is found in its internal database:

(define staff-data
(list
(list "scheme"
(cons
(list (cons "NAS-IP-Address" "127.0.0.1"))
(list (cons "Framed-MTU" "8096")))
(cons
'()
(list (cons "Framed-MTU" "256"))))))
(define (auth req check reply)
(let* ((username (assoc "User-Name" req))
(reqlist (assoc username req))
(reply-list '()))
(if username
(let ((user-data (assoc (cdr username) staff-data)))
(rad-log L_INFO (format #f "~A" user-data))
(if user-data
    (call-with-current-continuation
     (lambda (xx)
       (for-each
        (lambda (pair)
          (cond
           ((avl-match? req (car pair))
            (set! reply-list (avl-merge
                              reply-list
                              (cdr pair)))
            (xx #t))))
        (cdr user-data))
       #f)))))
(cons
#t
reply-list)))

To trigger the invocation of the Scheme authentication function, assign its name to the Scheme-Procedure attribute in the rhs of a corresponding ‘raddb/users’ profile. For example:

DEFAULT Auth-Type = SQL
Scheme-Procedure = "auth"

10.3.3 Accounting with Scheme

The Scheme accounting procedure must be declared as follows:

[Function Template] acct-function-name request-list

Its argument is:

request-list The list of a/v pairs from the incoming request

The function must return a boolean value. The accounting succeeds only if it has returned #t.

Here is an example of a Scheme accounting function. The function dumps the contents of the incoming request to a file:

  (define radius-acct-file "/var/log/acct/radius")
  (define (acct req)
    (call-with-output-file radius-acct-file
      (lambda (port)
        (for-each (lambda (pair)
                    (display (car pair) port)
                    (display "=" port)
                    (display (cdr pair) port)
                    (newline port))
                  req)
        (newline port)))
    #t)

10.3.4 Radius-Specific Functions

[Scheme Function] avl-delete av-list attr Delete from av-list the pairs with attribute attr.

[Scheme Function] avl-merge dst src

Merge src into dst.

[Scheme Function] avl-match? target list

Return #t if all pairs from list are present in target.

[Scheme Function] rad-dict-name->attr name

Return a dictionary entry for the given attribute name or #f if no such name was found in the dictionary. A dictionary entry is a list in the form

[Scheme List] dict-entry name-string attr-number type-number

vendor where the arguments are as follows:

name-string The attribute name

value-number The attribute number

type-number The attribute type

vendor The vendor PEC, if the attribute is a vendor-specific one, or #f otherwise.

[Scheme Function] rad-dict-value->name attr value

Returns the dictionary name of the given value for an integer-type attribute attr, which can be either an attribute number or its dictionary name.

[Scheme Function] rad-dict-name->value attr value

Convert a symbolic attribute value name into its integer representation.

[Scheme Function] rad-dict-pec->vendor pec

Convert a PEC to the vendor name.

[Scheme Function] rad-log-open prio

Open Radius logging to the severity level prio.

[Scheme Function] rad-log-close

Close a Radius logging channel opened by a previous call to rad-log-

open.
[Scheme Function] rad-rewrite-execute-string string

Interpret string as an invocation of a function in Rewrite language and execute it. Return value: return of the corresponding Rewrite call, translated to the Scheme data type.

[Scheme Function] rad-rewrite-execute arglist

Execute a Rewrite language function. (car arglist) is interpreted as a name of the Rewrite function to execute, and (cdr arglist) as a list of arguments to be passed to it. Return value: return of the corresponding Rewrite call, translated to the Scheme data type.

[Scheme Function] rad-openlog ident option facility Scheme interface to the system openlog() call.

[Scheme Function] rad-syslog prio text

Scheme interface to the system syslog() call.

[Scheme Function] rad-closelog

Scheme interface to the system closelog() call.

[Scheme Function] rad-utmp-putent status delay list radutmp file

radwtmp file Write the supplied data into the radutmp file. If radwtmp file is not nil, the constructed entry is also appended to wtmp file. list is:

[Scheme List] utmp-entry user-name orig-name port-id port-type session-id caller-id framed-ip nas-ip proto

user-name The user name

orig-name The original user name from the request

port-id The value of the NAS-Port-Id attribute

port-type A number or character indicating the port type session-id The session id caller-id The value of the Calling-Station-Id attribute from the

request

framed-ip The framed IP assigned to the user nas-ip The nas IP proto A number or character indicating the type of the connec-

tion

11 Utility Programs

11.1 radwho

Radwho displays the list of users currently logged in by the Radius server.

Default output information is made compatible with that of the standard unix finger(1) utility. For each user the following information is displayed: login name, name, connection protocol, nas port, login date, nas name, assigned IP or corresponding network name.

When used with ‘-l’ option, the long output format is used. In this format the following information is output:

‘Login’ Login name of the user

‘SessionID’

Unique session id assigned by the terminal server.

‘Proto’ Connection prototype.

‘Port’ Port number

‘When’ Login date and time

‘From’ Name of the nas that accepted the connection.

‘Location’

Framed IP or the corresponding network name.

‘Caller’ Caller station id ad reported by the nas.

‘Duration’

Duration of the session.

11.1.1 radwho Command Line Options

The following command line options can be used to modify the behavior of the program:

‘-A’

‘--all’ Display the information about logged-out users as well. The logged-out users are shown with ‘Proto’ field set to HUP.

‘-c’
‘--calling-id’

Display the calling station id in the second column. Equivalent

          to ‘--format clid’.
‘-d NAME’
‘--directory NAME’

Set the Radius configuration directory name.

‘-D fmt’
‘--date-format fmt’

Set the date representation. Th fmt is usual strftime(3) format string. It defaults to %a %H:%M, i.e. the abbreviated weekday name according to the current locale, and the hour and the minutes as two-digit decimal numbers.

‘-e STRING’
‘--empty STRING’

Display any empty field as STRING. This is useful when the output of radwho is fed to some analyzing program, as it helps to keep the same number of columns on each line of output.

‘-F’
‘--finger’

Start in fingerd mode. In this mode radwho emulates the behavior of the fingerd(8) utility. Use this option if starting radwho from the ‘/etc/inetd.conf’ line like this^{1}:

     finger stream tcp nowait nobody /usr/sbin/radwho
     radwho -fL

This mode is also enabled by default if radwho notices that its

name (argv[0]) is ‘fingerd’ or ‘in.fingerd’.
‘-H’
‘--no-header’

Don’t display header line.

‘-i’
‘--session-id’

Display session id instead of gecos in the second column. Equivalent to ‘--format sid’.

‘-I’
‘--ip-strip-domain’

Display hostnames without domain part.

‘-u’
‘--local-also’

Display information about local users from the system ‘utmp’ file. May prove useful when running radwho as a finger daemon.

‘-n’
‘--no-resolve’

Do not resolve IP.

1 In this example the statement has been split on two lines to fit the page width. It

must occupy a single line in the real configuration file.

‘-o format’
‘--format format’

Select customized output format. This can also be changed by setting the value of environment variable RADWHO_FORMAT. The format is either a symbolic name of one of the predefined formats or a format specification (see next subsection).

‘-s’
‘--secure’

Run in secure mode. Queries without a user name are rejected.

11.1.2 radwho Format Strings

A format string controls the output of every record from ‘radutmp’. It contains two types of objects: ordinary characters, which are simply copied to the output, and format specifications, each of which causes output of a particular piece of information from the ‘radutmp’ record.

Each format specification starts with an opening brace and ends with a closing brace. The first word after the brace is the name of the format specification. The rest of words are positional arguments followed by keyword arguments. Both are optional. The keyword arguments begin with a colon and must follow the positional arguments.

The full list of format specifications follows.

[Format Spec] newline [count]

Causes the newline character to be output. If the optional count is supplied, that many newlines will be printed

[Format Spec] tab [num]

Advance to the next tabstop in the output stream. If optional num is present, then skip num tabstops. Each tabstop is eight characters long.

The following specifications output particular fields of a ‘radutmp’ record. They all take two positional arguments: width and title.

The first argument, width sets the maximum output length for this specification. If the number of characters actually output is less than the width, they will be padded with whitespace either to the left or to the right, depending on the presence of the :right keyword argument. If the number of characters is greater than width, they will be truncated to fit. If width is not given, the exact data are output as is.

The second argument, title, gives the title of this column for the heading line. By default no title is output.

Every field specification accepts at least two keyword arguments. The keyword :right may be used to request alignment to the right for the data. This keyword is ignored if width is not given.

The keyword :empty followed by a string causes radwho to output that string if the resulting value for this specification would otherwise be empty.

[Format Spec] login width title [:empty repl][:right]

Print the user login name.

[Format Spec] orig-login width title [:empty repl][:right]

Print original login name as supplied with the request.

[Format Spec] gecos width title [:empty repl][:right]

The gecos field from the local ‘/etc/passwd’ corresponding to the login name. If the user does not have a local account, his login name is output.

[Format Spec] nas-port width title [:empty repl][:right]

nas port number

[Format Spec] session-id width title [:empty repl][:right]

The session ID.

[Format Spec] nas-address width title [:empty

repl][:right][:nodomain] The nas name or IP. The :nodomain keyword suppresses the output of the domain part of the name, i.e., the hostname is displayed only up to the first dot.

[Format Spec] framed-address width title [:empty

repl][:right][:nodomain] Framed IP assigned to the user, if any. The :nodomain keyword suppresses the output of the domain part of the name, i.e. the hostname is displayed only up to the first dot.

[Format Spec] protocol width title [:empty repl][:right]

Connection protocol as reported by Framed-Protocol attribute. If the symbolic value is found in the dictionary file, it will be displayed. Otherwise, the numeric value will be displayed as is.

[Format Spec] time width title [:empty repl][:right][:format

      date-format]

Date and time when the session started. The :format keyword introduces the strftime format string to be used when converting the date for printing. The default value is %a %H:%M.

[Format Spec] duration width title [:empty repl][:right]

Total time of the session duration.

[Format Spec] delay width title [:empty repl][:right]

Delay time (see Section 13.2.2 [Acct-Delay-Time], page 175).

[Format Spec] port-type width title [:empty repl][:right]

Port type as reported by the value of the NAS-Port-Type attribute. If the symbolic value is found in the dictionary file, it will be displayed. Otherwise, the numeric value will be displayed as is.

[Format Spec] clid width title [:empty repl][:right]

The calling station ID.

[Format Spec] realm width title [:empty repl][:right][:nodomain]

If the request was forwarded to a realm server, print the symbolic name of the realm from the ‘raddb/realms’ file. If no symbolic name is found, print the remote server IP or hostname. In the latter case, the :nodomain keyword may be used to suppress the output of the domain part of the name, i.e. to display the hostname only up to the first dot.

11.1.3 radwho Predefined Formats

The predefined formats are:

‘default’ Default output format. Each record occupies one line. The fields output are: login name, gecos name, connection protocol, port number, time when the connection was initiated, nas IP, and assigned framed IP. This corresponds to the following format specification (split in several lines for readability:

     (login 10 Login) (gecos 17 Name) \
     (protocol 5 Proto) (nas-port 5 TTY) \
     (time 9 When) (nas-address 9 From) \
     (framed-address 16 Location)

‘sid’ The same as ‘default’, except that the session ID is output in the second column.

‘clid’ The same as ‘default’, except that the calling station ID is output in the second column.

‘long’ Outputs all information from each ‘radutmp’ record. It is equivalent to specifying the following format string:

     (login 32 Login) (session-id 32 SID) \
     (protocol 5 Proto) (nas-port 5 Port) \
     (time 27 Date) (nas-address 32 NAS) \
     (clid 17 CLID) (duration 7 Duration) \
     (framed-address 16 Location) (realm 16 Realm)

‘gnu’ Each ‘radutmp’ record is represented as a table. It is equivalent to specifying the following format string:

     User: (login)(newline)\
     In real life: (gecos)(newline)\
     Logged in: (time)(newline)\
     NAS: (nas-address)(newline)\
     Port: (nas-port)(newline)\
     CLID: (clid)(newline)\
     Protocol: (protocol)(newline)\
     Session ID: (session-id)(newline)\
     Uptime: (duration)(newline)\
     Assigned IP: (framed-address)(newline)\
     Realm: (realm)(newline)"

11.2 radlast

The radlast utility lists sessions of specified users, nases, nas ports, and hosts, in reverse time order. By default, each line of output contains the login name, the nas short name and port number from where the session was conducted, the host IP or name, the start and stop times for the session, and the duration of the session. If the session is still continuing, radlast will so indicate.

When the ‘-l’ option is specified, radlast produces long output. It includes following fields:

• Login name

• nas short name

• Port number

• Connection protocol

• Port type

• Session id

• Caller id

• Framed IP address

• Session Start Time

• Session Stop Time

• Duration of the Session

11.2.1 radlast Command Line Options

Use following command line options to control the behavior of the radlast utility:

‘-number’
‘-c number’
‘--count number’

When given this option, radlast will output at most this many lines of information.

‘-f’
‘--file name’

Read the specified file instead of the default

          ‘/var/log/radwtmp’.
‘-h hostname’
‘--host hostname’

Report the logins from given host. Host can be either a name or a dotted-quad Internet address.

‘-n shortname’
‘--nas shortname’

Report the logins from the given nas.

‘-l’
‘--long-format’

Long output format. Report all the information stored in

          ‘radwtmp’ file.
‘-p number’
‘--port number’

Report the logins on a given port. The port may be specified either fully or abbreviated, e.g. radlast -p S03 or radlast -p

          3.
‘-s’
‘--show-seconds’

Report the duration of the login session in seconds instead of the default days, hours, and minutes.

‘-t’ The same as ‘-p’. This flag is provided for compatibility with

          last(1).
‘-w’

‘--wide’ Widen the duration field to show seconds as well as the default days, hours and minutes.

If multiple arguments are given, the logical or operation between them is assumed, i.e., the information selected by each argument is printed. This, however, does not apply to the ‘-c’ option. That option is always combined with the rest of command line by logical and.

The pseudo-user ‘~reboot’ logs in on every reboot of the network access server.

If radlast is interrupted, it indicates to what date the search had progressed.

11.3 radzap

radzap searches the Radius accounting database for matching login records and closes them.

At least one of the options ‘-n’, ‘-p’, or the user name must be specified. If they are used in conjunction, they are taken as if joined by the logical and operation.

radzap operates in two modes: silent and confirm. The silent mode is enabled by default. When run in this mode, radzap deletes every record that matches the search conditions given.

In confirm mode radzap will ask for a confirmation before zapping each matching record. Any line beginning with a ‘y’ is taken as a positive response; any other line is taken as a negative response.

The confirm mode is toggled by the command line option ‘-c’.

Syntax

     radzap [options] [username]

Options are:

‘-c’
‘--confirm’

Enable confirm mode.

‘-d dir’
‘--directory dir’

Specify alternate configuration directory. Default is

          ‘/usr/local/etc/raddb’.
‘-f file’
‘--file file’

Operate on file instead of the default ‘RADLOG/radutmp’.

‘-l dir’
‘--log-directory dir’

Search the file ‘radutmp’ in the given directory.

This option is deprecated. It is currently retained for backward compatibility with previous versions.

‘-q’

‘--quiet’ Disable confirm mode.

‘-h’

‘--help’ Display a short help summary, and exit.

‘-n name’
‘--nas name’

Specify nas name to zap user from.

‘-p port’
‘--port port’

Specify the port number of the session to be zapped. The port number can be specified either in its full form, e.g. radzap -p S02, or in its short form, e.g. radzap -p 2.

11.4 radgrep

This utility allows one to quickly look up the user in the Radius accounting database, using a regular expression match. radgrep scans the output of radwho utility and outputs only the lines that match given regular expressions.

Syntax

radgrep accepts two sets of options separated by ‘--’ (double hyphen). The first subset is passed as the command line to the radwho utility. The second one is passed to grep.

11.5 radping

This utility is a shell program that determines the user’s framed IP and runs ping on that address.

Syntax

     radping username
     radping -c calling-station-id

The second way of invoking the program allows one to use the calling station id to indicate the user.

11.6 radauth

The radauth utility sends the Radius server an Access-Request packet and displays the result it gets. If the server responds with Access-Accept radauth can also send an Accounting-Request thereby initiating user’s session.

The utility is a radtest program. See Section 12.2.12 [Sample Radtest Program], page 155, for the detailed discussion of its internals.

Invocation

     radauth [options] [command] user-name [password]

Options are:

‘-v’ Print verbose descriptions of what is being done.

‘-n nas-ip’

Set nas ip address

‘-s sid’ Set accounting session ID

‘-P port’ Set nas port number.

Valid commands are:

auth Send only Access-Request. This is the default.

acct Send Access-Request. If successfull, send Accounting-

          Request with Acct-Status-Type = Start.
start  Send Accounting-Request with Acct-Status-Type = Start.
stop  Accounting-Request with Acct-Status-Type = Stop.

The program determines which Radius server to use, the authentication port number, and the shared secret, following the procedure common to all client scripts (see Section 12.1 [client.conf], page 135).

11.7 radctl

Radctl is a control interface to the radiusd daemon. It allows the user running it to query radiusd about various aspects of its work and to issue administrative commands to it. The syntax is

     radctl command [args]

where command is a command telling radctl which actions to take, and args are optional arguments to the command. Only one command can be specified per invocation.

The valid commands are as follows:

start [args]

If radiusd is not running already, it is started. When present, args are passed as the command line to the server.

stop Stops running radiusd.

restart [args]

Stops the server and then starts it again. When present, args are passed as the command line to the server.

reload Causes the running radiusd server to reread its configuration files.

dumpdb Tells radiusd to dump its user hash table into the file ‘radlog/radius.parse’. This can be used for debugging configuration files.

which Reports the running version of radiusd. This command shows the line of ps(1) describing the running copy of radiusd program. The exact look depends on the version of operating system you are running. Please refer to “man ps” for more detail on ps output. Here is an example of what radctl which prints on GNU/Linux:

                 19692 ?        01:53:11 radiusd

Here, first field is the PID of the process, second field (‘?’) indicates that the running program has detached from the controlling terminal, the third field gives total amount of CPU time used by the program, and, finally, the last field shows the full name under which the command was invoked.

11.8 builddbm

Usage

builddbm converts the plaintext Radius users database into DBM files. Some versions of the Radius daemon have used this to speed up the access to the users database. However, with GNU Radius things go the other way around. The server reads the entire plaintext database, converts it into internal form, and stores into a hash table, which provides for fast access. Actually, using a DBM version of the users database slows down the access unless the machine that runs the Radius daemon is short of address space for the daemon to store the users database.

Syntax

When used without arguments, the builddbm utility attempts to convert the

file ‘raddb/users’ to ‘raddb/users.db’ or to the pair ‘raddb/users.pag’,

‘raddb/users.dir’, depending on the version of the DBM library used.

If used with one argument, that argument is taken as the name of the plaintext database file to operate upon.

Use the following command line options to modify the operation of

buildbm:

‘-d dir’ Specifies alternate directory for the Radius configuration files.

          This defaults to ‘/usr/local/etc/raddb’.

‘-h’ Outputs short usage summary and exits with 0 exit code.

11.9 radscm: A Guile Interface to Radius Functions

radscm is a Scheme interpreter based on Guile with the addition of special functions and variables for communicating with radiusd. This chapter concentrates on the special features provided by radscm. Refer to Guile documentation for information about Scheme and Guile (see section “Overview” in The Guile Reference Manual).

Variables

[Variable] %raddb-path

A path to the Radius configuration directory.

[Function] rad-server-list

A list of radius servers. Each element of the list is:

        (list id-str host-str secret-str auth-num acct-num
              cntl-num)

where the arguments are as follows:

id-str  Server id
host-str  Server hostname or IP
secret-str  Shared secret key to use
auth-num  Authentication port number
acct-num  Accounting port number
cntl-num  Control channel port number

Thus, each entry can be used as an argument to rad-client-set-server

  or rad-client-add-server.

Functions

[Function] rad-send-internal port-number code-number
      pair-list

Sends the request to currently selected server. Arguments are:

port-number Port number to use. These values are allowed:

0  Authentication port
1  Accounting port
2  Control port

The actual port numbers are those configured for the given server.

code-number Request code.

pair-list List of attribute-value pairs. Each pair is either

                         (cons attr-name-str value)

or

                         (cons attr-number value)

Return: On success,

             (list return-code-number pair-list)

On failure,

             '()
[Function] rad-send port-number code-number pair-list .
      verbose

Sends a radius request. Actually it does the same work as rad-send-internal, but if verbose is specified, the verbose report about interaction with the radius server is printed.

[Function] rad-client-list-servers

List currently configured servers. Two columns for each server are displayed: server id and IP.

[Function] rad-get-server

Returns the id of the currently selected server.

[Function] rad-client-set-server list

Selects for use the server described by list. Here list takes the form

     (list id-str host-str secret-str auth-num acct-num
           cntl-num)

where the elements are as follows:

id-str  Server id
host-str  Server hostname or IP
secret-str  Shared secret key to use

auth-num  Authentication port number
acct-num  Accounting port number
cntl-num  Control channel port number

[Function] rad-client-add-server list

Adds the server described by list to the list of active servers. Here list takes the form

     (list id-str host-str secret-str auth-num acct-num
           cntl-num)

where the elements are as follows:

id-str  Server id
host-str  Server hostname or IP
secret-str  Shared secret key to use
auth-num  Authentication port number
acct-num  Accounting port number
cntl-num  Control channel port number

[Function] rad-read-no-echo prompt-str

Prints the given prompt-str, disables echoing, reads a string up to the next newline character, restores echoing, and returns the string entered. This is the interface to the C getpass(3) function.

[Function] rad-client-source-ip ip-str

Sets the IP to be used as source. ip-str can be either an IP in dotted-quad form or a hostname.

[Function] rad-client-timeout number

Sets the timeout in seconds for waiting for a server reply.

[Function] rad-client-retry number

Sets the number of retries for sending requests to a Radius server.

[Function] rad-format-code dest-bool code-number

Format a radius reply code into a human-readable form. dest-bool has the same meaning as in format (see section “Formatted Output” in The Guile Reference Manual.)

[Function] rad-format-pair dest-bool pair

Format a radius attribute-value pair for output. dest-bool has the same meaning as in format. pair is either

                     (cons name-str value)

or

                     (cons attr-number value)

where value may be of any type appropriate for the given attribute.

[Function] rad-print-pairs dest-bool pair-list

Output the radius attribute-value pairs from pair-list. dest-bool has the same meaning as in format. pair-list is a list of pairs in the form

                     (cons name-str value)

or

                     (cons attr-number value)

where value may be of any type appropriate for the given attribute. All Reply-Message pairs from the list are concatenated and displayed as one.

[Function] rad-format-reply-msg pair-list . text

Concatenate and print text from all Reply-Message pairs from pair-list. If text is specified, it is printed before the concatenated text.

[Function] rad-list-servers

For each server from rad-server-list, print its id and hostname or IP.

[Function] rad-select-server ID-STR

Select the server identified by id-str as a current server. The server data are looked up in rad-server-list variable.

[Function] rad-add-server id-str

Add the server identified by id-str to the list of current servers. The server data are looked up in rad-server-list variable.

12 Client Package

Beside the Radius server and accompanying utilities, GNU Radius provides a set of utilities to be used as Radius clients.

The following sections describe in detail the parts of the Radius client package.

12.1 Client Configuration

All programs from the client package share the same configuration file: ‘raddb/client.conf’. The file uses simple line-oriented syntax. Empty lines are ignored; the command ‘#’ introduces an end-of-line comment.

The source IP is introduced with the source_ip statement. Its syntax is:

     source_ip ip-addr

where ip-addr must be the IP in dotted-quad notation.

The Radius server to send the requests to is introduced with server statement:

     server name ip-addr secret auth-port acct-port

Its parts are:

name The server name. It is reserved for further use.

ip-addr The server IP.

secret The shared secret to be used when sending requests to this server.

auth-port The authentication port number.

acct-port The accounting port number.

If several server statement are present, they are tried in turn until one of them replies to the request.

The amount of time a client program waits for the reply from a server is configured using the timeout statement:

     timeout number

If the program does not receive any response within number seconds, it assumes the server does not respond and either retries the transmission or tries the next available server. The number of retries is set with the retry statement:

     retry number

The example ‘raddb/client.conf’ follows:

     server first 10.11.10.1 secret 1645 1646
     server second 10.11.10.1 secret 1645 1646
     source_ip 127.0.0.1
     timeout 3
     retry 5

12.2 radtest

Radtest is a radius client shell, providing a simple and convenient language for sending requests to radius servers and analyzing their reply packets.

12.2.1 Invoking radtest

(This message will disappear, once this node revised.)

‘-a variable=value’
‘--assign=variable=value’

Assign a value to variable. See Section 12.2.4.5 [Assignment Options], page 142, for a detailed discussion.

‘-f file’
‘--file=file’

Read input from file. Stops further processing of the command line.

‘-i’
‘--no-interactive’

Disable interactive mode.

‘-n’
‘--dry-run’

Check the input file syntax and exit.

‘-q’

‘--quick’ Do not read the configuration file.

‘-r number’
‘--retry=number’

Set number of retries.

‘-s server’
‘--server=server’

Set radius server parameters.

‘-t number’
‘--timeout=number’

Set timeout

‘-v’
‘--verbose’

Verbose mode

‘-x debugspec’
‘--debug=debugspec’

Set debugging level

‘-d dir’
‘--directory dir’

Specify alternate configuration directory. Default is

          ‘/usr/local/etc/raddb’.

‘-L’
‘--license’

Print license and exit.

‘-?’

‘--help’ Print short usage summary ‘--usage’ Print even shorter usage summary.

‘-V’
‘--version’

Print program version.

12.2.2 Literal Values

There are four basic data types in radtest language: integer, ipaddr,

string and avlist.

12.2.2.1 Numeric Values

Integer means a signed integer value in the range -2147483648..2147483647.

Ipaddr is an unsigned integer value suitable for representing ipv4 addresses. These can be input either as decimal numbers or as ip addresss in usual “dotted-quad” notation.

As a convenience measure, radius request code names can be used in integer context. The following table lists currently defined request names with their integer codes:

Access-Request  1
Access-Accept  2
Access-Reject  3
Accounting-Request  4
Accounting-Response  5
Accounting-Status  6
Password-Request  7
Password-Ack  8
Password-Reject  9
Accounting-Message  10
Access-Challenge  11
Status-Server  12
Status-Client  13
Ascend-Terminate-Session  31
Ascend-Event-Request  33
Ascend-Event-Response  34
Ascend-Allocate-IP  51
Ascend-Release-IP  52

12.2.2.2 Character Strings

String is an arbitrary string of characters. Any input token consisting of letters of Latin alphabet, decimal digits, underscores dashes and dots and starting with a Latin alphabet letter or underscores is considered a string. To input strings containing other letters, surround them by double quotes. The following are valid strings:

     A-string
     "String, containing white space"

The double quote character ‘"’ must be preceeded by a backslash ‘\’ if it is part of a string:

     "Always quote \" character"

Generally speaking, ‘\’ is an escape character, that alters the meaning of the immediately following character. If it is located at the end of the line, it allows to input newline character to strings:

     "This string contains a \
     newline character."

Other special escape sequences are:

\a Audible bell character (ASCII 7)

\b Backspace (ASCII 8)

\e Escape character (ASCII 27)

\f Form feed (ASCII 12)

\n Newline (ASCII 10)

\r Carriage return (ASCII 13)

\t Horizontal tab (ASCII 9)

\\ Backslash

\ooo (‘o’ represents an octal digit) A character whose ASCII value is represented by the octal number ‘ooo’.

\xHH \XHH (‘H’ represents a hex digit) A character whose ASCII value is

represented by the hex number ‘HH’. If the character following the backslash is not one of those specified, the

backslash is ignored.

An important variant of string is a numeric string, or STRNUM for short. A numeric string is a string that can be converted to a number, for example "+2". This concept is used for type conversion between integer and string values.

Another way to represent strings is using here document syntax. Its format is as follows:

     <<[-]delimiter
       text
     delimiter

Delimiter is any word you choose to delimit the text, text represent the text of the string. If delimiter is prepended by a dash, any leading tabulation characters will be removed from text. This allows for natural indentation of ‘here document’ constructs.

The ‘here document’ construct is especially useful to represent strings containing embedded newlines, as shown in the example below:

     print <<EOT
     usage: foo [OPTIONS] [NAME...]
     OPTIONS are:
       -h            Print this help list.
     EOT

12.2.2.3 Lists of A/V pairs

Avlist are whitespace or comma-separated lists of radius attribute-value pairs. A syntax for a/v pair is

    name op value

where name is attribute name, op is a comparison operator (‘=’, ‘!=’, ‘<’, ‘<=’, ‘>’, ‘>=’), and value is any valid radtest data or expression. An a/v pair list must be enclosed in parentheses. This is an example of an a/v pair list consisting of two pairs:

     ( User-Name = "test" NAS-IP-Address = 10.10.10.1 )

An empty pair list is represented by a pair of parentheses: ().

12.2.3 Reserved Keywords

The following keywords are reserved in radtest:

     acct, and, auth, begin, break, case, continue,
     do, else, end, exit, expect, getopt, if,
     in, input, not, or, print, return, send,
     set, shift, while

The reserved keywords may be used as variable names, provided that the following requrements are met:

• In assignment, these names are quoted using single quotes.

         'case' = 1

• When dereferencing, the use of curly braces is obligatory:

         ${case} + 2

12.2.4 Variables

Variables are means of storing data values at one point of your program for using them in another parts of it. Variables can be assigned either in the program itself, or from the radtest command line.

12.2.4.1 Using Variables

The name of a variable must be a sequence of letters, digits, underscores and dashes, but it may not begin with a digit or dash. Notice, that in contrast to the majority of programming languages, use of dashes (minus signs) is allowed in user names. This is because traditionally radius attribute names contain dashes, so extending this practice to variable names makes radtest programs more consistent. On the other hand, this means that you should be careful when using minus sign as a subtraction operator (see [minusambiguity], page 141). Case is significant in variable names: a and A are different variables.

A name of a variable may coincide with one of radtest reserved keywords. See Section 12.2.3 [Reserved Keywords], page 139, for description on how to use such variables.

A few variables have special built-in meanings (see Section 12.2.4.6 [Builtin Variables], page 143). Such variables can be assigned and accessed just as any other ones. All built-in variables names are entirely upper-case.

Variables are never declared, they spring into existence when an assignment is made to them. The type of a variable is determined by the type of the value assigned to it.

12.2.4.2 Variable Assignments

An assignment stores a new value into a variable. It’s syntax is quite straightforward:

     variable = expression

As a result of the assignment, the expression is evaluated and its value is assigned to variable. If variable did not exist before the assignment, it is created. Otherwise, whatever old value it had before the assignment is forgotten.

It is important to notice that variables do not have permanent types. The type of a variable is the type of whatever value it currently holds. For example:

     foo = 1
     print $foo ⇒ 1
     foo = "bar"
     print $foo ⇒ bar
     foo = ( User-Name = "antonius" NAS-IP-Address = 127.0.0.1 )
     print $foo ⇒ ( User-Name = "antonius" NAS-IP-Address = 127.0.0.1 )

Another important point is that in radtest, assignment is not an expression, as it is in many other programming languages. So C programmers should resist temptation to use assignments in expressions. The following is not correct:

     x = y = 1

Finally, if the variable name coincides with one of radtest keywords, it must be enclosed in single quotes:

     'case' = 1

12.2.4.3 Dereferencing Variables

Dereferencing a variable means accessing its value. The simplest form of dereferencing is by prepending a dollar sign to the variable name:

     foo = 1
     print foo ⇒ foo
     print $foo ⇒ 1

Notice, that in the example above, the first print statement understands foo as a literal string, whereas the second one prints the value of the variable.

Dereferencing an undefined variable produces error message:

     print $x  error variable `x' used before definition

Optionally, the variable name may be surrounded by curly braces. Both $foo and ${foo} are equivalent. The use of the latter form is obligatory only when the variable name coincides with one of the reserved keywords (see Section 12.2.3 [Reserved Keywords], page 139). It also can be used to resolve ambiguity between using dash as a part of user name and as a subtraction operator:

     long-name = 2
     $long-name ⇒ 2
     $long-name-1  error variable `long-name-1' used before definition
     ${long-name}-1 ⇒ 1
     $long-name - 1 ⇒ 1

We recommend to always surround ‘-’ with whitespace when it is used as arithmetic operator.

The ${} notation also permits some operations similar to shell variable substitution.

${variable:-text} Use default values. If variable is unset, return text, otherwise return the value of the variable.

                 $x  error variable `x' used before definition
                 ${x:-1} ⇒ 1
                 x = 2
                 ${x:-1} ⇒ 2

${variable:=text} Assign default values. If variable is unset, text is assigned to it. The expression always returns the value of the variable.

                 $x  error variable `x' used before definition
                 ${x:=1} ⇒ 1
                 $x ⇒ 1

${variable:?text} Display error if unset. If variable is unset, text is written to the standard error (if text is empty, the default diagnostic message is used) and further execution of the program is aborted. Otherwise, the value of variable is returned.

                 $x  error variable `x' used before definition
                 ${x:?}  error x: variable unset
                 ${x:?foobar}  error foobar

${variable::text} Prompt for the value if unset. If variable is unset, radtest prints text (or a default message, if it is empty), reads the standard input up to the newline character and returns the value read. Otherwise, the value of the variable is returned. This notation provides a convenient way for asking user to supply default values.

     ${x::} ⊣ (<teletype>:1)x?
     ${x::Enter value of x: } ⊣ Enter value of x:

${variable:&text} Prompt for the value with echo turned off if unset. This is similar to the ${variable::text}, with the exception that the input value will not be echoed on the screen. This notation provides a convenient way for asking user to supply default values for variables (such as passwords, shared secrets, etc.) while preventing them from being compromised.

12.2.4.4 Accessing Elements of A/V Pair Lists

Elements of an avlist are accessed as if it were an array, i.e.:

$variable [ attribute-name ]

If the attribute attribute-name is of string data type and variable may contain more than one pair with this attribute, adding an asterisk after attribute-name returns concatenated values of all such pairs:

$variable [ attribute-name * ]

Examples:

x = (NAS-Port-Id = 127.0.0.1 \
Reply-Message = "a long"
Reply-Message = " string"
$x[NAS-Port-Id] ⇒ 127.0.0.1
$x[Reply-Message] ⇒ "a long"
$x[Reply-Message*] ⇒ "a long string"

12.2.4.5 Assignment Options

You can set any radtest variable from the command line. There are two ways of doing so.

First, you can use variable assignment option ‘--assign’ (or ‘-a’). Its syntax is:

--assign variable=text
-a variable=text

For example:

radtest -a foobar=5

Another way is useful when you load a radtest program by ‘--file’ or ‘-f’. This second way consists in including a variable assignment in the form

     variable=text

in the command line after the script name. For example:

     radtest -f myprog.rad foo=5 addr=127.0.0.1

This method is especially useful for executable scripts that are run using #! shell magic. Consider a simple script:

     #! /usr/local/bin/radtest -f
     print $addr

The value of addr can be given to the script from the command line as in the example below:

     myprog.rad addr=127.0.0.1

12.2.4.6 Built-in Variables

The following variables are predefined:

(an underscore character)

Contains the result of last evaluated expression.

REPLY CODE Contains the last reply code received from the radius server

          (integer).

REPLY Contains the a/v pairs lastly received from the radius server

          (avlist).

SOURCEIP Contains the source ip address of the radius client (ipaddr). By default, it equals the ip address set via source_ip statement in your ‘client.conf’ file (see Section 12.1 [client.conf], page 135).

INPUT The value of the input read by input statement (see Section 12.2.11 [Built-in Primitives], page 153).

OPTVAR The option obtained by the recent call to getopt (see Section 12.2.11 [Built-in Primitives], page 153).

OPTARG Argument to the option obtained by the recent call to getopt.

OPTIND Index of the next command line argument to be processed by getopt. If the last call to getopt returned false, OPTIND contains index of the first non-optional argument in the command line.

12.2.5 Positional Parameters

Normally radtest stops parsing its command line when it encounters either first non-optional argument (i.e. the one not starting with dash), or an argument consisting of two dashes. The rest of the command line starting from the first non-optional argument forms positional parameters. These parameters are said to form the top-level environment.

Similarly, when invoking a user-defined function (see Section 12.2.7 [Function Definitions], page 149), arguments passed to it are said to form the current environment of the function. These arguments are positional parameters for this function.

Positional parameters are assigned numbers starting from 1. To access (dereference) a positional parameter, the syntax $n is used, where n is the number of the parameter. Alternative forms, such as ${n} or ${n:-text}, can also be used. These work exactly as described in Section 12.2.4.3 [Dereferencing Variables], page 140).

The number of positional parameters can be accessed using a special notation $#.

Several things need to be mentioned:

• All top-level positional parameters have string data type, whereas the types of positional parameters in a function current environment are determined before inoking the function.

• Special notion $0 returns the name of the function being evaluated. When used in the top-level environment, it returns the name of radtest program as given by ‘--file’ (‘-f’) option.

• Dereferencing non-existing parameter returns empty string. This differs from dereferencing non-existing variable, which results in error.

• AWK programmers should note that assignments (see Section 12.2.4.5 [Assignment Options], page 142) are not included in the top level environment (see example below).

For example, suppose you run:

     radtest -f script.rad name foo=bar 5

Then, the top-level environment of program ‘script.rad’ consists of the following variables:

     $0 ⇒ script.rad
     $1 ⇒ name
     $2 ⇒ 5

12.2.6 Expressions

An expression evaluates to a value, which can be printed, assigned to a variable, used in a conditional statement or passed to a function. As in other languages, expressions in radtest include literals, variable and positional parameter dereferences, function calls and combinations of these with various operators.

12.2.6.1 Arithmetic Operations

Radtest provides the common arithmetic operators, which follow normal precedence rules (see Section 12.2.6.8 [Precedence], page 148), and work as you would expect them to. The only notable exception is subtraction operator (minus) which can be used as part of a variable or attribute name, and therefore expressions like $x-3 are ambiguous. This expression can be thought of either as a dereference of the variable x-3 (see Section 12.2.4.3 [Dereferencing Variables], page 140), or as subtraction of the value 3 from the value of the variable x. Radtest always resolves this ambiguityin the favor of variable dereference. Therefore we advise you to always surround minus sign by whitespace, if it is used as a subtraction operator. So, instead of $x-3, write $x - 3. For other methods of solving this ambiguity, See [minus-ambiguity], page 141.

This table lists the arithmetic operators in order from highest precedence to lowest:

- x Negation.

+ x Unary plus. This is equivalent to x.

x * y Multiplication.

x / y Division.

x % y Remainder.

x + y Addition.

x - y Subtraction.

Unary plus and minus have the same precedence, the multiplication, division and remainder all have the same precedence, and addition and subtraction have the same precedence.

If x and y are of different data types, their values are first coerced to a common data type, selected using a set of rules (see Section 12.2.6.6 [Conversion Between Data Types], page 147).

12.2.6.2 String Operations

There is only one string operation: concatenation. It is represented by plus sign, e.g.:

     "string" + "ent" ⇒ "stringent"

12.2.6.3 Operations on A/V Lists

(This message will disappear, once this node revised.)

The following operations are defined on A/V lists:

x + y Addition. The a/v pairs from y are added to x, honoring the respective pairs additivity (see Section 2.1 [additivity], page 7). For example:

                 ( User-Name = "foo" ) + ( Password = "bar" )
                 ⇒ ( User-Name = "foo" Password = "bar" )
                 ( User-Name = "foo" Service-Type = Login-User ) + \
                  ( Service-Type = Framed-User Password = "bar" )
                 ⇒ ( User-Name = "foo" \
                           Service-Type = Framed-User \
                           Password = "bar" )

x - y Subtraction. The result of this operation is an A/V list consist-

ing of pairs from x, which are not found in y.

                 ( User-Name = "foo" Service-Type = Login-User ) - \
                 ( Service-Type = Framed-User )
                 ⇒ ( User-Name = "foo" )

Notice, that only attribute name matters, its value is ignored.

x % y Intersection. The result of this operation is an a/v pair list

consisting of pairs from x which are also present in y.

                 ( User-Name = "foo" Service-Type = Login-User ) - \
                 ( Service-Type = Framed-User )
                 ⇒ ( Service-Type = Login-User )

12.2.6.4 Comparison Operations

Comparison expressions compare operands for relationships such as equality. They return boolean values, i.e. true or false. The comparison operations are nonassociative, i.e. they cannot be used together as in:

     # Wrong!
     1 < $x < 2

Use boolean operations (see Section 12.2.6.5 [Boolean Ops], page 147) to group comparisons together.

Comparison operations can only be used in conditional expressions.

This table lists all comparison operators in order from highest precedence to lowest (notice, however, the comment after it):

x = y True if x is equal to y. C and AWK programmers, please note single equal sign!

x != y True if x is not equal to y.

x < y True if x is less than y.

x <= y True if x is less than or equal to y.

x > y True if x is greater than y.

x >= y True if x is greater than or equal to y.

Operators = and != have equal precedence. Operators <, <=, >, >= have equal precedence.

Most operators are defined for all radtest data types. However, only = and != are defined for avlists. Using any other comparison operator with avlists produces error.

If x and y are of different data types, their values are first coerced to a common data type, selected using a set of rules (see Section 12.2.6.6 [Conversion Between Data Types], page 147).

12.2.6.5 Boolean Operations

A boolean operation is a combination of comparison expressions. Boolean operations can only be used in conditional expressions.

This table lists all comparison operators in order from highest precedence to lowest.

not x

! x True if x is false.

x and y True if both x and y are true. The subexpression y is evaluated only if x is true.

x or y True if at least one of x or y is true. The subexpression y is evaluated only if x is false.

12.2.6.6 Conversion Between Data Types

(This message will disappear, once this node revised.)

The unary negation operand is always converted to integer type:

     -(1 + 1) ⇒ -2
     -(127.0.0.1 + 2) ⇒ -2130706435
     - ("1" + "1") ⇒ -11
     - "text"  error cannot convert string to integer

The unary not operand is converted using the following rules:

1. If the operand is integer, no conversion is performed.

2. If the operand is STRNUM (see [STRNUM], page 138) or ipaddr, it is converted to integer.

3. If the operand is string (but is not STRNUM), the result of not is true only if the operand is an empty string.

4. If the operand is avl, the result of not is true if the list is empty.

Examples:

     not 0 ⇒ 1
     not 10 ⇒ 0
     not "23" ⇒ 0
     not "0" ⇒ 1
     not "text" ⇒ 0
     not "" ⇒ 1
     not 127.0.0.1 ⇒ 0
     not 0.0.0.0 ⇒ 1

When operands of two different data types are used in a binary operation, one of the operands is converted (cast) to another operand’s type according to the following rules:

1. If one of the operands is literal, radtest attemtps to convert another operand to the literal data type. If this attempt fails, it goes on to rule 2. If one of operands is STRNUM (see [STRNUM], page 138) and another

is of numeric data type (i.e. either integer or ipaddr), the latter is converted to string representation.

3. If one of the operands is ipaddr and another is integer, the latter is converted to ipaddr.

4. Otherwise, if one of the operands is string, the second operand is also converted to string.

5. Otherwise, the two operands are incompatible. Radtest prints appropriate diagnostics and aborts execution of the current statement.

12.2.6.7 Function Calls

A function is a name for a particular sequence of statements. It is defined using special definition syntax (see Section 12.2.7 [Function Definitions], page 149). Normally a function return some value. The way to use this value in an expression is with a function call expression, which consists of the function name followed by a comma-separated list of arguments in parentheses. The arguments are expressions which provide values for the function call environment (see Section 12.2.5 [Positional Parameters], page 143. When there is more than one argument, they are separated by commas. If there are no arguments, write just ‘()’ after the function name. Here are some examples:

     foo()             no arguments
     bar(1)            one argument
     bar(1, "string")  two arguments

12.2.6.8 Operator Precedence (How Operators Nest)

Operator precedence determines the order of executing operators, when different operators appear close by in one expression. For example, * has higher precedence than +; thus, a + b * c means to multiply b and c, and then add a to the product.

You can overrule the precedence of the operators by using parentheses. You can think of the precedence rules as saying where the parentheses are assumed to be if you do not write parentheses yourself. Thus the above example is equivalent to a + (b * c).

When operators of equal precedence are used together, the leftmost operator groups first. Thus, a - b + c groups as (a - b) + c.

This table lists radtest operators in order from highest precedence to the lowest:

$ Dereference.

(...) Grouping.

+ - not ! Unary plus, minus. Unary boolean negation.

* / % Multiplication, division, modulus.

+ - Addition, subtraction.

< <= = != > >=

Relational operators.

and Logical ‘and’.

or Logical ‘or’.

12.2.7 Function Definitions

A function is a name for a particular sequence of statements. The syntax for the function definition is:

     name
     begin
       ...
     end

where name is function name and ‘...’ represent a non-empty list of valid radtest statements.

Notice that newline characters are obligatory after name, begin and before the final end keyword.

If the function accepts arguments, these can be referenced in the function body using $n notation (see Section 12.2.5 [Positional Parameters], page 143). To return the value from the function return statement is used.

For example, here is a function that computes sum of the squares of its two arguments:

     hypo
     begin
             return $1*$1 + $2*$2
     end

12.2.8 Interacting with Radius Servers

Radtest provides two commands for interaction with remote radius servers.

Command send sends request to the server specified in

‘raddb/client.conf’. Its syntax is:
     send [flags] port-type code [expr-or-pair-list]

Optional flags can be used for fine-tuning the internals of send. You will seldom need to use these, unless you are developing GNU Radius. See Section 12.2.11 [Built-in Primitives], page 153, for the detailed description of these.

The first obligatory argument, port-type, specifies which radius port to send the request to. Specifying ‘auth’ will send the request to the authentication port (see Section 12.1 [client.conf], page 135); specifying ‘acct’ will send it to the accounting port (see Section 12.1 [client.conf], page 135).

Argument code gives the request code. It is either a number or a symbolic request code name (see Section 12.2.2.1 [Numeric Values], page 137).

The last argument, expr-or-pair-list is either a radtest expression evaluating to avlist or a list of a/v pairs. These pairs will be included in the request.

Here are several examples:

     # Send a Status-Server request without attributes.
     send auth Status-Server
     # Send an Access-Request with two attributes
     send auth Access-Request User-Name = "foo" User-Password = "bar"
     # Send an Accounting-Request, taking attributes from the variable
     # attr
     send acct Accounting-Request $attr

Command send stores the reply code into the variable REPLY_CODE and reply pairs into the variable REPLY (see Section 12.2.4.6 [Built-in Variables], page 143).

Another primitive is expect. Expect takes at most two arguments: a request code (either numeric or symbolic, (see Section 12.2.2.1 [Numeric Values], page 137)) and optional list of a/v pairs (similar to send expr-or-pair-list argument). Expect check if these match current REPLY_CODE and REPLY values and if so, prints the string ‘PASS’. Otherwise, it prints ‘FAIL’. This command is designed primarily for use in GNU Radius testsuite.

Expect is usually used right after send, as shown in the example below:

     send auth Access-Request User-Name = "foo" User-Password = "bar"
     expect Access-Accept Reply-Message = "Access allowed"

12.2.9 Conditional Statements

(This message will disappear, once this node revised.)

Radtest provides two kinds of conditional statements: if and case.

If statement

An if statement in its simplest form is:

     if cond stmt

where cond is a conditional expression and stmt is a valid radtest statement. Optional newline may be inserted between cond stmt.

In this form, if evaluates the condition and if it yields true, executes the statement. For example:

     if $REPLY[NAS-IP-Address] = 127.0.0.1
        print "Request from localhost"

More complex form of this statement allows to select between the two statements:

     if cond stmt-1 else stmt-2

Here, stmt-1 will be executed if cond evaluates to true, and stmt-2 will be executed if cond evaluates to false.

Notice, that an optional newline is allowed between cond and stmt-1 and right after else keyword. However, a newline before else constitutes an error.

If several statements should be executed in a branch of the if statement, use compound statement as in the example below:

     if $REPLY_CODE != Accounting-Response
     begin
       print "Accounting failed.\n"
       exit 1
     end else
       print "Accounting succeeded.\n"

If statements can be nested to any depth.

Case statement

Case statement allows select a statement based on whether a string expression matches given regular expression. The syntax of case statement is:

     case expr in
     expr-1 ) stmt-1
     expr-2 ) stmt-2
     ...
     expr-n ) stmt-n
     end

where expr is a control expression, expr-1, expr-2 etc. are expressions evaluating to extended POSIX regular expressions (for the detailed description of these see section “Regular Expression Library” in Regular Expression Library).

Case statement first evaluates expr and converts it to string data type. Then it evaluates each expr-n in turn and tests if the resulting regular expression matches expr. If so, the statement stmt-n is executed and the execution of case statement finishes.

The following example illustrates the concept:

     case $COMMAND in
     "auth.*")       authenticate($LIST, no)
     "acct")         authenticate($LIST, yes)
     ".*")           begin
                       print "Unknown command."
                       exit 1
                     end
     end

Bourne shell programmers should notice that:

• Case statement ends with end, not esac.

• There is no need to put ;; at the end of each branch,

• Boolean operations are not allowed in expr-n.

12.2.10 Loops

(This message will disappear, once this node revised.)

Two looping constructs are provided: while and do...while.

While loop

The syntax of a while loop is:

     while cond stmt

Newline after cond is obligatory.

Do...while loop

     dostmt while cond

As usual do...while loop differs from its while counterpart in that its stmt is executed at least once.

The looping constructs can be nested to any depth.

Two special statements are provided for branching within loop constructs.

These are break and continue.

Break statement stops the execution of the current loop statement and passes control to the statement immediately following it

     while $x < 10
     begin
       if $x < $y
          break
       ...
       x = $x + 1
     end
     print "OK\n"

In the example above, execution of break statement passes control to print statement.

Break may also take an argument: a literal number representing the number of nested loop statements to break from. For example, the break statement in the sample code below will exit from the outermost while:

     while $y < 10
     begin
       while $x < 10
       begin
         if $x < $y
            break 2
         ...
         x = $x + 1
       end
       ...
       y = $y + 1
     end
     print "OK\n"

Continue statement passes control to the condition of the current looping construct. When used with a numeric argument, the latter specifies the number of the nesting looping construct to pass control to (as with break, the innermost loop is considered to have number 1, so continue is equivalent

to continue 1).

12.2.11 Built-in Primitives

[Radtest built-in] getopt optstring [opt [arg [ind]]]

Getopt is used to break up command line options for subsequent parsing. The only mandatory argument, optstring is a list of short (one-character) options to be recognized. Each short option character in optstring may be followed by one colon to indicate it has a required argument, and by two colons to indicate it has an optional argument. Each subsequent invocation of getopt processes next command line argument. Getopt returns true if the argument is an option and returns false otherwise. It stores the retrieved option (always with a leading dash) in the variable opt (OPTVAR by default). If the option has an argument, the latter is stored in the variable arg (OPTARG by default). Index of the next command line argument to be processed is preserved in the variable ind (OPTIND by default). The usual way of processing command line options is by invoking getopt in a condition expression of while loop and analyzing its return values within the loop. For example:

     while getopt "hf:"
     case $OPTVAR in
     "-h")  print "Got -h option\n"
     "-f")  print "Got -f option. Argument is " $OPTARG "\n"
     ".*")  begin
               print "Unknown option: " $OPTVAR "\n"
               exit 1
            end
       end
     end

[Radtest statement] input [expr name]

Evaluates expr and prints its result on standard output. Then reads a line from standard input and assigns it to the variable name. If expr is given, name must also be present. If name is not given, variable INPUT is used by default.

[Radtest statement] set options

Sets radtest command line options. Options should be a valid radtest command line (see Section 12.2.1 [Invoking radtest], page 136).

[Radtest statement] shift [expr]

Shift positional parameters left by one, so that $2 becomes $1, $3 becomes $2 etc. $# is decremented. $0 is not affected.

If expr is given, it is evaluated, converted to integer and used as shift value. Thus shift 2 shifts all positional parameters left by 2.

[Radtest statement] return [expr]

Returns from the current function (see Section 12.2.7 [Function Definitions], page 149). If expr is present, it is evaluated and the value thus obtained becomes the function return value. It is an error to use return outside of a function definition.

[Radtest statement] break [n]

Exit from within a loop.If n is specified, break from number levels. n must be >= 1. If n is greater than the number of enclosing loops, an error message is issued. See Section 12.2.10 [Loops], page 152, for the detailed discussion of the subject.

[Radtest statement] continue [n]

Resume the next iteration of the enclosing loop. If n is specified, resume at the nth enclosing loop. n must be >= 1. If n is greater than the number of enclosing loops, an error message is issued. See Section 12.2.10 [Loops], page 152, for the detailed discussion of the subject.

[Radtest statement] exit [expr]

Exit to the shell. If expr is specified, it is evaluated and used as exit code. Otherwise, 0 is returned to the shell.

[Radtest statement] print expr-list

Evaluate and print expressions. Expr-list is whitespace or commaseparated list of expressions. Each expression is evaluated in turn and printed to the standard output.

[Radtest statement] send [flags] port-type code
      expr-or-pair-list

Send a request to the radius server and wait for the reply. Stores reply code in the variable REPLY_CODE and reply a/v pairs in the variable REPLY (see Section 12.2.8 [Interacting with Radius Servers], page 149). flags are a whitespace-separated list of variable assignments. Following variables are understood:

repeat=n Unconditionally resend the request n times.

id=n Specify the request ID.

keepauth=1

Do not alter request authenticator when resending the request.

port-type Specifies which port to use when sending the request. Use ‘auth’ to send the request to the authentication port (see Section 12.1 [client.conf], page 135), and ‘acct’ to send it to the accounting port (see Section 12.1 [client.conf], page 135).

code radius request code. Either numeric or symbolic (see Section 12.2.2.1 [Numeric Values], page 137).

expr-or-pair-list Specifies the a/v pairs to include in the request. This argument is either an expression evaluating to avlist, or an immediate avlist (see Section 12.2.2.3 [Avlists], page 139). In the latter case, the parentheses around the list are optional.

[Radtest statement] expect code [expr-or-pair-list]

Test if REPLY_CODE matches code and, optionally, if REPLY matches expror-pair-list. If so, print the string ‘PASS’, otherwise print ‘FAIL’. See Section 12.2.8 [Interacting with Radius Servers], page 149, for the detailed discussion of this statement.

12.2.12 Sample Radtest Program

As an example, let’s consider radauth program (see Section 11.6 [Radauth], page 129). Its main purpose is to send authentication request to the remote server, analyze its reply and if it is positive, send an appropriate accounting record, thereby initiating user’s session. Optionally, the script should also be able to send a lone accounting record.

In the discussion below, we will show and explain subsequent parts of the script text. For the ease of explanation, each line of program text will be prepended by its ordinal line number.

Parsing command line options

The script begins as follows:

       1 #! /usr/bin/radtest -f
       2
       3 while getopt "n:s:P:hv"
       4 begin
       5   case $OPTVAR in
       6   "-n") NASIP = $OPTARG
       7   "-s") SID = $OPTARG
       8   "-P") PID = $OPTARG
       9   "-v") set -v

1 It is a pragmatic comment informing shell that it should run radtest in order to interpret the program.

3 This line starts option processing loop. Getopt (see Section 12.2.11 [Built-in Primitives], page 153) in line 3 analyzes each subsequent command line argument and if it is an option checks whether it matches one of the option letters defined in its first argument. The option letter will be returned in OPTVAR variable, its argument (if any) – in OPTARG variable.

4 – 8 OPTARG value is analyzed using case statement. Lines 6 – 8 preserve OPTARG values in appropriate variables for later use. NASIP will be used as the value of NAS-IP-Address attribute, SID is the session id (Acct-Session-Id attribute), and PID is the port number (for NAS-Port-Id attribute.

9 This line sets ‘-v’ option to the radtest interpreter (see Section 12.2.1 [Invoking radtest], page 136).

The next piece of code handles ‘-h’ and erroneous options:

10   "-h") begin
11           print <<-EOT
12            usage: radauth [OPTIONS] [COMMAND] login [password]
13            Options are:
14            -v         Print verbose descriptions of what is being done
15            -n IP      Set NAS IP address
16            -s SID     Set session ID
17            -P PORT    Set NAS port number
18            COMMAND is one of:
19            auth       Send only Access-Request (default)
20            acct       Send Access-Request. If successfull, send
21                       accounting start request
22            start      Send accounting start request
23            stop       Send accounting stop request
24            EOT
25           exit 0
26         end
27   ".*") begin
28           print "Unknown option: " $OPTVAR "\n"
29           exit 1
30         end
31   end
32 end

10 – 26 Print short description and exit, if the program is given ‘-h’. Notice that ‘here document’ syntax is used to print the text (See Section 12.2.2.2 [Strings], page 137, for its description). The leading whitespace in lines 12 to 24 is composed of tabulation characters (ASCII 9), not usual space characters (ASCII 32), as required by ‘<<-’ construct.

27 – 30 These lines handle unrecognized options.

31 Closes case statement started on line 5

32 Closes compound statement started on line 4

Checking Command Line Consistency

33
34 shift ${OPTIND}-1
35
36 if $# > 3
37 begin
38         print "Wrong number of arguments."
39         print "Try radauth -h for more info"
40         exit 1
41 end

34 OPTIND keeps the ordinal number of the first non-optional argument. This line shifts off all the options processed by getopt, so that the first non-optional argument may be addressed by $1 notation. Notice use of curly braces to solve minus ambiguity (see [minus-ambiguity], page 141).

36 – 41 At this point we may have at most three arguments: command, user name, and password. If there are more, display the diagnostic message and exit the program.

Next piece of code:

42
43 case $1 in
44 "auth|acct|start|stop") begin
45                           COMMAND=$1
46                           shift 1
47                         end
48 ".*")   COMMAND="auth"
49 end
50
51 LOGIN=${1:?User name is not specified. Try radauth -h for more info.}
52
53 if ${NASIP:-} = ""
54         NASIP=$SOURCEIP
55
56 LIST = ( User-Name = $LOGIN NAS-IP-Address = $NASIP )

43 – 48 Check if a command is given. If so, store command name in the variable COMMAND and shift arguments by one, so login becomes argument $1. Otherwise, assume ‘auth’ command.

51 If the user login name is supplied, store it into LOGIN variable. Otherwise, print diagnostic message and exit.

53 – 54 Provide a default value for NASIP variable from the builtin variable SOURCEIP (see Section 12.2.4.6 [Built-in Variables], page 143)

56 The variable LIST will hold the list of A/V pairs to be sent to the server. This line initializes it with a list of two a/v pairs:

User-Name and NAS-IP-Address.

Defining Accounting Function

Accounting function will be used to send accounting requests to the server. It is supposed to take a single argument: an avlist of a/v pairs to be sent to the server.

      57
      58 'acct'
      59 begin
      60   if ${SID:-} = ""
      61     input "Enter session ID: " SID
      62   if ${PID:-} = ""
      63     input "Enter NAS port ID: " PID
      64   send auth Accounting-Request $1 + \
                 (Acct-Session-Id = $SID NAS-Port-Id = $PID)

58 – 59 These lines start the function definition. Notice quoting of the function name (‘acct’): it is necessary because it coincides with a reserved keyword (see Section 12.2.3 [Reserved Keywords], page 139).

60 – 61 If the value of SID (session ID) is not supplied, prompt the user to input it.

62 – 63 If the value of PID (port ID) is not supplied, prompt the user to input it.

64 Send accounting request. The list of a/v pairs to send is

          formed by concatenating Acct-Session-Id and NAS-Port-Id

attributes to the function’s first argument.

The final part of acct function analyzes the reply from the server:

      65   if $REPLY_CODE != Accounting-Response
      66   begin
      67     print "Accounting failed.\n"
      68     exit 1
      69   end
      70   print "Accounting OK.\n"
      71   exit 0
      72 end
      73

Notice, that acct never returns. Instead it exits with an error code indicating success or failure.

Defining Authentication Function

The purpose of the authentication function auth is to send an Access-Request to the server and perform some actions based on its reply.

The function will take three arguments:

$1 The list of a/v pairs to include in the request.

$2 User password.

$3 This argument indicates whether accounting request must be sent after successful authentication. String ‘yes’ means to send the accounting request, ‘no’ means not to send it.

The function is not expected to return. Instead it should exit to the shell with an appropriate error code.

      74 'auth'
      75 begin
      76   send auth Access-Request $1 + (User-Password = $2)

74 – 75 Begin the function definition. Notice quoting of the function name (‘auth’): it is necessary because it coincides with a reserved keyword (see Section 12.2.3 [Reserved Keywords], page 139).

76 Send the initial authentication request. The list of a/v pairs is formed by appending User-Password pair to the list given by the first argument to the function.

The rest of the function analyzes the reply from the server and takes appropriate actions. Notice that if the server replies with an Access-Challenge packet, we will have to send subsequent authentication requests, so this piece of code is enclosed within a while loop.

First, the function handles Access-Accept and Access-Reject replies:

      77   while 1
      78   begin
      79     if $REPLY_CODE = Access-Accept
      80     begin
      81       print "Authentication passed. " + $REPLY[Reply-Message*] + "\n"
      82       if ${3:-no} = no
      83         exit 0
      84       'acct'($1 + ( Acct-Status-Type = Start ))
      85     end else if $REPLY_CODE = Access-Reject
      86     begin
      87       print "Authentication failed. " + $REPLY[Reply-Message*] + "\n"
      88       break

77 Begin an “endless” while loop. It will eventually be exited either using break, or using exit (see below).

79 – 84 Hanlde Access-Accept replies: 81 Print the reply message. Notice the use of ‘*’ to print all the

instances of Reply-Message attribute from the reply packet (see Section 12.2.4.4 [Accessing Elements of A/V Pair Lists], page 142).

82 – 83 If the third argument is missing or is a string ‘no’, exit indicating success (see Section 12.2.4.3 [Dereferencing Variables], page 140).

84 Otherwise, call acct function to perform accounting. The a/v pairs included in the accounting request are formed by adding Acct-Status-Type attribute to the list given by the first argument to the function.

85 – 88 Handle Access-Reject replies. Print the reply message and break from the loop.

Next piece of code deals with Access-Challenge replies. For simplicity we assume that such replies always carry user menus (See Section 4.13 [menus directory], page 62, for the description of these). So, upon receiving an Access-Challenge we should print out the menu, read the users selection and send back an Access-Request to the server. This part is the only one that actually continues the loop at line 77.

      89     end else if $REPLY_CODE = Access-Challenge
      90     begin
      91       print $REPLY[Reply-Message*]
      92       input
      93       send auth Access-Request \
      94         (User-Name = $LOGIN User-Password = $INPUT \
                  State = $REPLY[State])

91 Print the menu contents carrieb by Reply-Message attributes. There may be several instances of the attribute, hence the use of ‘*’ to concatenate their values together.

92 Read the input from the user. The input will be stored in INPUT variable. See Section 12.2.11 [Built-in Primitives], page 153, for the description of input statement.

93 – 94 Send an Access-Request packet with three attributes. User-Password contains the user reply, State contains the menu state from the server reply packet.

Final part of the function:

      95     end else begin
      96       print "Authentication failed. Reply code " + $REPLY_CODE + "\n"
      97       break
      98     end
      99   end
     100   exit 1
     101 end
     102

95 – 98 Handle unknown reply codes.

99 Closes the while loop started on line 77.

100 Exit to the shell indicating failure. This statement will be reached only if a break is executed either on line 88 or on line 97.

101 Closes function definition started on lines 74 – 75

Final Part of Radauth Program

The final part selects an action based on the user command and executes it. It is equivalent to the main function in a C program:

     103 case ${COMMAND} in
     104 "auth")   'auth'($LIST, ${2:&Password: }, no)
     105 "acct")   'auth'($LIST, ${2:&Password: }, yes)
     106 "start")  'acct'($LIST+(Acct-Status-Type = Start))
     107 "stop")   'acct'($LIST+(Acct-Status-Type = Stop))
     108 ".*")       begin
     109               print "Unknown command. Try radauth -h for more info"
     110               exit 1
     111             end
     112 end
     113
     114 # End of radauth

103 Select an action based on the value of COMMAND variable.

104 – 105 Call auth function. If the second argument is given in the command line, its value is taken as user’s password. Otherwise, the user is prompted for the password with the string ‘Password: ’. The input is read with echo turned off to prevent the password from being compromised (the ‘:&’ construct, see Section 12.2.4.3 [Dereferencing Variables], page 140).

106 – 107 Call acct function for ‘start’ and stop commands.

108 – 111 Handle an unknown command verb.

112 Closes case statement from line 103.

12.3 radsession

radsession is a Guile script that sends authentication and accounting requests to the Radius server. To invoke the script, run

     radsession options action

Possible actions are:

‘--auth’ Send authentication request.

‘--start’ Send accounting start request.

‘--stop’ Send accounting stop request.

Options determine the contents of the request’s pairlist. They are:

‘-l STRING’
‘--login STRING’

Set login name.

‘-p STRING’
‘--passwd STRING’

Set password.

‘-n IP’

‘--nas IP’ Set the value of NAS-IP-Address attribute.

‘-s STRING’
‘--sid STRING’

Set the session id (Acct-Session-Id attribute).

‘-P NUMBER’
‘--port NUMBER’

Set the port number (NAS-Port-Id attribute).

‘-h’

‘--help’ Print a short usage message and exit.

‘-v’
‘--verbose’

Verbosely list the contents of the received reply.

12.4 nas.scm

nas.scm is a Guile program that allows one to convert a GNU/Linux box into a nas. It requires Guile version 1.4 or better and PPP version 2.3.7 or better.

To use it, you will basically need to do the following:

1. Create links:

         ln -s libexec/nas.scm /etc/ppp/ip-up
         ln -s libexec/nas.scm /etc/ppp/ip-down

Here, libexec denotes the location of your libexec directory, where nas.scm is installed. If not overridden at configure time, it defaults to ‘prefix/libexec’. These links assure that ppp will invoke nas.scm when the user’s session starts and ends, thus giving it a possibility to send accounting requests.

2. Configure the file ‘raddb/client.conf’.

3. Edit the file ‘raddb/nas.rc’. The supplied ‘nas.rc’ template is tailored to work in most environments. The only variables you may need to change are nas-log-facility, specifying the syslog facility to be used for logging, and pppd-args, keeping the arguments to be given to ppp.

4. Configure your ‘/etc/inittab’ and getty.

For example, if you use mgetty, then the ‘inittab’ entries for dial-up lines will look like:

         d0:345:respawn:/sbin/mgetty ttyS0 vt100
         d1:345:respawn:/sbin/mgetty ttyS1 vt100
         ...

mgetty’s ‘login.config’ will then contain the following line:

         *       -       -       /usr/local/libexec/nas.scm @

If you use agetty, then the ‘inittab’ will contain (with the long lines split for readability)

d0:345:respawn:/sbin/agetty -mt60 \
-l /usr/local/libexec/nas.scm 38400,19200,9600 \
ttyS0 vt100
d1:345:respawn:/sbin/agetty -mt60 \
-l /usr/local/libexec/nas.scm 38400,19200,9600 \
ttyS1 vt100
...
12.5 pam_radius.so

pam_radius.so is a PAM module for Radius authentication. The module understands the following command line options:

‘audit’ Enable audit information.

‘debug[=level]’

Enable debugging information. The higher level is, the more debugging info is output. When omitted, level defaults to 100. Debugging levels equal to or greater than 10 compromise users’ passwords, so use them sparingly.

‘use_authtok’

Use the authentication token passed from the previous module in the stack.

‘confdir=path’

Look for configuration files in path. The default is

‘$sysconfdir/etc/raddb’.

‘attr:’ This keyword marks the end of command line options. The part of the command line after it is parsed as a whitespace-separated list of a/v pairs to be sent with the request.

‘service_type=type’

This option is retained for compatibility with the 0.96 series of GNU Radius. It is equivalent to

             attr: Service-Type=type

The pam_radius.so module logs its messages under LOG_AUTH syslog facility.

164 GNU Radius Reference Manual

13 Attribute List

The following sections describe the most frequently used Radius attributes. Each attribute is described as follows:

     ATTRIBUTE name value type

        Users:  user-flags
        Hints:  hints-flags
        Huntgroups:  huntgroup-flags

Additivity:  additivity
Proxy propagated:  prop

These values have the following meaning:

name The attribute name.

value The attribute number.

type The attribute type.

user-flags Syntax flags defining in which part of a ‘raddb/users’ entry this attribute may be used. The flags consist of two letters: ‘L’ means the attribute can be used in the lhs, ‘R’ means it can be used in the rhs.

hints-flags Syntax flags defining in which part of a ‘raddb/hints’ entry this attribute may be used.

huntgroup-flags Syntax flags defining in which part of a ‘raddb/huntgroups’ entry this attribute may be used.

additivity The additivity of the attribute determines what happens if a rule attempts to add to the pair list an attribute that is already present in this list. Depending on its value, the actions of the server are:

Append New attribute is appended to the end of the list.

Replace New attribute replaces the old.

Drop New attribute is dropped. The old one remains in the list.

prop Is the attribute propagated back to the nas if the server works in proxy mode?

The entry N/A for any of this fields signifies “not applicable”.

13.1 Authentication Attributes

These are the attributes the nas uses in authentication packets and expects to get back in authentication replies. These can be used in matching rules.

13.1.1 CHAP-Password
ATTRIBUTE CHAP-Password 3 string

Users:  L-
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  No

This attribute indicates the response value provided by a PPP Challenge-Handshake Authentication Protocol (CHAP) user in response to the challenge. It is only used in Access-Request packets.

The CHAP challenge value is found in the CHAP-Challenge attribute (60) if present in the packet, otherwise in the request authenticator field.

13.1.2 Callback-Id
ATTRIBUTE Callback-Id 20 string

Users:  -R
Hints:  --
Huntgroups:  --
Additivity:  Replace
Proxy propagated:  No

This attribute indicates the name of a place to be called, to be interpreted by the nas. It may be used in Access-Accept packets.

13.1.3 Callback-Number
ATTRIBUTE Callback-Number 19 string

Users:  -R
Hints:  --
Huntgroups:  --
Additivity:  Replace
Proxy propagated:  No

This attribute indicates a dialing string to be used for callback. It may be used in Access-Accept packets. It may be used in an Access-Request packet as a hint to the server that a Callback service is desired, but the server is not required to honor the hint.

13.1.4 Called-Station-Id
ATTRIBUTE Called-Station-Id 30 string

Users:  L-
Hints:  -R
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

This attribute allows the nas to send in the Access-Request packet the phone number that the user called, using Dialed Number Identification (DNIS) or similar technology. Note that this may be different from the phone number the call comes in on. It is only used in Access-Request packets.

13.1.5 Calling-Station-Id
ATTRIBUTE Calling-Station-Id 31 string

Users:  L-
Hints:  -R
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

This attribute allows the nas to send in the Access-Request packet the phone number that the call came from, using automatic number identification (ANI) or similar technology. It is only used in Access-Request packets.

13.1.6 Class

ATTRIBUTE Class 25 string

Users:  LR
Hints:  LR
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

This attribute is available to be sent by the server to the client in an Access-Accept and should be sent unmodified by the client to the accounting server as part of the Accounting-Request packet if accounting is supported.

13.1.7 Framed-Compression
ATTRIBUTE Framed-Compression 13 integer

Users:  LR
Hints:  -R
Huntgroups:  LR
Additivity:  Replace
Proxy propagated:  Yes

VALUE      Framed-Compression  None                 0
VALUE      Framed-Compression  Van-Jacobson-TCP-IP  1

This attribute indicates a compression protocol to be used for the link. It may be used in Access-Accept packets. It may be used in an Access-Request packet as a hint to the server that the nas would prefer to use that compression, but the server is not required to honor the hint.

More than one compression protocol attribute may be sent. It is the responsibility of the nas to apply the proper compression protocol to appropriate link traffic.

13.1.8 Framed-IP-Address
ATTRIBUTE Framed-IP-Address 8 ipaddr

Users:  LR
Hints:  -R
Huntgroups:  LR
Additivity:  Replace
Proxy propagated:  No

This attribute indicates the address to be configured for the user. It may be used in Access-Accept packets. It may be used in an Access-Request packet as a hint by the nas to the server that it would prefer that address, but the server is not required to honor the hint.

The value 0xFFFFFFFF (255.255.255.255) indicates that the NAS should allow the user to select an address. The value 0xFFFFFFFE (255.255.255.254) indicates that the nas should select an address for the user (e.g. assigned from a pool of addresses kept by the nas). Other valid values indicate that the nas should use that value as the user’s IP.

When used in a rhs, the value of this attribute can optionally be followed by a plus sign. This usage means that the value of NAS-Port-Id must be added to this IP before replying. For example,

             Framed-IP-Address = 10.10.0.1+
13.1.9 Framed-IP-Netmask
     ATTRIBUTE Framed-IP-Netmask 9 ipaddr

Users:  LR
Hints:  -R
Huntgroups:  LR
Additivity:  Replace
Proxy propagated:  No

This attribute indicates the IP netmask to be configured for the user when the user is a router to a network. It may be used in Access-Accept packets. It may be used in an Access-Request packet as a hint by the nas to the server that it would prefer that netmask, but the server is not required to honor the hint.

13.1.10 Framed-MTU
     ATTRIBUTE Framed-MTU 12 integer

Users:  LR
Hints:  -R
Huntgroups:  -R
Additivity:  Replace
Proxy propagated:  Yes

This attribute indicates the maximum transmission unit to be configured for the user, when it is not negotiated by some other means (such as PPP). It is only used in Access-Accept packets.

13.1.11 Framed-Protocol
     ATTRIBUTE Framed-Protocol 7 integer

Users:  LR
Hints:  -R
Huntgroups:  LR
Additivity:  Replace
Proxy propagated:  Yes

VALUE      Framed-Protocol   PPP                  1
VALUE      Framed-Protocol   SLIP                 2

This attribute indicates the framing to be used for framed access. It may be used in both Access-Request and Access-Accept packets.

13.1.12 Framed-Route
ATTRIBUTE Framed-Route 22 string

Users:  -R
Hints:  --
Huntgroups:  --
Additivity:  Replace
Proxy propagated:  No

This attribute provides routing information to be configured for the user on the nas. It is used in the Access-Accept packet and can appear multiple times.

13.1.13 Framed-Routing
ATTRIBUTE Framed-Routing 10 integer

Users:  -R
Hints:  -R
Huntgroups:  -R
Additivity:  Replace
Proxy propagated:  No

VALUE      Framed-Routing    None                 0
VALUE      Framed-Routing    Broadcast            1
VALUE      Framed-Routing    Listen               2
VALUE      Framed-Routing    Broadcast-Listen     3

This attribute indicates the routing method for the user when the user is a router to a network. It is only used in Access-Accept packets.

13.1.14 Idle-Timeout
ATTRIBUTE Idle-Timeout 28 integer

Users:  -R
Hints:  --
Huntgroups:  --
Additivity:  Replace
Proxy propagated:  Yes

This attribute sets the maximum number of consecutive seconds of idle connection allowed to the user before termination of the session or prompt.

The server may send this attribute to the client in an Access-Accept or Access-Challenge.

13.1.15 NAS-IP-Address
     ATTRIBUTE NAS-IP-Address 4 ipaddr

Users:  L-
Hints:  -R
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

This attribute indicates the identifying IP of the nas which is requesting authentication of the user. It is only used in Access-Request packets. Each Access-Request packet should contain either a NAS-IP-Address or a NAS-Identifier attribute (Section 13.1.16 [NAS-Identifier], page 170).

13.1.16 NAS-Identifier
     ATTRIBUTE NAS-Identifier 32 string

Users:  L-
Hints:  -R
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

This attribute contains a string identifying the nas originating the access request. It is only used in Access-Request packets. Either NAS-IP-Address or NAS-Identifier should be present in an Access-Request packet.

See Section 13.1.15 [NAS-IP-Address], page 170.

13.1.17 NAS-Port-Id
     ATTRIBUTE NAS-Port-Id 5 integer

Users:  LR
Hints:  -R
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

This attribute indicates the physical port number of the nas that is authenticating the user. It is only used in Access-Request packets. Note that here we are using “port” in its sense of a physical connection on the nas, not in the sense of a tcp or udp port number.

Some nases try to encode various information in the NAS-Port-Id attribute value. For example, the max ascend terminal server constructs NAS-Port-Id by concatenating the line type (one digit), the line number (two digits), and the channel number (two digits), thus producing a five-digit port number. In order to normalize such encoded port numbers we recommend using a rewrite function (see Section 4.12 [rewrite file], page 61). A rewrite function for max ascend servers is provided in the distribution.

13.1.18 NAS-Port-Type
ATTRIBUTE NAS-Port-Type 61 integer

Users:  --
Hints:  --
Huntgroups:  --
Additivity:  Append
Proxy propagated:  No

VALUE      NAS-Port-Type     Async                0
VALUE      NAS-Port-Type     Sync                 1
VALUE      NAS-Port-Type     ISDN                 2
VALUE      NAS-Port-Type     ISDN-V120            3
VALUE      NAS-Port-Type     ISDN-V110            4

This attribute indicates the type of the physical port of the nas that is authenticating the user. It can be used instead of or in addition to the NAS-Port-Id (Section 13.1.17 [NAS-Port-Id], page 170) attribute. It is only used in Access-Request packets. Either NAS-Port or NAS-Port-Type or both should be present in an Access-Request packet, if the nas differentiates among its ports.

13.1.19 Reply-Message
ATTRIBUTE Reply-Message 18 string

Users:  -R
Hints:  --
Huntgroups:  --
Additivity:  Append
Proxy propagated:  Yes

This attribute indicates text that may be displayed to the user. When used in an Access-Accept, it is the success message. When used in an Access-Reject, it is the failure message. It may indicate

a dialog message to prompt the user before another Access-Request attempt.

When used in an Access-Challenge, it may indicate a dialog message to prompt the user for a response.

Multiple Reply-Message attributes may be included, and if any are displayed, they must be displayed in the same order as they appear in in the packet.

13.1.20 Service-Type
ATTRIBUTE Service-Type 6 integer

Users:  LR
Hints:  -R
Huntgroups:  LR
Additivity:  Replace
Proxy propagated:  Yes

VALUE      Service-Type      Login-User           1
VALUE      Service-Type      Framed-User          2

VALUE      Service-Type      Callback-Login-User  3
VALUE      Service-Type      Callback-Framed-User 4
VALUE      Service-Type      Outbound-User        5
VALUE      Service-Type      Administrative-User  6
VALUE      Service-Type      NAS-Prompt-User      7
VALUE      Service-Type      Authenticate-Only    8
VALUE      Service-Type      Call-Check           10

This attribute indicates the type of service the user has requested, or the type of service to be provided. It may be used in both Access-Request and Access-Accept packets.

When used in an Access-Request the service type represents a hint to the Radius server that the nas has reason to believe the user would prefer the kind of service indicated.

When used in an Access-Accept, the service type is an indication to the nas that the user must be provided this type of service.

The meaning of various service types is as follows:

Login-User

The user should be connected to a host.

Framed-User

A framed protocol, such as PPP or SLIP, should be started for the user. The Framed-IP-Address attribute (see Section 13.1.8 [Framed-IP-Address], page 167) will supply the IP to be used.

Callback-Login-User

The user should be disconnected and called back, then connected to a host.

Callback-Framed-User

The user should be disconnected and called back; then a framed protocol, such as PPP or SLIP, should be started for the user.

Outbound-User

The user should be granted access to outgoing devices.

Administrative-User

The user should be granted access to the administrative interface to the nas, from which privileged commands can be executed.

NAS-Prompt

The user should be provided a command prompt on the nas, from which nonprivileged commands can be executed.

Authenticate-Only

Only authentication is requested, and no authorization information needs to be returned in the Access-Accept.

Call-Check
Callback-NAS-Prompt

The user should be disconnected and called back, then provided a command prompt on the nas, from which nonprivileged commands can be executed.

13.1.21 Session-Timeout
ATTRIBUTE Session-Timeout 27 integer

Users:  -R
Hints:  --
Huntgroups:  --
Additivity:  Replace
Proxy propagated:  Yes

This attribute sets the maximum number of seconds of service to be provided to the user before termination of the session or prompt. The server may send this attribute to the client in an Access-Accept or Access-Challenge.

13.1.22 State

ATTRIBUTE State 24 string

Users:  LR
Hints:  LR
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

This attribute is available to be sent by the server to the client in an Access-Challenge and must be sent unmodified from the client to the server in the new Access-Request reply to that challenge, if any.

This attribute is available to be sent by the server to the client in an Access-Accept that also includes a Termination-Action attribute with the value RADIUS-Request. If the nas performs the termination action by sending a new Access-Request upon termination of the current session, it must include the State attribute unchanged in that Access-Request.

In either usage, no interpretation by the client should be made. A packet may have only one State attribute.

13.1.23 Termination-Action
ATTRIBUTE Termination-Action 29 integer

Users:  LR
Hints:  -R
Huntgroups:  -R
Additivity:  Replace
Proxy propagated:  No

VALUE      Termination-Action  Default              0
VALUE      Termination-Action  RADIUS-Request       1

This attribute indicates what action the nas should take when the specified service is completed. It is only used in Access-Accept packets.

13.1.24 User-Name
     ATTRIBUTE User-Name 1 string

Users:  LR
Hints:  -R
Huntgroups:  LR
Additivity:  Replace
Proxy propagated:  Yes

This attribute indicates the name of the user to be authenticated or accounted. It is used in Access-Request and Accounting attributes. The length of the user name is usually limited by some arbitrary value. By default, Radius supports user names up to 32 characters long. This value can be modified by redefining the RUT_USERNAME macro in the ‘include/radutmp.h’ file in the distribution directory and recompiling the program.

Some nases have peculiarities about sending long user names. For example, the Specialix Jetstream 8500 24-port access server inserts a ‘/’ character after the 10th character if the user name is longer than 10 characters. In such cases, we recommend applying rewrite functions in order to bring the user name to its normal form (see Section 4.12 [rewrite file], page 61).

13.1.25 User-Password
     ATTRIBUTE User-Password 2 string

Users:  L-
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  No

This attribute indicates the password of the user to be authenticated, or the user’s input following an Access-Challenge. It is only used in Access-Request packets.

On transmission, the password is hidden. The password is first padded at the end with nulls to a multiple of 16 octets. A one-way MD5 hash is calculated over a stream of octets consisting of the shared secret followed by the request authenticator. This value is xored with the first 16 octet segment of the password and placed in the first 16 octets of the String field of the User-Password attribute.

If the password is longer than 16 characters, a second one-way MD5 hash is calculated over a stream of octets consisting of the shared secret followed by the result of the first xor. That hash is xored with the second 16 octet segment of the password and placed in the second 16 octets of the string field of the User-Password attribute.

If necessary, this operation is repeated, with each xor result being used along with the shared secret to generate the next hash to xor the next segment of the password, up to no more than 128 characters.

13.1.26 Vendor-Specific

(This message will disappear, once this node revised.)

ATTRIBUTE Vendor-Specific 26 string

Users:  LR
Hints:  -R
Huntgroups:  -R
Additivity:  Append
Proxy propagated:  No

This attribute is available to allow vendors to support their own extended attributes not suitable for general usage.

13.2 Accounting Attributes

These are attributes the nas sends along with accounting requests. These attributes can not be used in matching rules.

13.2.1 Acct-Authentic
ATTRIBUTE Acct-Authentic 45 integer

Users:  --
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  N/A

VALUE           Acct-Authentic          RADIUS          1
VALUE           Acct-Authentic          Local           2
VALUE           Acct-Authentic          Remote          3

This attribute may be included in an Accounting-Request to indicate how the user was authenticated, whether by Radius, the nas itself, or another remote authentication protocol. Users who are delivered service without being authenticated should not generate accounting records.

13.2.2 Acct-Delay-Time
ATTRIBUTE Acct-Delay-Time 41 integer

Users:  --
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  N/A

This attribute indicates how many seconds the client has been trying to send this record for, and can be subtracted from the time of arrival on the server to find the approximate time of the event generating this Accounting-Request. (Network transit time is ignored.)

13.2.3 Acct-Input-Octets
ATTRIBUTE Acct-Input-Octets 42 integer

Users:  --
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  N/A

This attribute indicates how many octets have been received from the port over the course of this service being provided, and can only be present in Accounting-Request records where Acct-Status-Type is set to Stop.

13.2.4 Acct-Input-Packets
ATTRIBUTE Acct-Input-Packets 47 integer

Users:  --
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  N/A

This attribute indicates how many packets have been received from the port over the course of this service being provided to a framed user, and can only be present in Accounting-Request records where Acct-Status-Type is set to Stop.

13.2.5 Acct-Output-Octets
ATTRIBUTE Acct-Output-Octets 43 integer

Users:  --
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  N/A

This attribute indicates how many octets have been sent to the port in the course of delivering this service, and can only be present in Accounting-Request records where Acct-Status-Type is set to Stop.

13.2.6 Acct-Output-Packets
ATTRIBUTE Acct-Output-Packets 48 integer

Users:  --
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  N/A

This attribute indicates how many packets have been sent to the port in the course of delivering this service to a framed user, and can only be present in Accounting-Request records where Acct-Status-Type is set to Stop.

13.2.7 Acct-Session-Id
ATTRIBUTE Acct-Session-Id 44 string

Users:  --
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  N/A

This attribute is a unique accounting id to make it easy to match start and stop records in a log file. The start and stop records for a given session must have the same Acct-Session-Id. An Accounting-Request packet must have an Acct-Session-Id. An Access-Request packet may have an Acct-Session-Id; if it does, then the nas must use the same Acct-Session-Id in the Accounting-Request packets for that session.

13.2.8 Acct-Session-Time
ATTRIBUTE Acct-Session-Time 46 integer

Users:  --
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  N/A

This attribute indicates how many seconds the user has received service for, and can only be present in Accounting-Request records where Acct-

Status-Type is set to Stop.
13.2.9 Acct-Status-Type
ATTRIBUTE Acct-Status-Type 40 integer

Users:  --
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  N/A

VALUE    Acct-Status-Type    Start              1
VALUE    Acct-Status-Type    Stop               2
VALUE    Acct-Status-Type    Alive              3
VALUE    Acct-Status-Type    Accounting-On      7
VALUE    Acct-Status-Type    Accounting-Off     8

This attribute indicates whether this Accounting-Request marks the beginning of the user service (Start) or the end (Stop).

It may also be used to mark the start of accounting (for example, upon booting) by specifying Accounting-On and to mark the end of accounting (for example, just before a scheduled reboot) by specifying Accounting-Off.

A special value Alive or Interim-Update indicates the packet that contains some additional data to the initial Start record or to the last Alive record.

13.2.10 Acct-Terminate-Cause
ATTRIBUTE Acct-Terminate-Cause 49 integer

Users:  --
Hints:  --
Huntgroups:  --
Additivity:  N/A
Proxy propagated:  N/A

VALUE    Acct-Terminate-Cause    User-Request            1
VALUE    Acct-Terminate-Cause    Lost-Carrier            2
VALUE    Acct-Terminate-Cause    Lost-Service            3
VALUE    Acct-Terminate-Cause    Idle-Timeout            4
VALUE    Acct-Terminate-Cause    Session-Timeout         5
VALUE    Acct-Terminate-Cause    Admin-Reset             6
VALUE    Acct-Terminate-Cause    Admin-Reboot            7
VALUE    Acct-Terminate-Cause    Port-Error              8
VALUE    Acct-Terminate-Cause    NAS-Error               9
VALUE    Acct-Terminate-Cause    NAS-Request             10
VALUE    Acct-Terminate-Cause    NAS-Reboot              11
VALUE    Acct-Terminate-Cause    Port-Unneeded           12
VALUE    Acct-Terminate-Cause    Port-Preempted          13
VALUE    Acct-Terminate-Cause    Port-Suspended          14
VALUE    Acct-Terminate-Cause    Service-Unavailable     15
VALUE    Acct-Terminate-Cause    Callback                16
VALUE    Acct-Terminate-Cause    User-Error              17
VALUE    Acct-Terminate-Cause    Host-Request            18

This attribute indicates how the session was terminated, and can only be present in Accounting-Request records where Acct-Status-Type is set to

Stop.

13.3 Radius Internal Attributes

These are attributes used by GNU Radius during the processing of a request. They are never returned to the nas. Mostly, they are used in matching rules.

13.3.1 Acct-Ext-Program
ATTRIBUTE Acct-Ext-Program 2008 string

Users:  --
Hints:  -R
Huntgroups:  --
Additivity:  Replace

Proxy propagated:  N/A

The Acct-Ext-Program attribute can be used in rhs of an ‘raddb/hints’ to require the execution of an external accounting program or filter. If the attribute value starts with a vertical bar (‘|’), then the attribute specifies the filter program to be used. If it starts with a slash (‘/’), then it is understood as the full pathname and arguments for the external program to be executed. Using any other character as the start of this string results in error.

The command line can reference any attributes from both check and reply pairlists using attribute macros (see Section 4.14 [Macro Substitution], page 64).

Before the execution of the program, radiusd switches to the uid and gid of the user daemon and the group daemon. You can override these de-

faults by setting variables exec-program-user and exec-program-group in

configuration file to proper values (see Section 4.1.1 [The option statement], page 22).

The accounting program must exit with status 0 to indicate a successful accounting.

13.3.2 Acct-Type
     ATTRIBUTE Acct-Type 2003 integer

Users:  L-
Hints:  -R
Huntgroups:  -R
Additivity:  Append
Proxy propagated:  N/A

     VALUE           Acct-Type               None    0
     VALUE           Acct-Type               System  1
     VALUE           Acct-Type               Detail  2
     VALUE           Acct-Type               SQL     3

The Acct-Type allows one to control which accounting methods must be used for a given user or group of users. In the absence of this attribute, all currently enabled accounting types are used. See Chapter 7 [Accounting], page 81, for more information about accounting types.

13.3.3 Auth-Failure-Trigger

This attribute specifies an external program or a Scheme expression to be run upon an authentication failure. The handling of this attribute depends upon its value:

If the value of Auth-Failure-Trigger begins with ‘/’, it is taken to contain a command line for invoking an external program. In this case radiusd invokes the program much the same way it does when handling Exec-Program attribute, i.e. the program is invoked with standard input closed, its standard output and standard error are captured and redirected to ‘radlog/radius.stderr’ file, the return value of the program is ignored.

If the value of Auth-Failure-Trigger begins with ‘(’, it is executed it as a Scheme expression. The return value of the expression is ignored.

This attribute is designed as a means to provide special handling for authentication failures. It can be used, for example, to increase failure counters and to block accounts after a specified number of authentication failures occurs. See Section 6.10 [Auth Probing], page 76, for the detailed discussion of its usage.

13.3.4 Auth-Data
     ATTRIBUTE Auth-Data 2006 string

Users:  L-
Hints:  -R
Huntgroups:  -R
Additivity:  Replace
Proxy propagated:  N/A

The Auth-Data can be used to pass additional data to the authentication methods that need them. In version 1.3 of GNU Radius, this attribute may be used in conjunction with the SQL and Pam authentication types. When used with the Pam authentication type, this attribute holds the name of the PAM service to use. This attribute is temporarily appended to the authentication request, so its value can be referenced to as %C{Auth-Data}. See Section 4.11.2 [Authentication Server Parameters], page 54, for an example of of using the Auth-Data attribute in ‘raddb/sqlserver’:

13.3.5 Auth-Type
     ATTRIBUTE Auth-Type 1000 integer

Users:  L-
Hints:  -R
Huntgroups:  -R
Additivity:  Append
Proxy propagated:  No

     VALUE      Auth-Type         Local                0
     VALUE      Auth-Type         System               1
     VALUE      Auth-Type         Crypt-Local          3
     VALUE      Auth-Type         Reject               4
     VALUE      Auth-Type         SQL                  252
     VALUE      Auth-Type         Pam                  253
     VALUE      Auth-Type         Accept               254

This attribute tells the server which type of authentication to apply to a particular user. It can be used in the lhs of the user’s profile (see Chapter 6 [Authentication], page 71.)

Radius interprets values of Auth-Type attribute as follows:

Local The value of the User-Password attribute from the record is taken as a cleantext password and is compared against the User-Password value from the input packet.

System This means that a user’s password is stored in a system password type. Radius queries the operating system to determine if the user name and password supplied in the incoming packet are O.K.

Crypt-Local

The value of the User-Password attribute from the record is taken as an MD5 hash on the user’s password. Radius generates MD5 hash on the supplied User-Password value and compares the two strings.

Reject Authentication fails.

Accept Authentication succeeds.

SQL

Mysql The MD5-encrypted user’s password is queried from the sql database (Section 6.6 [SQL Auth], page 73). Mysql is an alias maintained for compatibility with other versions of Radius.

Pam The user-name–password combination is checked using PAM.

13.3.6 Crypt-Password
     ATTRIBUTE Crypt-Password 1006 string

Users:  L-
Hints:  --
Huntgroups:  --
Additivity:  Append
Proxy propagated:  No

This attribute is intended to be used in user’s profile lhs. It specifies the MD5 hash of the user’s password. When this attribute is present, Auth-

Type = Crypt-Local is assumed. If both Auth-Type and Crypt-Password

are present, the value of Auth-Type is ignored.

See Section 13.3.5 [Auth-Type], page 180.

13.3.7 Exec-Program-Wait
     ATTRIBUTE Exec-Program-Wait 1039 string

Users:  -R
Hints:  --
Huntgroups:  --
Additivity:  Replace
Proxy propagated:  No

When present in the rhs, the Exec-Program-Wait attribute specifies the program to be executed when the entry matches. If the attribute value string starts with vertical bar (‘|’), then the attribute specifies the filter program to be used. If it starts with slash (‘/’), then it is understood as the full pathname and arguments for the external program to be executed. Using any other character as the start of this string results in error.

13.3.7.1 Running an External Program

The command line can reference any attributes from both check and reply pairlists using attribute macros see Section 4.14 [Macro Substitution], page 64.

Before the execution of the program, radiusd switches to uid and gid of the user daemon and the group daemon. You can override these defaults by setting the variable exec-program-user in the configuration file to a proper value. See Section 4.1.1 [The option statement], page 22.

The daemon will wait until the program terminates. The return value of its execution determines whether the entry matches. If the program exits with a nonzero code, then the match fails. If it exits with a zero code, the match succeeds. In this case the standard output of the program is read and parsed as if it were a pairlist. The attributes thus obtained are added to the entry’s reply attributes.

Example.

Suppose the ‘users’ file contains the following entry:

     DEFAULT Auth-Type = System,
                     Simultaneous-Use = 1
             Exec-Program-Wait = "/usr/local/sbin/telauth \
                                  %C{User-Name} \
                                  %C{Calling-Station-Id}"

Then, upon successful matching, the program ‘/usr/local/sbin/telauth’ will be executed. It will get as its arguments the values of the User-Name and Calling-Station-Id attributes from the request pairs.

The ‘/usr/local/sbin/telauth’ can, for example, contain the following:

     #! /bin/sh
     DB=/var/db/userlist
     if grep "$1:$2" $DB; then
         echo "Service-Type = Login,"
         echo "Session-Timeout = 1200"
         exit 0
     else
         echo "Reply-Message = \
               \"You are not authorized to log in\""
         exit 1
     fi

It is assumed that ‘/var/db/userlist’ contains a list of username:callerid pairs for those users that are authorized to use login service.

13.3.7.2 Using an External Filter

If the value of Exec-Program-Wait attribute begins with ‘|’, radiusd strips this character from the value and uses the resulting string as a name of the predefined external filter. Such filter must be declared in ‘raddb/config’ (see Section 4.1.10 [filters], page 38).

Example.

Let the ‘users’ file contain the following entry:

     DEFAULT Auth-Type = System,
                     Simultaneous-Use = 1
             Exec-Program-Wait = "|myfilter"

and let the ‘raddb/config’ contain the following^{1}:

     filters {
         filter myfilter {
             exec-path "/usr/libexec/myfilter";
             error-log "myfilter.log";
             auth {
                 input-format "%C{User-Name}
                               %C{Calling-Station-Id}";
                 wait-reply yes;
             };
         };
     };

Then,  upon  successful  authentication,  the  program

/usr/libexec/myfilter will be invoked, if it hasn’t already been started for this thread. Any output it sends to its standard error will be redirected to the file ‘myfilter.log’ in the current logging directory. A string consisting of the user’s login name and his calling station id followed by a newline will be sent to the program.

The following is a sample /usr/libexec/myfilter written in the shell:

     #! /bin/sh
     DB=/var/db/userlist
     while read NAME CLID
     do
         if grep "$1:$2" $DB; then
             echo "0 Service-Type = Login, Session-Timeout = 1200"
         else
             echo "1 Reply-Message = \
                   \"You are not authorized to log in\""
         fi
     done
13.3.8 Exec-Program
     ATTRIBUTE Exec-Program 1038 string

Users:  -R
Hints:  --

1 In this example the input-format statement has been split on two lines to fit the page

width. It must occupy a single line in the real configuration file.

Huntgroups:  --
Additivity:  Replace
Proxy propagated:  No

When present in the rhs, the Exec-Program attribute specifies the full pathname and arguments for the program to be executed when the entry matches.

The command line can reference any attributes from both check and reply pairlists, using attribute macros (see Section 4.14 [Macro Substitution], page 64).

Before the execution of the program, radiusd switches to the uid and gid of the user daemon and the group daemon. You can override these de-

faults by setting variables exec-program-user and exec-program-group

in configuration file to proper values Section 4.1.1 [The option statement], page 22.

The daemon does not wait for the process to terminate.

Example

Suppose the ‘users’ file contains the following entry:

     DEFAULT Auth-Type = System,
                     Simultaneous-Use = 1
             Exec-Program = "/usr/local/sbin/logauth \
                             %C{User-Name} \
                             %C{Calling-Station-Id}"

Then, upon successful matching, the program ‘/usr/local/sbin/logauth’ will be executed. It will get as its arguments the values of the User-Name and Calling-Station-Id attributes from the request pairs.

13.3.9 Fall-Through
     ATTRIBUTE Fall-Through 1036 integer

Users:  LR
Hints:  LR
Huntgroups:  --
Additivity:  Append
Proxy propagated:  No

     VALUE      Fall-Through      No                   0
     VALUE      Fall-Through      Yes                  1

The Fall-Through attribute should be used in the reply list. If its value is set to Yes in a particular record, that tells Radius to continue looking up other records even when the record at hand matches the request. It can be used to provide default values for several profiles.

Consider the following example. Let’s suppose the ‘users’ file contains the following:

     johns   Auth-Type = SQL
                     Framed-IP-Address = 11.10.10.251,

          Fall-Through = Yes
smith   Auth-Type = SQL
          Framed-IP-Address = 11.10.10.252,
          Fall-Through = Yes
DEFAULT NAS-IP-Address = 11.10.10.1
  Service-Type = Framed-User,
          Framed-Protocol = PPP

Then after successful matching of a particular user’s record, the matching will continue until it finds the DEFAULT entry, which will add its rhs to the reply pairs for this request. The effect is that, if user ‘johns’ authenticates successfully she gets the following reply pairs:

  Service-Type = Framed-User,
  Framed-Protocol = PPP,
  Framed-IP-Address = 11.10.10.251

whereas user smith gets

  Service-Type = Framed-User,
  Framed-Protocol = PPP,
  Framed-IP-Address = 11.10.10.252

Note that the attribute Fall-Through itself is never returned to the nas.

13.3.10 Group

ATTRIBUTE Group 1005 string

Users:  L-
Hints:  L-
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

13.3.11 Hint

ATTRIBUTE Hint 1040 string

Users:  L-
Hints:  -R
Huntgroups:  -R
Additivity:  Append
Proxy propagated:  No

Use the Hint attribute to specify additional matching criteria depending on the hint (see Section 4.6 [hints file], page 49).

Let the ‘hints’ file contain

DEFAULT         Prefix = "S", Strip-User-Name = No
          Hint = "SLIP"

and the ‘users’ file contain

DEFAULT Hint = "SLIP",
          NAS-IP-Address = 11.10.10.12,

                     Auth-Type = System
             Service-Type = Framed-User,
                     Framed-Protocol = SLIP

Then any user having a valid system account and coming from nas ‘11.10.10.12’ will be provided SLIP service if his user name starts with ‘S’.

13.3.12 Huntgroup-Name
     ATTRIBUTE Huntgroup-Name 221 string

Users:  L-
Hints:  -R
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

The Huntgroup-Name can be used either in the lhs of the ‘users’ file record or in the rhs of the ‘huntgroups’ file record.

When encountered in a lhs of a particular ‘users’ profile, this attribute indicates the huntgroup name to be matched. Radius looks up the corresponding record in the ‘huntgroups’ file. If such a record is found, each a/v pair from its reply list is compared against the corresponding pair from the request being processed. The request matches only if it contains all the attributes from the specified huntgroup, and their values satisfy the conditions listed in the huntgroup pairs.

For example, suppose that the authentication request contains the following attributes:

     User-Name = "john",
     User-Password = "guess",
     NAS-IP-Address = 10.11.11.1,
     NAS-Port-Id = 24

Let us further suppose that the ‘users’ file contains the following entry:

     john    Huntgroup-Name = "users_group",
                     Auth-Type = System
             Service-Type = Login

and, finally, ‘huntgroups’ contains the following entry:

     users_group     NAS-IP-Address = 10.11.11.1
                     NAS-Port-Id < 32

Then the authentication request will succeed, since it contains NAS-Port-Id attribute and its value is less than 32.

See Section 4.7 [huntgroups file], page 50.

13.3.13 Log-Mode-Mask
     ATTRIBUTE Log-Mode-Mask 2007 integer

Users:  L-
Hints:  -R
Huntgroups:  -R

Additivity:  Append
Proxy propagated:  N/A

     VALUE           Log-Mode-Mask           Log-Auth                1
     VALUE           Log-Mode-Mask           Log-Auth-Pass           2
     VALUE           Log-Mode-Mask           Log-Failed-Pass         4
     VALUE           Log-Mode-Mask           Log-Pass                6
     VALUE           Log-Mode-Mask           Log-All                 7

Log-Mode-Mask is used to control the verbosity of authentication log messages for given user or class of users. The meaning of its values is:

Log-Auth Do not log successful authentications.

Log-Auth-Pass

Do not show the password with the log message from a successful authentication.

Log-Failed-Pass

Do not show a failed password.

Log-Pass Do not show a plaintext password, either failed or succeeded.

Log-All Do not log authentications at all.

Technical details: After authentication, the server collects all Log-Mode-Mask attributes from the incoming request and lhs of the user’s entry. The values of these attributes ored together form a mask, which is applied via an xor operation to the current log mode. The value thus obtained is used as effective log mode.

13.3.14 Login-Time
     ATTRIBUTE Login-Time 1042 string

Users:  L-
Hints:  --
Huntgroups:  --
Additivity:  Append
Proxy propagated:  No

The Login-Time attribute specifies the time range over which the user is allowed to log in. The attribute should be specified in the lhs.

The format of the Login-Time string is the same as that of UUCP time ranges. The following description of the time range format is adopted from the documentation for the Taylor UUCP package:

A time string may be a list of simple time strings separated with vertical bars ‘|’ or commas ‘,’.

Each simple time string must begin either with a day-of-week abbreviation (one of ‘Su’, ‘Mo’, ‘Tu’, ‘We’, ‘Th’, ‘Fr’, ‘Sa’), or ‘Wk’ for any day from Monday to Friday inclusive, or ‘Any’ or ‘Al’ for any day.

Following the day may be a range of hours separated with a hyphen, using 24-hour time. The range of hours may cross 0; for example ‘2300-0700’ means any time except 7 AM to 11 PM. If no time is given, calls may be made at any time on the specified day(s).

The time string may also be the single word ‘Never’, which does not match any time.

Here are a few sample time strings with an explanation of what they mean.

‘Wk2305-0855,Sa,Su2305-1655’

This means weekdays before 8:55 AM or after 11:05 PM, any time Saturday, or Sunday before 4:55 PM or after 11:05 PM. These are approximately the times during which night rates apply to phone calls in the U.S.A. Note that this time string uses, for example, ‘2305’ rather than ‘2300’; this will ensure a cheap rate even if the computer clock is running up to five minutes ahead of the real time.

‘Wk0905-2255,Su1705-2255’

This means weekdays from 9:05 AM to 10:55 PM, or Sunday from 5:05 PM to 10:55 PM. This is approximately the opposite of the previous example.

‘Any’

This means any day. Since no time is specified, it means any time on any day.

13.3.15 Match-Profile
     ATTRIBUTE Match-Profile 2004 string

Users:  LR
Hints:  -R
Huntgroups:  -R
Additivity:  Append
Proxy propagated:  No

The Match-Profile attribute can be used in lhs and rhs lists of a user profile. Its value is the name of another user’s profile (target profile). When Match-Profile is used in the lhs, the incoming packet will match this profile only if it matches the target profile. In this case the reply pairs will be formed by concatenating the rhs lists from both profiles. When used in the rhs, this attribute causes the reply pairs from the target profile to be appended to the reply from the current profile if the target profile matches the incoming request.

For example:

     IPPOOL  NAS-IP-Address = 10.10.10.1
                     Framed-Protocol = PPP,
                     Framed-IP-Address = "10.10.10.2"
     IPPOOL  NAS-IP-Address = 10.10.11.1
                     Framed-Protocol = PPP,

          Framed-IP-Address = "10.10.11.2"
guest   Auth-Type = SQL
          Service-Type = Framed-User,
  Match-Profile = IPPOOL

In this example, when user guest comes from nas 10.10.10.1, he is assigned IP 10.10.10.2, otherwise if he is coming from nas 10.10.11.1 he is assigned IP 10.10.11.2.

13.3.16 Menu

ATTRIBUTE Menu 1001 string

Users:  -R
Hints:  --
Huntgroups:  --
Additivity:  Replace
Proxy propagated:  No

This attribute should be used in the rhs. If it is used, it should be the only reply item.

The Menu attribute specifies the name of the menu to be presented to the user. The corresponding menu code is looked up in the ‘RADIUS_DIR/menus/’ directory (see Section 4.13 [menus directory], page 62).

13.3.17 Pam-Auth
ATTRIBUTE Pam-Auth 1041 string

Users:  L-
Hints:  -R
Huntgroups:  -R
Additivity:  Append
Proxy propagated:  No

The Pam-Auth attribute can be used in conjunction with

Auth-Type = Pam

to supply the PAM service name instead of the default ‘radius’. It is ignored if Auth-Type attribute is not set to Pam.

13.3.18 Prefix

ATTRIBUTE Prefix 1003 string

Users:  L-
Hints:  L-
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

The Prefix attribute indicates the prefix that the user name should contain in order for a particular record in the profile to be matched. This attribute should be specified in the lhs of the ‘users’ or ‘hints’ file.

For example, if the ‘users’ file contained

DEFAULT Prefix = "U", Auth-Type = System
          Service-Type = Login-User

then the user names ‘Ugray’ and ‘Uyoda’ would match this record, whereas ‘gray’ and ‘yoda’ would not.

Both Prefix and Suffix attributes may be specified in a profile. In this case the record is matched only if the user name contains both the prefix and the suffix specified.

See Section 13.3.27 [Suffix], page 193, and Section 13.3.26 [Strip-User-Name], page 192.

13.3.19 Proxy-Replied
ATTRIBUTE Proxy-Replied 2012 integer

Users:  L-
Hints:  L-
Huntgroups:  L-
Additivity:  Replace
Proxy propagated:  N/A

VALUE      Proxy-Replied     No                   0
VALUE      Proxy-Replied     Yes                  1

radiusd adds this attribute to the incoming request if it was already processed by a remote radius server.

13.3.20 Realm-Name

(This message will disappear, once this node revised.)

ATTRIBUTE Realm-Name 2013 string

Users:  L-
Hints:  L-
Huntgroups:  L-
Additivity:  Append
Proxy propagated:  No

13.3.21 Replace-User-Name
ATTRIBUTE Replace-User-Name 2001 string

Users:  LR
Hints:  LR
Huntgroups:  --
Additivity:  Append
Proxy propagated:  No

VALUE      Replace-User-Name  No                   0
VALUE      Replace-User-Name  Yes                  1

Use this attribute to modify the user name from the incoming packet. The Replace-User-Name can reference any attributes from both lhs and rhs pairlists using attribute macros (Section 4.14 [Macro Substitution], page 64).

For example, the ‘users’ entry

guest   NAS-IP-Address = 11.10.10.11,
          Calling-Station-Id != ""
          Auth-Type = Accept
  Replace-User-Name = "guest#%C{Calling-Station-Id}",
          Service-Type = Framed-User,
          Framed-Protocol = PPP

allows the use of PPP service for user name guest, coming from nas ‘11.10.10.11’ with a nonempty Calling-Station-Id attribute. A string consisting of a ‘#’ character followed by the Calling-Station-Id value is appended to the user name.

13.3.22 Rewrite-Function
ATTRIBUTE Rewrite-Function 2004 string

Users:  LR
Hints:  LR
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

The Rewrite-Function attribute specifies the name of the rewriting function to be applied to the request. The attribute may be specified in either pairlist in the entries of the ‘hints’ or ‘huntgroups’ configuration file.

The corresponding function should be defined in ‘rewrite’ as

integer name()

i.e., it should return an integer value and should not take any arguments.

See Section 4.12 [Packet rewriting rules], page 61, Section 4.6 [hints file], page 49; Section 4.7 [huntgroups file], page 50.

13.3.23 Scheme-Acct-Procedure
ATTRIBUTE Scheme-Acct-Procedure 2010 string

Users:  --
Hints:  -R
Huntgroups:  --
Additivity:  Replace
Proxy propagated:  N/A

The Scheme-Acct-Procedure attribute is used to set the name of the Scheme accounting procedure. See Section 10.3.3 [Accounting with Scheme], page 117, for information about how to write Scheme accounting procedures.

13.3.24 Scheme-Procedure
ATTRIBUTE Scheme-Procedure 2009 string

Users:  -R
Hints:  --
Huntgroups:  --
Additivity:  Append
Proxy propagated:  N/A

The Scheme-Procedure attribute is used to set the name of the Scheme authentication procedure. See Section 10.3.2 [Authentication with Scheme], page 116, for information about how to write Scheme authentication procedures.

13.3.25 Simultaneous-Use
     ATTRIBUTE Simultaneous-Use 1034 integer

Users:  L-
Hints:  -R
Huntgroups:  -R
Additivity:  Append
Proxy propagated:  No

This attribute specifies the maximum number of simultaneous logins a given user is permitted to have. When the user is logged in this number of times, any further attempts to log in are rejected.

See Section 6.9 [Multiple Login Checking], page 74.

13.3.26 Strip-User-Name
     ATTRIBUTE Strip-User-Name 1035 integer

Users:  LR
Hints:  LR
Huntgroups:  -R
Additivity:  Append
Proxy propagated:  No

     VALUE      Strip-User-Name   No                   0
     VALUE      Strip-User-Name   Yes                  1

The value of Strip-User-Name indicates whether Radius should strip any prefixes/suffixes specified in the user’s profile from the user name. When it is set to Yes, the user names will be logged and accounted without any prefixes or suffixes.

A user may have several user names for different kind of services. In this case differentiating the user names by their prefixes and stripping them off before accounting would help keep accounting records consistent.

For example, let’s suppose the ‘users’ file contains

     DEFAULT Suffix = ".ppp",
                     Strip-User-Name = Yes,
                     Auth-Type = SQL
             Service-Type = Framed-User,
                     Framed-Protocol = PPP
     DEFAULT Suffix = ".slip",
                     Strip-User-Name = Yes,
                     Auth-Type = SQL
             Service-Type = Framed-User,
                     Framed-Protocol = SLIP

Now, user ‘johns’, having a valid account in the sql database, logs in as ‘johns.ppp’. She then is provided the PPP service, and her PPP session is accounted under user name ‘johns’. Later on, she logs in as ‘johns.slip’. In this case she is provided the SLIP service and again her session is accounted under her real user name ‘johns’.

13.3.27 Suffix

     ATTRIBUTE Suffix 1004 string

Users:  L-
Hints:  L-
Huntgroups:  LR
Additivity:  Append
Proxy propagated:  No

The Suffix attribute indicates the suffix that the user name should contain in order for a particular record in the profile to be matched. This attribute should be specified in lhs of the ‘users’ or ‘hints’ file.

For example, if the ‘users’ file contained

     DEFAULT Suffix = ".ppp", Auth-Type = System,
                     Strip-User-Name = Yes
             Service-Type = Framed-User,
                     Framed-Protocol = PPP

then the user names ‘gray.ppp’ and ‘yoda.ppp’ would match this record, whereas ‘gray’ and ‘yoda’ would not.

Both Prefix and Suffix attributes may be specified in a profile. In this case the record is matched only if the user name contains both the prefix and the suffix specified.

See Section 13.3.18 [Prefix], page 189, and Section 13.3.26 [Strip-User-Name], page 192.

13.3.28 Termination-Menu
     ATTRIBUTE Termination-Menu 1002 string

Users:  -R
Hints:  --
Huntgroups:  --
Additivity:  Replace
Proxy propagated:  No

This attribute should be used in the rhs. If it is used, it should be the only reply item.

The Termination-Menu specifies the name of the menu file to be presented to the user after finishing his session. The corresponding menu code is looked up in the ‘RADIUS_DIR/menus/’ directory (see Section 4.13 [menus directory], page 62).

194 GNU Radius Reference Manual

14 Reporting Bugs

It is possible you will encounter a bug in one of the Radius programs. If this happens, we would like to hear about it. As the purpose of bug reporting is to improve software, please be sure to include maximum information when reporting a bug. The information needed is:

• Conditions under which the bug appears.

• Version of the package you are using.

• Compilation options used when configuring the package.

• If the bug is found in radiusd daemon, run ‘radiusd -v’ and include the output it produces.

• Contents of Radius configuration directory (‘/usr/local/etc/raddb’ or whatever you have set it to while configuring).

• Log messages produced.

Send your report to bug-gnu-radius@gnu.org. Allow us a couple of days to answer.

196 GNU Radius Reference Manual

15 Where to Get Information about GNU Radius

The two places to look for news regarding GNU Radius are the Radius home-

page at http://www.gnu.org/software/radius and the Radius project
page at http://savannah.gnu.org/projects/radius.

The following mailing lists are related to GNU Radius:

info-gnu-radius@gnu.org

This list distributes announcements and progress reports on GNU Radius. This is a moderated list. Please do not send bug reports or requests for help to this list; there exist special mailing lists for these purposes. To subscribe to the list, visit

http://mail.gnu.org/mailman/listinfo/info-gnu-radius.
help-gnu-radius@gnu.org

This list is the place for users and installers of GNU Radius to ask for help. The list is not moderated, but postings are allowed for list members only. To subscribe to the list, visit

http://mail.gnu.org/mailman/listinfo/help-gnu-radius.
bug-gnu-radius@gnu.org

This list distributes bug reports, bug fixes, and suggestions for improvements in Radius. User discussion of Radius bugs also occurs here. The list is not moderated; postings are allowed for anybody. To subscribe to the list, visit

http://mail.gnu.org/mailman/listinfo/bug-gnu-radius.

198 GNU Radius Reference Manual

How to Obtain Radius

GNU Radius is free software; this means that everyone is free to use it and free to redistribute it on certain conditions. GNU Radius is not in the public domain; it is copyrighted and there are restrictions on its distribution, but these restrictions are designed to permit everything that a good cooperating citizen would want to do. What is not allowed is to try to prevent others from further sharing any version of GNU Radius that they might get from you. The precise conditions are found in the GNU General Public License that comes with Radius and also appears following this section.

One way to get a copy of GNU Radius is from someone else who has it. You need not ask for our permission to do so, or tell any one else; just copy it. If you have access to the Internet, you can get the latest distribution version of GNU Radius by anonymous FTP. It is available at

ftp://ftp.gnu.org/pub/gnu/radius

200 GNU Radius Reference Manual

Radius Glossary

Throughout this document the following terms are used:

radius (small capitals)

The Remote Authentication Dial In User Service protocol as described in RFC 2138, 2865, and 2866.

nas A network access server, that is, a computer or a special device designed to provide access to the network. For example, it can be a computer connected to the network and equipped with several modems. Such a nas will allow a user connecting to one of its modems to access the network.

Service A service, such as PPP, SLIP, or telnet, provided to a user by the nas.

Session Each instance of a service. Sessions start when the service is first provided and close when the service is ended. A user may be allowed to have multiple sessions active simultaneously.

Session id The session identifier: a string of characters uniquely identifying the session.

a/v pair Attribute-value pair: see Section 2.1 [Attributes], page 7.

Dial-in or dial-up user A user connecting to a service through the modem line.

User database A database where a radius server keeps information about users, their authentication information, etc.

User’s profile A record in the user database describing a particular user for purposes of authentication and authorization, i.e., how the user should be authenticated as well as which services he is allowed to be provided and parameters of these services.

202 GNU Radius Reference Manual

Acknowledgements

I would like to acknowledge Oswaldo Aguirre and Francisco Obispo, who invested a lot of time and effort to debug and test the program. They also wrote web-radact — a web interface to the radius database.

Alexandre Oliva provided a lot of good advice and offered valuable help in testing Radius on various platforms.

The following people provided many helpful comments, bug reports and patches: Dustin Mitchell, Jody Owens, Andrey Y. Mosienko, Oleg Gawriloff, Adrian P. van Bloois, Michael Samuel, Michael Smirnov, Andrey Pavlenko, Michael Weiser, Eric Salomé, Clement Gerouville, Dave Restall, Vlad Lungu, Robert Abbate, Jaime Tellez Sanchez, Cornel Cristea, Krzysztof Kopera, and David Friedman.

Additional people need to be thanked for their assistance in producing this manual. Lisa M. Goldstein coordinated its preparation and Joseph C. Fineman and Daniel Barowy did a remarkable job of editing.

And of course, thanks to Richard M. Stallman for founding the FSF and starting the GNU project.

204 GNU Radius Reference Manual

16 New Configuration Approach (draft)

(This message will disappear, once this node revised.)

This document presents a draft describing new approach for processing radius requests. It is intended as a request for comments, and, in the long run, as a guide for GNU Radius developers. In its current state it is far from being complete. Please check http://www.gnu.org/software/ radius/manual for updated versions. Feel free to send your comments and

suggestions to bug-gnu-radius@gnu.org.

16.1 A brief description of Currently Used Approach

When I started to write GNU Radius, back in 1998, I had two major aims. The first and primary aim was to create a flexible and robust system that would follow the principle of Jon Postel:

Be liberal in what you accept and conservative in what you send.

This, I believe, is the main principle of any good software for Internet. The second aim was to be backward compatible with the implementations

that already existed back then. This seemed to be important (and the time has proved it was), because it would allow users to easily switch from older radius daemon to GNU Radius.

An important part of every complex program is its configuration file. Traditional implementations of radius servers (beginning from Livingston Radius) used a configuration suite consisting of several files, usually located in ‘/etc/raddb’ subdirectory. Its main components were:

‘dictionary’

A file containing translations of symbolic names of radius attributes and attribute values to their integer numbers as specified by radius protocol.

‘hints’ This file was intended to separate incoming requests in groups, based on the form of their login name. Traditionally such separation was performed basing on common prefixes and/or suffixes of login names.

‘huntgroups’

The purpose of this file was to separate incoming requests depending on their source, i.e. on the nas that sent them and the port number on that nas. It also served as a sort of simplified access control list.

‘users’ This file contained a users database. It described criteria for authentication and reply pairs to be sent back to requesting nases.

Among these files, the first two were used for requests of any kind, whereas ‘users’ was used only for Access-Request packets.

Though this configuration system suffered from many inconsistencies, the second aim required GNU Radius to use this approach.

To compensate for its deficiencies and to fulfill the first aim, this configuration system was extended, while preserving its main functionality. A number of additional internal attributes were added, that control radiusd behavior. A new language was created whose main purpose was to modify incoming requests (see Section 10.2 [Rewrite], page 98). The support for GNU’s Ubiquitous Intelligent Language for Extensions (see Section 10.3 [Guile], page 115) was added, that allowed to further extend GNU Radius functionality.

The present operation model^{1} of GNU Radius and its configuration file system^{2} emerged as a result of the two development aims described above. Since 1998 up to present, GNU Radius users contributed a lot of ideas and code to the further development of the system.

However, it became obvious that this system presents strong obstacles to the further development. The next section addresses its deficiencies.

16.2 Deficiencies of Current Operation Model and Configuration Suite

The main deficiencies are inherited with the traditional configuration file suite. The rules for processing each request are split among three files, each of which is processed differently, despite of their external similarity. The administrator has to keep in mind a set of exotic rules when configuring the system^{3}. When matching incoming requests with configuration file entries (LHS, see Section 2.3 [Matching Rule], page 11), some attributes are taken verbatim, whereas others are used to control radiusd behavior and to pass additional data to other rules (see Section 13.3 [Radius Internal Attributes], page 178). The things become even more complicated when radius realms come into play (see Section 2.4.2.1 [Proxy Service], page 13). Some attributes are meaningful only if used in a certain part of a certain configuration file rule.

So, while being a lot more flexible than the approach used by other radius implementations, the current system is quite difficult to maintain.

Another deficiency is little control over actions executed on different events. For example, it is often asked how can one block a user account after a predefined number of authentication failures? Currently this can only be done by writing an external authentication procedure (either in Scheme, using Guile, or as a standalone executable, using Exec-Program-Wait). The proper solution would be to have a set of user-defined triggers for every radius event (in this case, for authentication failure).

1 See Chapter 2 [Operation], page 7.

2 See Chapter 4 [Configuration Files], page 21.

3 ‘Hints’ is processed for each request... Authentication requests first pass ‘hints’, then

‘huntgroups’, then ‘users’... Accounting requests use only ‘hints’ and ‘huntgroups’...

‘Huntgroups’ entries may also be used (sometimes inadvertently) to create ACL rules,

etc, etc...

Another commonly asked question is how to make radiusd execute several SQL queries when processing a request. While GNU Radius is not supposed to compensate for deficiencies of some SQL implementations that do not allow for nested queries, such a feature could come quite handy.

16.3 Proposed Solution

(This message will disappear, once this node revised.)

Processing of incoming requests is controlled by request-processing program. Request-processing program is a list-like structure, consisting of instructions.

16.3.1 Request-processing Instruction

Request-processing program consists of instructions. There are seven basic instruction types:

grad_instr_conditional_t

This instruction marks a branch point within the program.

grad_instr_call_t

Represents a call of a subprogram

grad_instr_action_t

Invokes a Rewrite function

grad_instr_proxy_t

Proxies a request to the remote server

grad_instr_forward_t

Forwards a request to the remote server

grad_instr_reply_t

Replies back to the requesting nas.

Consequently, an instruction is defined as a union of the above node types:

[Instruction] grad_instr_t

enum grad_instr_type
{
grad_instr_conditional,
grad_instr_call,
grad_instr_return,
grad_instr_action,
grad_instr_reply,
grad_instr_proxy,
grad_instr_forward
};
typedef struct grad_instr grad_instr_t;
struct grad_instr
{
enum grad_instr_type type;
grad_instr_t *next;
union
  {
    grad_instr_conditional_t cond;
    grad_instr_call_t call;
    grad_instr_action_t action;
    grad_instr_reply_t reply;
    grad_instr_proxy_t proxy;
    grad_instr_forward_t forward;
  } v;
};

Type member contains type of the instruction. The evaluator uses type to determine which part of union v, holds instruction-specific data. Next points to the next instruction. The evaluator will go to this instruction unless the present one changes the control flow. Finally, v contains instruction-specific data. These will be discussed in the following subsections.

16.3.2 grad instr conditional

(This message will disappear, once this node revised.)

[Instruction] grad_instr_conditional_t cond iftrue iffalse

struct grad_instr_conditional
{
grad_entry_point_t cond;  /* Entry point to the compiled
                             Rewrite condition */
grad_instr_t *iftrue;     /* Points to the ``true'' branch  */
grad_instr_t *iffalse;    /* Points to the ``false'' branch  */
};
typedef struct grad_instr_conditional grad_instr_conditional_t;

Instructions of type grad_instr_conditional_t indicate branching. Upon encountering an grad_instr_conditional_t, the engine executes a Rewrite expression pointed to by cond. If the expression evaluates to true, execution branches to instruction iftrue. Otherwise, if iffalse is not NULL, execution branches to that instruction. Otherwise, the control flow passes to grad_instr_t.next, as described in the previous section.

RPL representation

[RPL defun] COND expr if-true [if-false]

expr Textual representation of Rewrite conditional expression or its entry point.

if-true RPL expression executed if expr evaluates to t.

if-true Optional RPL expression that is executed if expr evaluates

             to nil.

Example

COND with two arguments:

     (COND "%[User-Name] ~= \"test-.*\""
           (REPLY Access-Reject ("Reply-Message" . "Test accounts disabled")))

COND with three arguments:

     (COND "%[Hint] == "PPP" && authorize(PAM)"
           (REPLY Access-Accept
                  ("Service-Type" . "Framed-User")
                  ("Framed-Protocol" . "PPP"))
           (REPLY Access-Reject
                  ("Reply-Message" . "Access Denied")))

16.3.3 grad instr call

(This message will disappear, once this node revised.)

[Instruction] grad_instr_call_t entry

        struct grad_instr_call {
               grad_instr_t *entry;
        };
        typedef struct grad_instr_call grad_instr_call_t;

Instructions of type grad_instr_call instruct the engine to call the given subprogram. The engine pushes the current instruction to the return point stack and branches to instruction entry. Execution of the subprogram ends when the engine encounters an instruction of one of the following types:

grad_instr_return, grad_instr_reply or grad_instr_proxy.

If grad_instr_return is encountered, the engine pops the instruction from the top of the return point stack and makes it current instruction, then it branches to the next node.

   If grad_instr_reply or grad_instr_proxy is encountered, the engine,

after executing corresponding actions, finishes executing the program.

RPL representation

[RPL defun] CALL list

[RPL defun] CALL defun-name

In the first form, the argument list is the RPL subprogram to be executed. In the second form defun-name is a name of the RPL subprogram defined

  by defun.

Examples

First form:

     (CALL (ACTION "myfun(%[User-Name])")
           (REPLY Access-Reject
                  ("Reply-Message" . "Access Denied")))

Second form:

     (CALL process_users)

16.3.4 grad instr return

(This message will disappear, once this node revised.)

An instruction of type grad_instr_return indicates a return point from the subprogram. If encountered in a subprogram (i.e. a program entered by grad_instr_call node), it indicates return to the calling subprogram (see the previous subsection). Otherwise, if grad_instr_return is encountered within the main trunk, it ends evaluating of the program.

Instructions of this type have no data associated with them in union v.

RPL representation

[RPL defun] RETURN

Examples

     (RETURN)

16.3.5 grad instr action

(This message will disappear, once this node revised.)

[Instruction] grad_instr_reply_t expr
        struct grad_instr_action {
               grad_entry_point_t expr;    /* Entry point to the compiled
                                              Rewrite expression */
        };
        typedef struct grad_instr_action grad_instr_reply_t;

The machine executes a Rewrite expression with entry point expr. Any return value from the expression is ignored.

RPL representation

[RPL defun] ACTION expr
[RPL defun] ACTION entry-point

Examples

     (ACTION "%[NAS-IP-Address] = request_source_ip()")

16.3.6 grad instr reply

(This message will disappear, once this node revised.)

[Instruction] grad_instr_reply_t return code

        struct grad_instr_reply {
               u_char reply_code;         /* Radius request code */
        };
        typedef struct grad_instr_reply grad_instr_reply_t;

grad_instr_reply instructs radiusd to send to the requesting nas a reply with code reply_code. Any reply pairs collected while executing the program are attached to the reply.

After executing grad_instr_reply instruction, the engine stops executing of the program.

Any execution path will usually end with this instruction.

RPL representation

[RPL defun] REPLY reply-code [attr-list]

Arguments:

reply-code Radius reply code.

attr-list List of a/v pairs to be added to the reply. Each a/v pair is

             represented as a cons: (name-or-number . value).

Example

     (REPLY Access-Accept
            ("Service-Type" . "Framed-User")
            ("Framed-Protocol" . "PPP"))

16.3.7 grad instr proxy

(This message will disappear, once this node revised.)

[Instruction] grad_instr_proxy_t realm
        struct grad_instr_proxy
        {
          grad_realm_t realm;
        };
        typedef struct grad_instr_proxy grad_instr_proxy_t;

This instruction tells radius to proxy the request to the server defined in realm. In other words, the engine executes proxy_send. Further processing of the program is stopped.

RPL representation

[RPL defun] PROXY realm-name

Realm-name is name of the realm as defined in ‘raddb/realms’.

Examples

.

16.3.8 grad instr forward

(This message will disappear, once this node revised.)

[Instruction] grad_instr_forward_t server list

        struct grad_instr_forward
        {
          grad_list_t server_list;
        };
        typedef struct grad_instr_forward grad_instr_forward_t;

This node forwards the request to each servers from server_list. Forwarding differs from proxying in that the requests are sent to the remote servers and processed locally. The remote server is not expected to reply. See Section 4.1.3 [auth], page 28, for more information on this subject.

In contrast to grad_instr_proxy, this instruction type does not cause the execution to stop.

   Elements of server_list are of type grad_server_t.

Currently forwarding is performed by forward_request function (‘forward.c’), which could be used with little modifications. Namely, it will be rewritten to get server list as argument, instead of using static variable forward_list. Consequently, the functions responsible for creating and initializing this static variable will disappear along with the variable itself. .

16.3.9 RPL representation

[RPL defun] FORWARD server-list

16.4 Changes to Rewrite Language

(This message will disappear, once this node revised.)

16.5 Support for Traditional Configuration Files.

(This message will disappear, once this node revised.)

Within the new configuration system, the traditional “trio” ‘hints-huntgroups-users’ will be translated to the following program:

Chapter 16: New Configuration Approach (draft) 213

(defprog main
 (CALL hints)
 (CALL huntgroups)
 (COND "request_code() == Access-Request"
       (CALL users))
 (REPLY Access-Reject
        (Reply-Message . "\nAccess denied\n")))

For example, consider the following configuration:

# raddb/hints:
DEFAULT Prefix = "PPP" Hint = PPP

This will produce the following program:

(defprog hints
 (COND "%[Prefix] == \"PPP\"")
       (ACTION "%[Hint] = \"PPP\""))
#raddb/huntgroups
DEFAULT NAS-IP-Address = 10.10.4.1      Suffix = "staff"
DEFAULT NAS-IP-Address = 10.10.4.2      Huntgroup-Name = "second"

Will produce

(defprog huntgroups
 (COND "%[NAS-IP-Address] == 10.10.4.1 && !(%[Suffix] == \"staff\")"
       (REPLY Access-Reject
              ("Reply-Message" . "Access Denied by Huntgroup")))
 (COND "%[NAS-IP-Address] == 10.10.4.2"
       (ACTION "%[Huntgroup-Name] = \"second\"")))

Finally, ‘users’:

#raddb/users
DEFAULT Hint = "PPP",
               Auth-Type = PAM
        Service-Type = Framed-User,
               Framed-Protocol = PPP
DEFAULT Huntgroup-Name = "second",
               Auth-Type = PAM
        Service-Type = "Authenticate-Only",
               Reply-Message = "Authentity Confirmed"

will produce

(defprog users
 (COND "%[Hint] == "PPP" && authorize(PAM)"
       (REPLY Access-Accept
             (Service-Type . Framed-User)
             (Framed-Protocol . PPP))
       (REPLY Access-Reject
             (Reply-Message . "Access Denied")))
 (COND "%[Huntgroup-Name] == \"second\" && authorize(PAM)"
       (REPLY Access-Accept
              (Service-Type . "Authenticate-Only")
              (Reply-Message . "Authentity Confirmed"))))

214 GNU Radius Reference Manual

16.6 New Configuration Files

(This message will disappear, once this node revised.)

Appendix A GNU Free Documentation License

Version 1.2, November 2002

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000,2001,2002 Free Software Foundation, Inc. 59 Temple Place, Suite 330, Boston, MA 02111-1307, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct

from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified

Version, as the publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adja-

cent to the other copyright notices. F. Include, immediately after the copyright notices, a license notice

giving the public permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections

and required Cover Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add

to it an item stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included in the Modified Version.

N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in title with any Invariant Section.

O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties— for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one.

The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. See

http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

A.0.1 ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

       Copyright (C)  year  your name.
       Permission is granted to copy, distribute and/or modify this document
       under the terms of the GNU Free Documentation License, Version 1.2
       or any later version published by the Free Software Foundation;
       with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts.
       A copy of the license is included in the section entitled ``GNU
       Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with...Texts.” line with this:

         with the Invariant Sections being list their titles, with
         the Front-Cover Texts being list, and with the Back-Cover Texts
         being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Appendix A: Index 223

Index

$

$INCLUDE (dictionary) . . . . . . . . . . . . . . 40

%

%raddb-path . . . . . . . . . . . . . . . . . . . . . . . 131

_. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

A

A/V pair . . . . . . . . . . . . . . . . . . . . . . . . 7, 201 Accept Authentication Type . . . . . . . . . . 71 access-denied . . . . . . . . . . . . . . . . . . . . . . 36 ‘access.deny’ file . . . . . . . . . . . . . . . . . . . . 53 account-closed . . . . . . . . . . . . . . . . . . . . . 36 Accounting directory . . . . . . . . . . . . . . . . . . 5 Accounting requests. . . . . . . . . . . . . . . . . . . 9 Accounting service parameters . . . . . . . . 30 Accounting Types. . . . . . . . . . . . . . . . . . . . 81 Accounting with Scheme . . . . . . . . . . . . 117 acct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 acct statement. . . . . . . . . . . . . . . . . . . . . . 30 Acct-Authentic . . . . . . . . . . . . . . . . . . . . 175 Acct-Delay-Time . . . . . . . . . . . . . . . . . . . 175 acct-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Acct-Ext-Program . . . . . . . . . . . . . . . . . . 178 acct-function-name . . . . . . . . . . . . . . . . 117 Acct-Input-Octets . . . . . . . . . . . . . . . . . 176 Acct-Input-Packets . . . . . . . . . . . . . . . . 176 Acct-Output-Octets . . . . . . . . . . . . . . . . 176

Acct-Output-Packets . . . . . . . . . . . . . . . 176

Acct-Session-Id . . . . . . . . . . . . . . . . . . . 177 Acct-Session-Time . . . . . . . . . . . . . . . . . 177 Acct-Status-Type . . . . . . . . . . . . . . . . . . 177

Acct-Terminate-Cause. . . . . . . . . . . . . . 178

Acct-Type . . . . . . . . . . . . . . . . . . . . . . . . . 179 acl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 ACTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 Additivity of an attribute. . . . . . . . . . . . . . 7 ALIAS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 allow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

Analyzing SNMP output . . . . . . . . . . . . . 104

Attribute . . . . . . . . . . . . . . . . . . . . . . . . 7, 201 ATTRIBUTE. . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Attribute-value pair . . . . . . . . . . . . . . . . . . . 7 Attribute-Value pair . . . . . . . . . . . . . . . . 201 auth. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 38 Auth-Data . . . . . . . . . . . . . . . . . . . . . . . . . 180

Auth-Failure-Trigger. . . . . . . . . . . . . . 179

auth-function . . . . . . . . . . . . . . . . . . . . . 116 Auth-Type . . . . . . . . . . . . . . . . . . . . . . . . . 180 auth_db . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 auth_failure_query . . . . . . . . . . . . . . . . . 54 auth_query . . . . . . . . . . . . . . . . . . . . . . . . . 54 auth_success_query . . . . . . . . . . . . . . . . . 54 Authentication . . . . . . . . . . . . . . . . . . . . . . 16 authentication probes . . . . . . . . . . . . . . . . 76 Authentication requests . . . . . . . . . . . . . . . 8 Authentication service parameters. . . . . 28 Authentication with Scheme . . . . . . . . . 116 avl-delete . . . . . . . . . . . . . . . . . . . . . . . . 118 avl-match? . . . . . . . . . . . . . . . . . . . . . . . . 118 avl-merge . . . . . . . . . . . . . . . . . . . . . . . . . 118

B

BEGIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 BEGIN-VENDOR . . . . . . . . . . . . . . . . . . . . . . . 42 break . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 buildbm. . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 Built-in functions, Rewrite . . . . . . . . . . 111

C

CALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 Callback-Id . . . . . . . . . . . . . . . . . . . . . . . 166 Callback-Number . . . . . . . . . . . . . . . . . . . 166 Called-Station-Id . . . . . . . . . . . . . . . . . 166 Calling-Station-Id . . . . . . . . . . . . . . . . 167 category. . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 channel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 CHAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 CHAP-Password . . . . . . . . . . . . . . . . . . . . . 166 check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

Checking UNIX finger output . . . . . . . 104
checkrad-assume-logged. . . . . . . . . . . . . 28

Class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 clid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 Client Configuration . . . . . . . . . . . . . . . . 135 Client Package. . . . . . . . . . . . . . . . . . . . . . 135 `client.conf' . . . . . . . . . . . . . . . . . . . . . 135 ‘clients’ file . . . . . . . . . . . . . . . . . . . . . . . . 44 `clients' file. . . . . . . . . . . . . . . . . . . . . . 45 common . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 community. . . . . . . . . . . . . . . . . . . . . . . . . . . 32

compare-atribute-flag . . . . . . . . . . 28, 30

Comparing the requests . . . . . . . . . . . . . . 67 COND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 Configuration directory. . . . . . . . . . . . . . . . 5 Configuration files (radiusd) . . . . . . . . . . 21 continue. . . . . . . . . . . . . . . . . . . . . . . . . . . 154 Controlling Authentication Probes …. 76 Crypt-Password . . . . . . . . . . . . . . . . . . . . 181 Custom Accounting Types. . . . . . . . . . . . 83 Custom Authentication Types . . . . . . . . 73 Customizing accounting service . . . . . . . 30 Customizing authentication server. . . . . 28 Customizing Radiusd Guile interface . . 35 Customizing reply messages . . . . . . . . . . 36 Customizing SNMP server . . . . . . . . . . . . 32

D

d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 Data directory . . . . . . . . . . . . . . . . . . . . . . . . 5 Data types, Rewrite. . . . . . . . . . . . . . . . . 105 ‘datadir’, directory for shared data files

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 DBM: enabling . . . . . . . . . . . . . . . . . . . . . . 32 debug . . . . . . . . . . . . . . . . . . . . . . . . . . . 35, 93 Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . 89 Declarations, Rewrite . . . . . . . . . . . . . . . 108 delay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 Deleting hung user sessions . . . . . . . . . . 127 deny . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 detail . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 30 Detailed Request Accounting . . . . . . . . . 81 dgettext. . . . . . . . . . . . . . . . . . . . . . . . . . . 114 Dial-in user . . . . . . . . . . . . . . . . . . . . . . . . 201 Dial-up user . . . . . . . . . . . . . . . . . . . . . . . . 201 dict-entry . . . . . . . . . . . . . . . . . . . . . . . . 118 ‘dictionary’ file . . . . . . . . . . . . . . . . . . . . . 40 Disabling user accounts . . . . . . . . . . . . . . 71 dngettext . . . . . . . . . . . . . . . . . . . . . . . . . 114 doauth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 duplicate requests, checking. . . . . . . . . . . 12 duration. . . . . . . . . . . . . . . . . . . . . . . . . . . 124

E

Enabling DBM . . . . . . . . . . . . . . . . . . . . . . 32 Encrypted Password Authentication Type

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 END . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 END-VENDOR . . . . . . . . . . . . . . . . . . . . . . . . . 42 eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Exec-Program . . . . . . . . . . . . . . . . . . . . . . 183 exec-program-user . . . . . . . . . . . . . . . . . . 22 Exec-Program-Wait . . . . . . . . . . . . . . . . . 181 exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 expect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

                                        Extended comparison . . . . . . . . . . . . . . . . 68

Extended Comparison . . . . . . . . . . . . . . . . 67 Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . 95

F

Fall-Through . . . . . . . . . . . . . . . . . . . . . . 184 FDL, GNU Free Documentation License

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26, 32 filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 filters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Filters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 FORWARD. . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 framed-address . . . . . . . . . . . . . . . . . . . . 124 Framed-Compression . . . . . . . . . . . . . . . . 167 Framed-IP-Address . . . . . . . . . . . . . . . . . 167 Framed-IP-Netmask . . . . . . . . . . . . . . . . . 168 Framed-MTU . . . . . . . . . . . . . . . . . . . . . . . . 168 Framed-Protocol . . . . . . . . . . . . . . . . . . . 168 Framed-Route . . . . . . . . . . . . . . . . . . . . . . 169 Framed-Routing . . . . . . . . . . . . . . . . . . . . 169

G

g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 gc-interval . . . . . . . . . . . . . . . . . . . . . . . . 35 gecos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 getopt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 gettext. . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 Group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 group_query . . . . . . . . . . . . . . . . . . . . . . . . 54 gsub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

                                        Guest accounts, setting up . . . . . . . . . 71

guile . . . . . . . . . . . . . . . . . . . . . . . . . . . 35, 91 Guile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 Guile interface. . . . . . . . . . . . . . . . . . . . . . 131 Guile interface configuration . . . . . . . . . . 35 Guile, representation of Radius data. . 115

H

Hint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 Hints . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 50 `hints' file. . . . . . . . . . . . . . . . . . . . . . . . 50 hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 htonl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 htons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Huntgroup-Name . . . . . . . . . . . . . . . . . . . . 186 Huntgroups . . . . . . . . . . . . . . . . . . . . . . 15, 50 `huntgroups' file . . . . . . . . . . . . . . . . . . 50

I

ident . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Identifiers, Rewrite. . . . . . . . . . . . . . . . . . 108 Idle-Timeout . . . . . . . . . . . . . . . . . . . . . . 169 index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 inet_aton . . . . . . . . . . . . . . . . . . . . . . . . . 112 inet_ntoa . . . . . . . . . . . . . . . . . . . . . . . . . 112 input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 input-format . . . . . . . . . . . . . . . . . . . . . . . 38

Invoking Scheme authentication

function . . . . . . . . . . . . . . . . . . . . . . 117 Invoking the radius daemon . . . . . . . . . . 17

IP pools for MAX Ascend . . . . . . . . . . . . 105

L

Label, Matching Rule . . . . . . . . . . . . . . . . 11 length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 LHS, Matching Rule . . . . . . . . . . . . . . . . . 11 listen . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 30 load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 load-module . . . . . . . . . . . . . . . . . . . . . . . . 35 load-path. . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Local Password Auth . . . . . . . . . . . . . . . . 71 Log directory . . . . . . . . . . . . . . . . . . . . . . . . . 5 log-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Log-Mode-Mask . . . . . . . . . . . . . . . . . . . . . 186 logging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Logging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 Logging category . . . . . . . . . . . . . . . . . . . . 26 Logging channel . . . . . . . . . . . . . . . . . . . . . 26 Logging hook . . . . . . . . . . . . . . . . . . . . . . . . 25 logging statement . . . . . . . . . . . . . . . . . . 27 Logging, ‘config’ statement . . . . . . . . . . 24 login . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

Login verification functions . . . . . . 104

Login-Time . . . . . . . . . . . . . . . . . . . . . . . . 187 logit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

M

                                        master-read-timeout . . . . . . . . . . . . . . . . 22
                                        master-write-timeout. . . . . . . . . . . . . . . 22

Match-Profile . . . . . . . . . . . . . . . . . . . . . 188 Matching Rule. . . . . . . . . . . . . . . . . . . . . . . 11 MAX Ascend, broken passwords . . . . . . 45 max-nas-count . . . . . . . . . . . . . . . . . . . . . . 32 max-port-count . . . . . . . . . . . . . . . . . . . . . 32 max-processes . . . . . . . . . . . . . . . . . . . . . . 22 max-requests . . . . . . . . . . . . . 22, 28, 30, 32 Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 menu, syntax . . . . . . . . . . . . . . . . . . . . . . . . 62 `menus' file. . . . . . . . . . . . . . . . . . . . . . . . 62 ‘menus’, configuration subdirectory …. 62 message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Messages: configuring . . . . . . . . . . . . . . . . 36 mlc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Multiple Login Checking . . . . . . . . . . . . . 74 multiple-login . . . . . . . . . . . . . . . . . . . . . 36

N

Naming conventions . . . . . . . . . . . . . . . . . . . 5 nas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 201 NAS types, standard . . . . . . . . . . . . . . . . . 49 nas-address . . . . . . . . . . . . . . . . . . . . . . . 124 NAS-Identifier . . . . . . . . . . . . . . . . . . . . 170 NAS-IP-Address . . . . . . . . . . . . . . . . . . . . 170 nas-port. . . . . . . . . . . . . . . . . . . . . . . . . . . 124 NAS-Port-Id . . . . . . . . . . . . . . . . . . . . . . . 170 NAS-Port-Type . . . . . . . . . . . . . . . . . . . . . 171 nas.scm. . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 ‘naslist’ file . . . . . . . . . . . . . . . . . . . . . . . . 45 `naslist' file. . . . . . . . . . . . . . . . . . . . . . 46 ‘nastypes’ file . . . . . . . . . . . . . . . . . . . . . . . 47 `nastypes' file . . . . . . . . . . . . . . . . . . . . 48 ‘nastypes’ file, syntax of . . . . . . . . . . . . . 47 network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Network Access Server . . . . . . . . . . . . 7, 201 newline. . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 ngettext. . . . . . . . . . . . . . . . . . . . . . . . . . . 114 ‘NOREALM’, special realm name. . . . . . . . . 50 ntohl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 ntohs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

O

option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 orig-login . . . . . . . . . . . . . . . . . . . . . . . . 124 outfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

P

PAM Authentication Type . . . . . . . . . . . 73 Pam-Auth. . . . . . . . . . . . . . . . . . . . . . . . . . . 189 pam_radius.so . . . . . . . . . . . . . . . . . . . . . 163

password-expire-warning . . . . . . . . 28, 36

password-expired . . . . . . . . . . . . . . . . . . . 36 perms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 port . . . . . . . . . . . . . . . . . . . . . . . . . 28, 30, 32 port-type . . . . . . . . . . . . . . . . . . . . . . . . . 124 Prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 prefix-hook . . . . . . . . . . . . . . . . . . . . . 25, 26 print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 print-auth . . . . . . . . . . . . . . . . . . . . . . . . . 26 print-category . . . . . . . . . . . . . . . . . . . . . 26 print-cons . . . . . . . . . . . . . . . . . . . . . . . . . 26 print-failed-pass . . . . . . . . . . . . . . . . . . 26 print-level . . . . . . . . . . . . . . . . . . . . . . . . 26 print-milliseconds . . . . . . . . . . . . . . . . . 26 print-pass . . . . . . . . . . . . . . . . . . . . . . . . . 26 print-pid. . . . . . . . . . . . . . . . . . . . . . . . . . . 26 print-priority . . . . . . . . . . . . . . . . . . . . . 26 print-tid. . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Problem Tracking . . . . . . . . . . . . . . . . . . . . 87

process-idle-timeout. . . . . . . . . . . . . . . 22

Processing requests . . . . . . . . . . . . . . . . . . 12 Propagation of an attribute . . . . . . . . . . . . 7 Properties of an attribute. . . . . . . . . . . . . . 7 PROPERTY. . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 protocol. . . . . . . . . . . . . . . . . . . . . . . . . . . 124 PROXY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 Proxy Service. . . . . . . . . . . . . . . . . . . . . . . . 13 Proxy-Replied . . . . . . . . . . . . . . . . . . . . . 190 Proxying . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

Q

q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 qprn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 query-nas. . . . . . . . . . . . . . . . . . . . . . . . . . . 91 quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 quote_string . . . . . . . . . . . . . . . . . . . . . . 113

R

r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 rad-add-server . . . . . . . . . . . . . . . . . . . . 134

rad-client-add-server. . . . . . . . . . . . . 133
rad-client-list-servers . . . . . . . . . . 132

rad-client-retry . . . . . . . . . . . . . . . . . . 133

rad-client-set-server. . . . . . . . . . . . . 132
rad-client-source-ip. . . . . . . . . . . . . . 133

rad-client-timeout . . . . . . . . . . . . . . . . 133 rad-closelog . . . . . . . . . . . . . . . . . . . . . . 119

rad-dict-name->attr . . . . . . . . . . . . . . . 118
rad-dict-name->value. . . . . . . . . . . . . . 119
rad-dict-pec->vendor. . . . . . . . . . . . . . 119
rad-dict-value->name. . . . . . . . . . . . . . 118

rad-format-code . . . . . . . . . . . . . . . . . . . 133 rad-format-pair . . . . . . . . . . . . . . . . . . . 133

rad-format-reply-msg. . . . . . . . . . . . . . 134

rad-get-server . . . . . . . . . . . . . . . . . . . . 132 rad-list-servers . . . . . . . . . . . . . . . . . . 134 rad-log-close . . . . . . . . . . . . . . . . . . . . . 119 rad-log-open . . . . . . . . . . . . . . . . . . . . . . 119 rad-openlog . . . . . . . . . . . . . . . . . . . . . . . 119 rad-print-pairs . . . . . . . . . . . . . . . . . . . 133 rad-read-no-echo . . . . . . . . . . . . . . . . . . 133

rad-rewrite-execute . . . . . . . . . . . . . . . 119
rad-rewrite-execute-string . . . . . . . 119

rad-select-server . . . . . . . . . . . . . . . . . 134 rad-send. . . . . . . . . . . . . . . . . . . . . . . . . . . 132 rad-send-internal . . . . . . . . . . . . . . . . . 132 rad-server-list . . . . . . . . . . . . . . . . . . . 131 rad-syslog . . . . . . . . . . . . . . . . . . . . . . . . 119 rad-utmp-putent . . . . . . . . . . . . . . . . . . . 119 ‘radacct’, accounting directory . . . . . . . . 5 radauth. . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 radctl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 ‘raddb’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘raddb/access.deny’ file . . . . . . . . . . . . . 53 ‘raddb/client.conf’. . . . . . . . . . . . . . . . 135 ‘raddb/clients’ file. . . . . . . . . . . . . . . . . . 44 ‘raddb/config’ file . . . . . . . . . . . . . . . . . . . 22 ‘raddb/hints’ file . . . . . . . . . . . . . . . . . . . . 50 ‘raddb/huntgroups’ file . . . . . . . . . . . . . . 50 ‘raddb/menus’, configuration subdirectory

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 ‘raddb/naslist’ file. . . . . . . . . . . . . . . . . . 45 ‘raddb/realms’ file . . . . . . . . . . . . . . . . . . . 50 ‘raddb/rewrite’, configuration file …. 61 ‘raddb/sqlserver’ file. . . . . . . . . . . . . . . . 53 ‘raddb/users’ file . . . . . . . . . . . . . . . . . . . . 52 radgrep. . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 Radius daemon invocation . . . . . . . . . . . . 17 Radius dictionary . . . . . . . . . . . . . . . . . . . . 40 Radius-Specific Scheme Functions …. 118 radiusd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Radiusd configuration . . . . . . . . . . . . . . . . 22 Radiusd configuration files. . . . . . . . . . . . 21 radiusd-user . . . . . . . . . . . . . . . . . . . . . . . 22 radlast. . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 radlast, options . . . . . . . . . . . . . . . . . . 126 ‘radlog’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 radping. . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 radscm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 radsession . . . . . . . . . . . . . . . . . . . . . . . . 161 radtest. . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 radwho . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 radwho, command line options . . . . . . . 121 radwho, format strings . . . . . . . . . . . . . . 123 radwho, predefined formats . . . . . . . . . . 125 radzap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 rd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 realm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 Realm-Name . . . . . . . . . . . . . . . . . . . . . . . . 190 realm-quota . . . . . . . . . . . . . . . . . . . . . . . . 36 Realms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ‘realms’ file . . . . . . . . . . . . . . . . . . . . . . . . . 50 `realms' file. . . . . . . . . . . . . . . . . . . . . . . 52 Regular Expressions, Rewrite . . . . . . . . 110 Reject Authentication Type . . . . . . . . . . 71 Replace-User-Name . . . . . . . . . . . . . . . . . 190 REPLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 Reply-Message . . . . . . . . . . . . . . . . . . . . . 171 Request . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 request queue, configuring . . . . . . . . . . . . 69

request-cleanup-delay . . . . . . . 28, 30, 32

request-define . . . . . . . . . . . . . . . . . . . . . 93 request-print . . . . . . . . . . . . . . . . . . . . . . 94 request_code . . . . . . . . . . . . . . . . . . . . . . 115

request_code_string . . . . . . . . . . . . . . . 113

request_id . . . . . . . . . . . . . . . . . . . . . . . . 115 request_source_ip . . . . . . . . . . . . . . . . . 114

request_source_port . . . . . . . . . . . . . . . 115

Requests, accounting . . . . . . . . . . . . . . . . . . 9 Requests, authentication . . . . . . . . . . . . . . 8 resolve . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 return . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 RETURN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 rewrite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Rewrite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 Rewrite functions . . . . . . . . . . . . . . . . . 100 Rewrite identifiers . . . . . . . . . . . . . . . . . . 108 Rewrite language settings . . . . . . . . . . . . 35 Rewrite, applying functions . . . . . . . . . . . 99 Rewrite, attribute creation functions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 ‘rewrite’, configuration file . . . . . . . . . . . 61 Rewrite, data types . . . . . . . . . . . . . . . . . 105 Rewrite, Logging Hook Functions …. 105 Rewrite, login verification functions . . 103 Rewrite, quick start introduction . . . . . . 98 Rewrite, symbols . . . . . . . . . . . . . . . . . . . 106 Rewrite, syntax of the language . . . . . . 105 Rewrite, syntax overview . . . . . . . . . . . . . 98 Rewrite, usage . . . . . . . . . . . . . . . . . . . . . . . 99 Rewrite-Function . . . . . . . . . . . . . . . . . . 191 rewrite-stack . . . . . . . . . . . . . . . . . . . . . . 92 Rewriting incoming requests . . . . . . . . . . 99 RHS, Matching Rule . . . . . . . . . . . . . . . . . 11 rindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 rp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 rs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 Rule Tracing . . . . . . . . . . . . . . . . . . . . . . . . 87 run-rewrite . . . . . . . . . . . . . . . . . . . . . . . . 92 Run-time options (radiusd) . . . . . . . . . . . 22

S

s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

                                        Scheme accounting function . . . . . . . . 118
                                        Scheme authentication function. . . . 116
                                        Scheme authentication function,

invocation . . . . . . . . . . . . . . . . . . . . 117

                                        Scheme-Acct-Procedure. . . . . . . . . . . . . 191

Scheme-Procedure . . . . . . . . . . . . . . . . . . 191 second-login . . . . . . . . . . . . . . . . . . . . . . . 36 send . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 Service. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 Service-Type . . . . . . . . . . . . . . . . . . . . . . 171 Session. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 Session ID. . . . . . . . . . . . . . . . . . . . . . . . . . 201 session-id . . . . . . . . . . . . . . . . . . . . . . . . 124 Session-Timeout . . . . . . . . . . . . . . . . . . . 173 set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Simultaneous logins, checking for. . . . . . 74 Simultaneous-Use . . . . . . . . . . . . . . . . . . 192 snmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 SNMP service parameters . . . . . . . . . . . . 32 source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 source-ip. . . . . . . . . . . . . . . . . . . . . . . . . . . 22 sql . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 SQL Accounting . . . . . . . . . . . . . . . . . . . . . 83 SQL accounting query templates . . . . . . 58 SQL accounting query templates, writing

of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 SQL Authentication Type . . . . . . . . . . . . 73 ‘sqlserver’ file. . . . . . . . . . . . . . . . . . . . . . 53 State . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 Statements, Rewrite . . . . . . . . . . . . . . . . 109 storage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 strip-names . . . . . . . . . . . . . . . . . . . . . . . . 28 Strip-User-Name . . . . . . . . . . . . . . . . . . . 192 substr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Suffix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 suffix-hook . . . . . . . . . . . . . . . . . . . . . 25, 26 Symbols, Rewrite . . . . . . . . . . . . . . . . . . . 106 Syntax of ‘nastypes’ . . . . . . . . . . . . . . . . . 47 syslog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 System Accounting . . . . . . . . . . . . . . . . . . 81 System Authentication Type. . . . . . . . . . 72

T

t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 tab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 Termination-Action . . . . . . . . . . . . . . . . 173 Termination-Menu . . . . . . . . . . . . . . . . . . 193 Test Mode . . . . . . . . . . . . . . . . . . . . . . . . . . 90 textdomain . . . . . . . . . . . . . . . . . . . . . . . . 114 time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 time-to-live . . . . . . . . . . . . . . . . 28, 30, 32 timespan. . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 timespan-violation . . . . . . . . . . . . . . . . . 36 tolower. . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 toupper. . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

U

unquote_string . . . . . . . . . . . . . . . . . . . . 113 usedbm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 User Profiles. . . . . . . . . . . . . . . . . . . . . . . . . 16 User-Name . . . . . . . . . . . . . . . . . . . . . . . . . 174 User-Password . . . . . . . . . . . . . . . . . . . . . 174 username-chars . . . . . . . . . . . . . . . . . . . . . 22 `users' file. . . . . . . . . . . . . . . . . . . . . . . . 52 Utility Programs. . . . . . . . . . . . . . . . . . . . 121 utmp-entry . . . . . . . . . . . . . . . . . . . . . . . . 119

V

VALUE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 VENDOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Vendor-Specific . . . . . . . . . . . . . . . . . . . 175

W

wait-reply . . . . . . . . . . . . . . . . . . . . . . . . . 38 Writing SQL accounting query templates

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

Short Contents

Introduction to Radius. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Naming Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2 How Radius Operates . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3 How to Start the Daemon. . . . . . . . . . . . . . . . . . . . . . . . . 17 4 Radius Configuration Files . . . . . . . . . . . . . . . . . . . . . . . . 21 5 Request Comparison Methods . . . . . . . . . . . . . . . . . . . . . . 67 6 Authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 7 Accounting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 8 Logging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 9 Problem Tracking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 10 Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 11 Utility Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 12 Client Package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 13 Attribute List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 14 Reporting Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 15 Where to Get Information about GNU Radius . . . . . . . . . . 197 How to Obtain Radius . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 Radius Glossary. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 16 New Configuration Approach (draft) . . . . . . . . . . . . . . . . 205 A GNU Free Documentation License . . . . . . . . . . . . . . . . . . 215 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 ii GNU Radius Reference Manual

Table of Contents

Introduction to Radius . . . . . . . . . . . . . . . . . . . . . . . . 1

Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Naming Conventions . . . . . . . . . . . . . . . . . . . . . . . 5

2 How Radius Operates . . . . . . . . . . . . . . . . . . . . . . 7

2.1 Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2 radius Requests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2.2.1 Authentication Requests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2.2 Accounting Requests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2.3 Matching Rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.4 Processing Requests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

2.4.1 Checking for Duplicate Requests . . . . . . . . . . . . . . . . . . . . . . . . 12 2.4.2 Proxying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

2.4.2.1 Proxy Service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.4.2.2 Realms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

2.4.3 Hints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.4.4 Huntgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.4.5 User Profiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

3 How to Start the Daemon. . . . . . . . . . . . . . . . . 17

4 Radius Configuration Files . . . . . . . . . . . . . . . . 21

4.1 Run-Time Configuration Options — ‘raddb/config’ . . . . . . . . . . 22 4.1.1 option block . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.1.2 logging block . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

4.1.2.1 Logging hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.1.2.2 category statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.1.2.3 channel statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.1.2.4 Example of the logging statement . . . . . . . . . . . . . . . . . 27

4.1.3 auth statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 4.1.4 acct statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 4.1.5 usedbm statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 4.1.6 snmp statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 4.1.7 rewrite statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.1.8 guile statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.1.9 message statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.1.10 filters statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.1.11 mlc statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

4.2 Dictionary of Attributes — ‘raddb/dictionary’ . . . . . . . . . . . . . . 39 4.2.1 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.2.2 $INCLUDE Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.2.3 VENDOR Statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.2.4 ATTRIBUTE statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.2.5 Blocks of Vendor-Specific Attributes. . . . . . . . . . . . . . . . . . . . . 42 4.2.6 ALIAS statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.2.7 PROPERTY statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.2.8 VALUE Statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

4.3 Clients List — ‘raddb/clients’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

4.3.1 Example of ‘clients’ file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 4.4 NAS List — ‘raddb/naslist’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

4.4.1 Example of ‘naslist’ file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 4.5 NAS Types — ‘raddb/nastypes’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

4.5.1 Syntax of ‘raddb/nastypes’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 4.5.2 Example of nastypes file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 4.5.3 Standard nas types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

4.6 Request Processing Hints — ‘raddb/hints’ . . . . . . . . . . . . . . . . . . 49

4.6.1 Example of ‘hints’ file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.7 Huntgroups — ‘raddb/huntgroups’ . . . . . . . . . . . . . . . . . . . . . . . . . 50

4.7.1 Example of ‘huntgroups’ file. . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.8 List of Proxy Realms — ‘raddb/realms’ . . . . . . . . . . . . . . . . . . . . . 50

4.8.1 Example of ‘realms’ file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 4.9 User Profiles — ‘raddb/users’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

4.9.1 Example of ‘users’ file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 4.10 List of Blocked Users — ‘raddb/access.deny’ . . . . . . . . . . . . . . 53 4.11 SQL Configuration — ‘raddb/sqlserver’ . . . . . . . . . . . . . . . . . . . 53

4.11.1 SQL Client Parameters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 4.11.2 Authentication Server Parameters. . . . . . . . . . . . . . . . . . . . . . 54 4.11.3 Authorization Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 4.11.4 Accounting Parameters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

4.11.4.1 Writing SQL Accounting Query Templates . . . . . . . . . 58 4.12 Rewrite functions — ‘raddb/rewrite’ . . . . . . . . . . . . . . . . . . . . . . 61 4.13 Login Menus — ‘raddb/menus’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

4.13.1 A menu file syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 4.13.2 An example of menu files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

4.14 Macro Substitution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

5 Request Comparison Methods . . . . . . . . . . . . . 67

5.1 Extended Comparison . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 5.1.1 An example of extended comparison configuration . . . . . . . . 68 5.1.2 List of attributes that can be declared comparable. . . . . . . . 68

5.2 Fine-Tuning the Request Queue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

6 Authentication . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

6.1 Accept Authentication Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 6.2 Reject Authentication Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 6.3 Local Password Authentication Type . . . . . . . . . . . . . . . . . . . . . . . . 71 6.4 Encrypted Password Authentication Type . . . . . . . . . . . . . . . . . . . . 72 6.5 System Authentication Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 6.6 SQL Authentication Type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 6.7 PAM Authentication Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 6.8 Defining Custom Authentication Types . . . . . . . . . . . . . . . . . . . . . . 73 6.9 Multiple Login Checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

6.9.1 Retrieving Session Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 6.9.2 Verifying Active Sessions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

6.10 Controlling Authentication Probes. . . . . . . . . . . . . . . . . . . . . . . . . . 76

7 Accounting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

7.1 System Accounting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 7.2 Detailed Request Accounting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 7.3 sql Accounting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 7.4 Defining Custom Accounting Types. . . . . . . . . . . . . . . . . . . . . . . . . . 83

8 Logging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

9 Problem Tracking. . . . . . . . . . . . . . . . . . . . . . . . . 87

9.1 Rule Tracing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 9.2 Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 9.3 Test Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

10 Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

10.1 Filters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 10.1.1 Getting Acquainted with Filters . . . . . . . . . . . . . . . . . . . . . . . 95 10.1.2 Declaring the Filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 10.1.3 Invoking the Filter from a User Profile . . . . . . . . . . . . . . . . . 96 10.1.4 Adding Reply Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 10.1.5 Accounting Filters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 10.1.6 Invoking the Accounting Filter. . . . . . . . . . . . . . . . . . . . . . . . . 98

10.2 Rewrite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 10.2.1 Syntax Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 10.2.2 Quick Start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 10.2.3 Interaction with Radius . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 10.2.4 Rewriting Incoming Requests . . . . . . . . . . . . . . . . . . . . . . . . . . 99

10.2.4.1 Examples of Various Rewrite Functions . . . . . . . . . . . 100 10.2.5 Login Verification Functions . . . . . . . . . . . . . . . . . . . . . . . . . . 103

10.2.5.1 Examples of Login Verification Functions . . . . . . . . . 104 10.2.6 Attribute Creation Functions . . . . . . . . . . . . . . . . . . . . . . . . . 104 10.2.7 Logging Hook Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 10.2.8 Full Syntax Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

10.2.8.1 Rewrite Data Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 10.2.8.2 Rewrite Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 10.2.8.3 Rewrite Identifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 10.2.8.4 Rewrite Declarations . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 10.2.8.5 Rewrite Statements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 10.2.8.6 Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 10.2.8.7 Rewrite Built-in Functions . . . . . . . . . . . . . . . . . . . . . . . 111

10.3 Guile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 10.3.1 Data Representation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 10.3.2 Authentication with Scheme . . . . . . . . . . . . . . . . . . . . . . . . . . 116 10.3.3 Accounting with Scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 10.3.4 Radius-Specific Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

11 Utility Programs . . . . . . . . . . . . . . . . . . . . . . . 121

11.1 radwho . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 11.1.1 radwho Command Line Options . . . . . . . . . . . . . . . . . . . . . . 121 11.1.2 radwho Format Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 11.1.3 radwho Predefined Formats. . . . . . . . . . . . . . . . . . . . . . . . . . . 125

11.2 radlast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

11.2.1 radlast Command Line Options . . . . . . . . . . . . . . . . . . . . . 126 11.3 radzap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 11.4 radgrep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 11.5 radping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 11.6 radauth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 11.7 radctl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 11.8 builddbm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 11.9 radscm: A Guile Interface to Radius Functions . . . . . . . . . . . . . 131

12 Client Package . . . . . . . . . . . . . . . . . . . . . . . . . 135

12.1 Client Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 12.2 radtest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

12.2.1 Invoking radtest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 12.2.2 Literal Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

12.2.2.1 Numeric Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 12.2.2.2 Character Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 12.2.2.3 Lists of A/V pairs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

12.2.3 Reserved Keywords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 12.2.4 Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

12.2.4.1 Using Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 12.2.4.2 Variable Assignments. . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 12.2.4.3 Dereferencing Variables . . . . . . . . . . . . . . . . . . . . . . . . . . 140 12.2.4.4 Accessing Elements of A/V Pair Lists . . . . . . . . . . . . 142 12.2.4.5 Assignment Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

12.2.4.6 Built-in Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 12.2.5 Positional Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 12.2.6 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144

12.2.6.1 Arithmetic Operations. . . . . . . . . . . . . . . . . . . . . . . . . . . 144 12.2.6.2 String Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 12.2.6.3 Operations on A/V Lists . . . . . . . . . . . . . . . . . . . . . . . . 145 12.2.6.4 Comparison Operations. . . . . . . . . . . . . . . . . . . . . . . . . . 146 12.2.6.5 Boolean Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 12.2.6.6 Conversion Between Data Types . . . . . . . . . . . . . . . . . 147 12.2.6.7 Function Calls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 12.2.6.8 Operator Precedence (How Operators Nest) . . . . . . . 148

12.2.7 Function Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 12.2.8 Interacting with Radius Servers . . . . . . . . . . . . . . . . . . . . . . . 149 12.2.9 Conditional Statements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 12.2.10 Loops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 12.2.11 Built-in Primitives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 12.2.12 Sample Radtest Program. . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

12.3 radsession . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 12.4 nas.scm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 12.5 pam_radius.so . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163

13 Attribute List . . . . . . . . . . . . . . . . . . . . . . . . . . 165

13.1 Authentication Attributes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 13.1.1 CHAP-Password. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 13.1.2 Callback-Id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 13.1.3 Callback-Number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 13.1.4 Called-Station-Id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 13.1.5 Calling-Station-Id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 13.1.6 Class. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 13.1.7 Framed-Compression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 13.1.8 Framed-IP-Address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 13.1.9 Framed-IP-Netmask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 13.1.10 Framed-MTU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 13.1.11 Framed-Protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 13.1.12 Framed-Route. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 13.1.13 Framed-Routing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 13.1.14 Idle-Timeout. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 13.1.15 NAS-IP-Address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 13.1.16 NAS-Identifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 13.1.17 NAS-Port-Id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 13.1.18 NAS-Port-Type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 13.1.19 Reply-Message. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 13.1.20 Service-Type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 13.1.21 Session-Timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 13.1.22 State . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 13.1.23 Termination-Action . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 13.1.24 User-Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 13.1.25 User-Password. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 13.1.26 Vendor-Specific . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175

13.2 Accounting Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 13.2.1 Acct-Authentic. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 13.2.2 Acct-Delay-Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 13.2.3 Acct-Input-Octets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 13.2.4 Acct-Input-Packets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 13.2.5 Acct-Output-Octets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 13.2.6 Acct-Output-Packets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 13.2.7 Acct-Session-Id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 13.2.8 Acct-Session-Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 13.2.9 Acct-Status-Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 13.2.10 Acct-Terminate-Cause . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178

13.3 Radius Internal Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 13.3.1 Acct-Ext-Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 13.3.2 Acct-Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 13.3.3 Auth-Failure-Trigger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 13.3.4 Auth-Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 13.3.5 Auth-Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 13.3.6 Crypt-Password. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 13.3.7 Exec-Program-Wait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

13.3.7.1 Running an External Program . . . . . . . . . . . . . . . . . . . 182 13.3.7.2 Using an External Filter . . . . . . . . . . . . . . . . . . . . . . . . . 182

13.3.8 Exec-Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 13.3.9 Fall-Through . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 13.3.10 Group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 13.3.11 Hint. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 13.3.12 Huntgroup-Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 13.3.13 Log-Mode-Mask. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 13.3.14 Login-Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 13.3.15 Match-Profile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 13.3.16 Menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 13.3.17 Pam-Auth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 13.3.18 Prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 13.3.19 Proxy-Replied. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 13.3.20 Realm-Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 13.3.21 Replace-User-Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 13.3.22 Rewrite-Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 13.3.23 Scheme-Acct-Procedure . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 13.3.24 Scheme-Procedure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 13.3.25 Simultaneous-Use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 13.3.26 Strip-User-Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 13.3.27 Suffix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 13.3.28 Termination-Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193

14 Reporting Bugs . . . . . . . . . . . . . . . . . . . . . . . . 195

15 Where to Get Information about GNU Radius . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197

How to Obtain Radius . . . . . . . . . . . . . . . . . . . . . . 199

Radius Glossary . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201

Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . 203

16 New Configuration Approach (draft). . . . . 205

16.1 A brief description of Currently Used Approach . . . . . . . . . . . . 205 16.2 Deficiencies of Current Operation Model and Configuration Suite

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 16.3 Proposed Solution. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207

16.3.1 Request-processing Instruction. . . . . . . . . . . . . . . . . . . . . . . . 207 16.3.2 grad instr conditional. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 16.3.3 grad instr call. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 16.3.4 grad instr return . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 16.3.5 grad instr action . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 16.3.6 grad instr reply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 16.3.7 grad instr proxy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 16.3.8 grad instr forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 16.3.9 RPL representation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212

16.4 Changes to Rewrite Language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 16.5 Support for Traditional Configuration Files. . . . . . . . . . . . . . . . . 212 16.6 New Configuration Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213

Appendix A GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215

A.0.1 ADDENDUM: How to use this License for your documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

x GNU Radius Reference Manual